A Complete Bibliography of ACM Transactions on Applied Perception

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

04 October 2025 Version 1.67

Title word cross-reference

2 [KPSL10, KHJK13, LPR06]. 3 [AW15, AWR18, BSW10, BSHW14, EBPJ16, GLT05b, GVC^{+}17, HOH15, HAHG17, JDR08, KPSL10, LDDR18, Lav09, MSS^{+}22, NFD^{+}21, OR04, Ste15, SK16, VSKL17, WPDH14, WBNF06]. ^{∘} [WIA24]. K [KDK^{+}16].

1994 [Bar05a]. 1996 [FTB05, WK05b]. 1997 [Bre05, MPC05]. 1998 [BM05, EM05a, Fer05].

2000 [Vic05]. 2001 [SCS05]. 2002 [HR05b]. 2003 [BS05a, GLT05a]. 2004 [vdD05a]. 2010 [MB10]. 2012 [MS12]. 2013 [GS13]. 2015 [KMS15]. 2016 [BT16]. 2018 [RK18].

2019 [HC19]. 2020 [GGFE20]. 2021 [JO21]. 2022 [SBC22]. 2024 [LP24]. 2025 [DB25]. 20s [KVE16].

Abstract [SNW16, ZZ13]. Abstraction [SGHL^{+}19]. abstractions [MMS13]. Acceleration [RKS16]. accelerations [BSPB10]. Acceptability [MP20]. Accuracy [UHWT21]. Achieved [FvdMV25]. acoustic [FR08, LZG^{+}13]. Acquisition [BFCP^{+}25, GHS^{+}20]. Across [AONB17, TGJ08]. Acting [JOY^{+}18]. Action [BB21, GRM^{+}21, JOZ^{+}21, PR24, BSW10, ECOG11, NAB^{+}11, ZNWK12, ZAAC12]. action-based [ZNWK12]. Actions [NP15, NOSS17, KWSS08]. activation [DKR^{+}05]. activation-based [DKR^{+}05]. activations [VSCM12]. active [TSRD07]. activity [BWG12]. actual [RFR09]. Actuators [ÆÁP^{+}22]. Adaptation [PUC^{+}24, RBC14]. Adaptive [BMB19, FO23, LFM12]. addressing [GTAE04]. ADHD [JCM^{+}25]. Adjustment [JWA19, JDKN18]. Adult [JAA^{+}16]. Adults [MVL25]. advantages [WWA11]. Aesthetic [ZYZ^{+}17, TJL^{+}11]. Aesthetics [HZD23, ZYJ^{+}22]. Affect [CR22, LCC15, GNP^{+}10]. affected [TSC13]. Affective [LCC15, SCM18, WKM^{+}17, KWI09, MDR10]. Affects [LVV^{+}20, ZOH^{+}15, ZNWK12]. Affordance [LRB15]. After [FM05, SB12]. Agency [BLKD24]. Agent [AONB17, EBA^{+}21, FGWdS^{+}25, JOZ^{+}21]. aid [JSHG08, RDF11]. aided [SDW05]. Air [VVC^{+}22, HDF^{+}23, MHPM23, PCK08, SHHF^{+}22]. al [BM05, FTB05, SCS05]. Albums [KPL^{+}19]. algorithm [MP09, SGF^{+}10]. Algorithmic [CPVC19, WKM^{+}15]. Algorithms [HBM^{+}14, HFJS09, KYL^{+}07, OAD^{+}12, PJN^{+}11]. Alignment [USA20]. Allowing [BDR^{+}21]. Alphabet [PVK20]. Alty [Vic05]. Ambient [DBR21, KDK^{+}16, DCR06]. Ambient/Focal [DBR21, KDK^{+}16]. Ambiguities [WBHP20]. American [FWS^{+}23, MVL25]. Among [FWS^{+}23, MVL25]. Amplitude [BFCP^{+}25, MTÆ^{+}24]. Analysis [ASG^{+}18, BCD15, BH17, CLR10, EKL^{+}21, FZL20, FK19, MMSO15, NOSS17, RTSW18, TMM17, BMGC05, CWB10, FCH09, FBT05, MB04, MP09, TKK^{+}13]. Analyzing [NGJT13]. Angiography [ABK^{+}15]. Angle [Ste15, HS12, TSRD07]. Animals [SNW16]. Animated [HCKH16, BAMB13, HJO^{+}10, WBCB08]. animating [TCMH11]. animation [RO09]. Animations [KMH^{+}19, KVDE19, VHBO14]. anomalies [HJO^{+}10]. Anonymization [WIA24].

APGV [MB10]. APGV07 [FL09].

Appearance

[APK15, FWS^{+}23, KHH17, WDLC24]. Application [SHHF^{+}22, AMR06, Lav09]. applications [KHJK13]. Applied [FL09]. Applying [MB04, PI08, YBC13]. Approach [SJ18, TB24, BGL^{+}08, GEMA13, MO09b]. Appropriate [IOYK19]. approximate [YCK^{+}09]. approximating [JDR08]. Area [MW15]. arm [VGBF10]. arousal [VSCM12]. Arrow [USA20]. Art [SNW16, SMO^{+}10]. Arteriovenous [ABK^{+}15]. artifact [TKK^{+}13]. Artifacts [HOH15, PMS17, TMM17, MGM12]. Artistic [WBHP20]. Artists [SNW16]. Arts [PHRE15]. assess [SAB07]. assessed [VCR08]. Assessing [BLKD24, HBW11, HBF16, TCGC19, WTWN16]. Assessment [BMB19, GVC^{+}17, NFD^{+}21, VSKL17, ZLQ^{+}19, APP07, DCN^{+}06, WK05a]. Assistants [RL17]. Assisted [DCRS15]. Association [KNL22]. assumptions [MRT^{+}10]. asymmetric [SMI06]. Atmosphere [YO25]. Attacks [MJP^{+}24]. Attention [BFSV16, BH17, DBR21, HHNOP19, KDK^{+}16, O’S05, ZYJ^{+}22, FRC10, GMT09, HCS10, RTPG11]. Attentional [RTPG11]. Attraction [FWS^{+}23]. Attractiveness [OEMO16, ZNO^{+}20]. Attributes [FK19, LZL17]. Audio [GBTB25, MSHLR16, RBCK12, BSVDD10, LZG^{+}13]. audio-visual [BSVDD10, LZG^{+}13, RBCK12]. audiovisual [BJK13, GLT05b]. Auditory [AL15, CR22, FR08, GWFL22, OR04, RFR09, AZ10, Bar05a, Bar05b, BMGC05, BM05, EM05a, GDBP13, GLT05b, KW05, MB04, NW08, SCSG05, SCS05, WK05a, DFJ^{+}20]. augmentations [KMHO13]. Augmented [AL15, GRM^{+}21, GSCRB23, RVH^{+}19, WC22, WDLC24, HFJS09, KWSS08, KMHO13, SDW05]. augmented-reality [KWSS08]. auralization [VA05, Vic05]. Author [GLT05a, MPC05, HR05b, Vic05]. Automated [HBK^{+}21, KDCM15, SMO^{+}10]. Automatic [PMS17, McN06, RVB05, TVR^{+}11]. auxiliary [KMHO13]. Avatar [HHB^{+}24, KS12]. avoidance [FFW07]. Awareness [OMN25, MBCW10, ZCRTW12].

Background [LMM^{+}22]. Balance [KKC19, WBN^{+}11]. Balancing [KDCM15]. Barrass [Bar05a]. Base [RSM^{+}15]. Based [BLKD24, BFSV16, BYB18, CSUN05, GBA17, KPDP19, KHH17, KVDE19, LRB15, LZL^{+}18, NPKR23, OMN25, RTJ^{+}21, SYF^{+}25, SXCS15, SJ18, SWT^{+}23, TUG^{+}20, WL21, BS06, BMB19, DKR^{+}05, ENC^{+}08, HR05b, HVM06, HDH10, JWA19, Kaw19, KBP^{+}13, MI07, MC05, MTCR^{+}07, MDT09, RLH^{+}08, SD22, SZ22, SMS13, WMA12, ZC06, ZLO13, ZNWK12, ZLQ^{+}19, vdD05b]. Bathing [MNG^{+}25]. Bayesian [ECOG11]. be [FBT05, RVB05]. Beauty [CXZ14]. Behavior [ASG^{+}18, BBE16, BBE22, KPR22, KPL^{+}19, RPH10]. Behavioral [MVL25, OOAY22]. Behaviors [EBPJ16, SBR07]. believable [BSPB10]. belt [VVJD05]. Benefits [EKL^{+}21, Ste15]. Better [IOYK19]. between [NGJT13, PD17, SNW16, SZ22, WNW^{+}07]. bias [HCS10]. Biases [WZA^{+}23]. bidirectional [FCH09]. Big [CEN^{+}23, JSCR^{+}15]. Bimodal [BGW11, BSVDD10]. binning [ECOG11]. Biological [NOSS17]. Biomechanics [JLS^{+}17]. Biometric [RKS16, KHJK13]. Biometrics [KHKP15]. Black [MVL25]. Blending [FO23, CLR12]. Blind [BDR^{+}21, RSM^{+}15]. Blindness [KT21]. blinks [TCMH11]. Blur [LVV^{+}20, SK16, BJK13, LFM12, NCSG11]. board [WBN^{+}11]. bodies [MEDO09].

Body

[GFD^{+}15, KHW^{+}15, KLL24, KSLM15, BWG12, MJH^{+}09, MJM^{+}09, NGJT13]. body-worn [BWG12]. Bonebright [BM05]. both [NAB^{+}11]. Bouncing [DFJ^{+}20]. Braille [LPHL05]. Brain [TB24, VCA16]. Brain-Computer [VCA16]. Brewster [Bre05]. Brightness [CKAD18]. browsing [Fer05, LZG^{+}13]. Brungart [BS05a]. bumps [LBT08, VVHV10]. Button [KBL14].

Calibrated [GSCRB23]. Calibration [KSLM15, MTCR^{+}07, WKM^{+}17, DCR06, KTCR09]. Camera [DKM21, SMI06]. Can [BRS^{+}25, CR22, FKM17, RVB05, RSPA^{+}06]. Canvas [VVC^{+}22]. Capacity [TNE20]. Cartoon [FMM21]. Cartoon-Rendered [FMM21]. Case [GFD^{+}15]. Categorization [BCS17, VSWB07]. Caudell [MPC05]. Caused [HOH15, MTÆ^{+}24]. causes [KBP^{+}13]. cellphones [ZAAC12]. center [KBP^{+}13]. Centeredness [HZD23]. challenging [OAD^{+}12]. Change [KT21, PSB^{+}23, NZG^{+}11]. Changes [BFCP^{+}25, MTÆ^{+}24, NPKR23, RKP22, SDW05]. Character [CCV^{+}25, WTWN16]. Characteristics [JKB17, KHKP15]. Characterization [RRM^{+}16, MAYKM13]. characterize [GEMA13]. Characters [FMM21, HCKH16, JDKN18, HJO^{+}10, LSRS10, MJH^{+}09]. check [AG06]. Chief [IG15]. Child [JAA^{+}16]. Children [SNW16]. Chromatic [BPFP16]. chromaticity [LXXB10]. circular [RVSP09]. Clarke [Bre05]. Classification [DK19, DFJ^{+}20, TCGC19, MGM12]. Classifier [PK20]. Classifying [KWI09]. Client [SXCS15]. Client-Server [SXCS15]. Clouds [AWR18]. Clustering [GEMA13, KPL^{+}19, SMS13]. Co [CCV^{+}25, SDG25]. Co-Solving [CCV^{+}25]. Co-Speech [SDG25]. Coach [GFD^{+}15]. Code [PVK20, SF23]. Coding [HAHG17, SMI06]. Coefficient [KDK^{+}16]. Cognitive [LZL^{+}18, MBCW10, RSPA^{+}06, ZLQ^{+}19, FRC10, HMS09, KWSS08]. Collaborative [HHB^{+}24]. collected [MP09]. collection [BMGC05]. College [FGWdS^{+}25]. Collisions [O’S05, BB13]. Color [APK15, AK16, CSUN05, KHH17, KVE16, YWL^{+}24, ZLWZ24, DBS^{+}09, LPT^{+}06, VSWB07]. Colorization [ZLWZ24]. Combination [VVC^{+}22, LXXB10]. Combining [OOAY22]. Comfort [CEN^{+}23, JWA19, SWA14]. Comments [Bar05a, BM05, Bre05, BS05a, EM05a, Fer05, FTB05, SCS05, WK05b, vdD05a, GLT05a, HR05b, MPC05, Vic05]. Communication [AONB17, SXCS15]. Comparative [BMB19, DK19, HBM^{+}14, TMM17, AMR06]. Comparing [NCVW10, OAD^{+}12]. Comparison [BYB18, GATM18, GLT05b, KCK^{+}18, NPKR23, NFD^{+}21, SNW16, SB12]. Compensating [SXCS15]. Competitive [YNYK25]. Complex [LMM^{+}22, VCR08]. Complexity [FvdMV25, MP20, HMS09]. Components [CKWB05, RLV^{+}10]. Composite [TLS^{+}15]. composites [FHC04]. Composition [WKM^{+}15, WKM^{+}17, LME10]. Comprehension [BGK17]. Comprehensive [ATSD23, Bar05a]. compressed [SMI06]. computation [TGJ08]. Computational [FRC10, HZD23, KT21, WL21, ZYZ^{+}17, BGL^{+}08, PI08]. Computer [BRS^{+}25, DCRS15, FWN^{+}14, FWS^{+}23, HBF16, JOZ^{+}21, SGHL^{+}19, VCA16, SDW05, SII04, WBC^{+}07]. computer-aided-drawing [SDW05]. Computer-Assisted [DCRS15]. Computer-Generated [BRS^{+}25, FWN^{+}14, HBF16, SGHL^{+}19, JOZ^{+}21, WBC^{+}07]. concurrent [MB04]. Conditions [BYB18]. conferences [KW05]. configuration [BS05b]. Consequences [ZB17]. considerations [NCSG11]. Considering [ZYJ^{+}22]. Consistency [KNL22]. Constancy [BSH18, CWT^{+}05]. Constant [Ste15]. Consumer [VCA16]. Consumer-Grade [VCA16]. Contact [CYK^{+}21, SAB07, WH08]. Contact-type [CYK^{+}21]. Content [BFSV16, FO23, RLH^{+}08]. Content-adaptive [FO23]. Context [BDR^{+}21, KPL^{+}19, EPO11]. contingent [MDT09, WMA12]. Contour [WL21]. Contrast [MGVM16, WDLC24, YO25, MI07, MMS06]. Contribution [AWR18]. Contributions [RM16]. Control [ALN^{+}21, KHH17, APP07, BOK10, KPAA10, LPO09, MM13, MAYKM13, NVW13, SGF^{+}10, WWA11, ZZ13]. controlled [HU11, LBWP07]. Controlling [BCB20]. Conventional [MGVM16, NCVW10]. Conversational [CKWB05, EBA^{+}21]. conversations [MEDO09]. Convolutional [HBK^{+}21, PSB^{+}23, WZA^{+}23]. Cooperative [KM17]. Coordination [BB21]. Cornsweet [AR08]. correction [DCR06]. Correlates [WKM^{+}15]. correspondence [NGJT13]. Cortex [WL21]. Craik [AR08]. Creating [SJ18, ZYJ^{+}22, FHC04]. Creatures [FKM17]. Crossing [JOY^{+}18, JOZ^{+}21]. crossmodal [GMT09, RTPG11]. Crowd [BLKD24]. crowds [EPO11]. Crystallization [HR05a]. cue [FCH^{+}07, KBP^{+}13]. cue-based [KBP^{+}13]. Cueing [BJK13, RTPG11]. Cues [AWR18, CEN^{+}23, FB05, KHW^{+}15, MSHLR16, RKS16, USA20, EML13, JSHG08, LSRR13, PCK08, RVB05, RFR09]. Cunningham [SCS05]. cursor [LBT08]. curvature [AASH^{+}12]. Curve [AASH^{+}12]. Curved [ZB17]. curving [KBL^{+}06]. Cutaneous [PTP14]. Cybersickness [KLL24, LVV^{+}20]. CyberWalk [SGS^{+}11].

D [AW15, AWR18, BSW10, BSHW14, EBPJ16, GLT05b, GVC^{+}17, HOH15, HAHG17, JDR08, KPSL10, KHJK13, LPR06, LDDR18, Lav09, MSS^{+}22, NFD^{+}21, OR04, Ste15, SK16, VSKL17, WPDH14, WBNF06]. Dark [EKL^{+}21]. Data [BMGC05, BH17, FvdMV25, FBT05, NW08, TB24, Bar05b, FTB05, HDH10, LME10, MP09]. data-driven [HDH10]. Dataflow [BDW^{+}23]. datasets [HR05a]. Day [KRV^{+}14]. Day-for-Night [KRV^{+}14]. Daylighting [MMSO15]. Decisions [FKM17]. Decoding [TB24]. Decreases [JWA19]. Deep [HBK^{+}21, PSB^{+}23]. Defeat [YNYK25]. definition [HMS09]. Deformation [KFSN16, AJML13, LPHL05]. Degraded [TCGC19]. degree [ZLO13]. Delay [SXCS15]. Density [WKM^{+}15, NW08]. dependencies [VCR08]. Dependent [MSS^{+}22]. depiction [WBNF06]. Depth [AWR18, HOH15, HAHG17, JWA19, Kaw19, WRHS18, ZOH^{+}15, EML13, HHL10, KMHO13, WP10, YBC13]. Depth-based [JWA19]. Depth-Enhanced [HAHG17]. Depth-of-Field [WRHS18]. derive [BSW10]. Description [CBB^{+}14]. Descriptions [YO25]. Design [BPIC24, BCD15, BMB19, FWS^{+}23, FGWdS^{+}25, KHH17, BC05, EM05a, EM05b, GMT09, RDF11, WK05b, WMVO05]. Designed [SYF^{+}25]. Designer [FGWdS^{+}25]. designing [AMR06]. Desktop [FTB05, FBT05]. desynchronization [MEDO09]. detail [PDZ05]. Detection [EDAM^{+}24, HDF^{+}23, MGM16, NJS06, PMS17, VCA16, GDBP13, KSM^{+}05, LZG^{+}13, WMA12, WMS08]. Determine [CKWB05]. Development [HJ07]. Device [CYK^{+}21, LKTH06, SHBK05]. Devices [MSHLR16, AAM08]. diagrams [WB04]. Difference [DCRS15, UHWT21, LPT^{+}06].

Differences [RNLH16, RPH10, WDLC24, FCH^{+}07, TGT^{+}09, ZCRTW12]. Different [BYB18, JAA^{+}16, TMM17, TSC13]. differently [ZNWK12]. Difficulty [PK20]. Diffuse [TGG^{+}20]. Dimensional [FvdMV25, LZL17, HR05a, WMVO05]. Dimensions [TGG^{+}20, WM08]. direct [FM05]. direction [BSH^{+}06, MBG09, PI08]. Disc [BTDB20]. Discerning [KDK^{+}16]. Discrete [MW15]. Discrimination [CEM24, RM16, SK16, TSRD07, BSH^{+}06, HJ07, VGBF10]. Dismounted [GHS^{+}20]. Disorders [SYF^{+}25]. Disparity [KRV^{+}14, CLR10]. Displacement [CYK^{+}21, KBP^{+}13, MGM16, DFZ^{+}05]. Display [BBE22, IOYK19, LPHL05, MSS^{+}22, OOAY22, RTJ^{+}21, Bar05a, Bar05b, BS05b, BS05a, DFZ^{+}05, EML13, GDBP13, GBLR10, HJ07, KW05, SCS05, WMVO05, WCCRT09]. Display-Size [MSS^{+}22]. Displays [ATSD23, BSH18, EKL^{+}21, HOH15, HAHG17, JWA19, LLBM15, MD05, MGVM16, NMVRB20, SWA14, WPDH14, WC22, WRHS18, BM05, CLR10, HHL10, LFM12, MLK^{+}06, PCK08, SCSG05, WK05a, ZNWK12]. Disposition [BLKD24]. distal [RTPG11]. Distance [AL15, BYB18, KCK^{+}18, LWK18, MD05, PR24, PKCR05, FR08, FLKB07, GNP^{+}10, KTCR09, NAB^{+}11, NZG^{+}11, RBCK12, SCRTW05, WCCRT09, ZNWK12]. Distances [GRM^{+}21, LLBM15]. Distinctiveness [BGL^{+}08, OEMO16]. Distinguishing [SNW16]. Distortion [AW15, KBP^{+}13]. Distortions [DK19, CLR10]. distractions [MBG09]. DNN [DK19]. Do [EBA^{+}21, RDF11, SB12]. Doel [vdD05a]. Does [AAZMF21, DBR21, DFJ^{+}20, GNP^{+}10]. Dome [GATM18]. don’t [HU11]. dots [LPHL05]. Down [MGM16]. Drawing [PD17, SDW05]. Drawings [CPVC19]. driven [HDH10]. Driving [BBE16, BBE22, BDR^{+}21]. Duration [CR22]. Durations [RTJ^{+}21]. During [EBPJ16, BSHW14, CCV^{+}25, FFW07, FCH^{+}07, GTAE04, HHB^{+}24, JWB12, LAE09, MGVM16, VSCM12]. Dynamic [APLR17, BBE22, HNT^{+}22, JWA19, KFSN16, EML13, LSRS10, MMS06, NCVW10, RDLTS04].

earcon [MB04]. earcons [MB04]. Edge [EDAM^{+}24]. edges [AČMS10]. Editing [VHBO14]. Editorial [Bod21, Int06, MB10, PK07, RB04, RB08, Rus05, Tho07, BO09, CRM09, FL09, HE05]. Editors [IG15]. Editors-in-Chief [IG15]. Edwards [EM05a]. EEG [MGM12]. Effect [BRM23, CEM24, FWS^{+}23, HHNOP19, LGS25, MJP^{+}24, NP15, RO09, SXCS15, WDLC24, ZHRM15, ZNO^{+}20, AJML13, CWT^{+}05, MJH^{+}09, PJN^{+}11]. Effectiveness [KWSS08, PW10, ZCRTW12]. Effects [BBE16, BBE22, CCV^{+}25, FGWdS^{+}25, JLS^{+}17, JOZ^{+}21, KMH^{+}19, KLL24, KSLM15, LKTH06, LWK18, LMM^{+}22, MSS^{+}22, MR18, NW08, NZG^{+}11, OMN25, OOAY22, PJGE21, PR24, RTJ^{+}21, RRM^{+}16, SMI06, SWA14, EPO11, GTAE04, KTCR09, WCCRT09]. efficacy [LPO09]. Efficient [BDW^{+}23, LZG^{+}13]. Effort [KNL22]. Egocentric [LLBM15, RBCK12]. Elastic [AJML13]. elasticity [AJML13]. electro [WMA12]. electro-ocular-graph-based [WMA12]. electromyographic [NJS06]. Electrostatic [IOYK19, OOAY22]. Electrostatic-friction [OOAY22]. Electrotactile [BRM23]. Electrovibration [BRM23]. Elements [KNL22]. embodied [SBR07]. Embodiment [EBA^{+}21, KSLM15]. Embodiments [AONB17]. Emotion [GFD^{+}15, SZ22, TB24, MJM^{+}09]. Emotional [OMN25, WKM^{+}15, NTKA12]. Emotionally [WKM^{+}17]. Emotions [ZHRM15, GEMA13]. Empirical [EBPJ16, TNE20, VA05, BSW10]. employing [MRT^{+}10]. Emulating [ECOG11]. Enabling [SGS^{+}11]. encoded [WMS08]. Encoding [BDW^{+}23]. Engagement [OMN25]. Engine [PMS17]. enhance [RVSP09]. Enhanced [HAHG17, BC05, Bre05]. Enhancement [ABK^{+}15, MI07]. Enhancing [KSM^{+}05, PCK08, ZAAC12]. enough [ONS12]. Entertainment [GBTB25]. entities [SMS13]. Entropy [KDS^{+}15, ZZ13]. Environment [AL15, GATM18, JOY^{+}18, KLL24, RSM^{+}15, APP07, LSRS10, MTCR^{+}07, MRT^{+}10, RPH10, WBN^{+}11]. Environments [BSHW14, BYB18, EBPJ16, GWFL22, HHNOP19, JKB17, KPR22, LRB15, RBC14, RRM^{+}16, SXCS15, BB13, BSVDD10, FFW07, FCH^{+}07, FLKB07, GNP^{+}10, HBW11, JWB12, KBP^{+}13, KCRT08, LPO09, LBWP07, MBCW10, MC05, NAB^{+}11, NZG^{+}11, PK07, PI08, PKCR05, RBCK12, SCRTW05, SAB07, SGS^{+}11, WCCRT09, WNW^{+}07]. Equation [MVL25]. error [LPO09]. Errors [FvdMV25, YON25, RO09]. Establishing [TUG^{+}20]. Estimates [EDAM^{+}24]. Estimating [GRM^{+}21, RDLTS04]. Estimation [BYB18, CEN^{+}23, FvdMV25, FLKB07, LZL17, RNLH16, GDBP13, GNP^{+}10, LXXB10, NW08, RLH^{+}08]. Estimations [RNLH16]. EuroHaptics [HE05]. Evaluating [APK15, ALN^{+}21, AK16, BGK17, HBK^{+}21, HHO05, HCKH16, KPSL10, KYL^{+}07, LCC15, MJH^{+}09, RLV^{+}10, SD22, WBCB08, ZCRTW12, BMGC05]. Evaluation [BMB19, BM05, BTDB20, BB21, EML13, EBPJ16, JDR08, LPT^{+}06, MLK^{+}06, MR18, OMN25, RTSW18, VGBF10, VHBO14, WBC^{+}07, WBN^{+}11, YWL^{+}24, ZYZ^{+}17, AR08, BBD^{+}09, BC05, DFZ^{+}05, HJ07, HDH10, LME10, SGF^{+}10]. evaluations [NCVW10]. event [LZG^{+}13]. Events [DFJ^{+}20]. Evidence [WAH^{+}15]. EvoFIT [FHC04]. evoked [KWI09, MAYKM13]. evolutionary [FHC04]. Examining [DBR21]. Example [CSUN05, KHH17, SD22]. Example-Based [CSUN05, KHH17, SD22]. exerting [DFZ^{+}05]. expect [SB12]. Experience [RKP22, YO25]. experiences [VSCM12]. Experiencing [YNYK25]. experiment [RPH10]. Experimental [ARAP^{+}18, WK05a]. Experiments [OR04, GB08, WMVO05]. Experts [DCRS15]. exploiting [HCS10]. Exploration [PJGE21, Ste15]. exploratory [CWB10]. explore [HBW11, WBN^{+}11]. Exploring [AZ10, FGWdS^{+}25, Fer19, GWFL22, KLL24, McN06, NOSS17, PR24, ZHRM15]. Exposure [SWA14]. Expressibility [AONB17]. Expression [OMN25, SDG25]. Expression-Based [OMN25]. expressionism [SMO^{+}10]. Expressions [CKWB05, NP15, GEMA13, KWI09, WBC^{+}07, WBCB08]. Expressive [ZMM19, MDR10]. Expressiveness [DCRS15]. Extended [CEN^{+}23, EKL^{+}21, ZLO13]. Extensible [SYF^{+}25]. Extension [JKB17]. extraction [WBNF06]. Eye [DKM21, FvdMV25, LLBM15, PK20, RKS16, SF23, KVJG10, LFM12, LME10, MP09, NTKA12, TCMH11]. Eye-tracking [FvdMV25, LME10]. Eyeball [JDKN18]. eyegaze [AMR06].

Fabric [KHH17]. Face [BRS^{+}25, FWN^{+}14, FZL20, HBK^{+}21, MJP^{+}24, NP15, VCA16, ZC06, BS06, OAD^{+}12, PJN^{+}11, SB12, ZLO13]. Faces [BRS^{+}25, BGL^{+}08, TGJ08]. Facial [CXZ14, CEN^{+}23, CKWB05, FKM17, FMM21, HCKH16, NP15, OMN25, WIA24, FHC04, KWI09, NJS06, SB12, VSCM12, WBC^{+}07, WBCB08]. Facilitate [ABK^{+}15, AC11]. facilitated [RFR09]. facilitation [BGW11]. factors [RSPA^{+}06]. Fake [BRS^{+}25]. False [AK16]. familiarity [ZLO13]. fast [ONS12]. faster [LZG^{+}13]. faster-than-real-time [LZG^{+}13]. feasibility [LPHL05]. Feature [BH17, FMM21, GBA17, PK20, WKM^{+}17, ZLQ^{+}19, CKWB06]. Feature-Based [GBA17]. Features [FKM17, SWT^{+}23, ZC06]. FechDeck [Fer19]. Feedback [ARAP^{+}18, BCB20, BB21, EBPJ16, FZL20, KBL14, PTP14, YNYK25, AZ10, KL06, LKTH06, LBWP07]. feedback-controlled [LBWP07]. Fernström [Fer05]. Fidelity [APK15, BTDB20, HCS10, McN06, MRT^{+}10]. Field [AAZMF21, BB21, CKAD18, JKB17, WC22, WRHS18, ZOH^{+}15, BTDB20, LPR06, NAB^{+}11, WCCRT09]. Field-of-View [AAZMF21, JKB17]. Film [WAH^{+}15, CST^{+}10]. Films [BH17]. Filter [Fau17]. Filtering [WRHS18]. Filters [EDAM^{+}24]. finger [CEM24, TSRD07]. First [DFZ^{+}05]. Fitt [BTDB20]. Fixations [PHRE15, KVJG10, MP09, SMS13]. fixed [LPO09]. FixTag [MP09]. flow [PW10]. Flowers [FTB05]. Focal [DBR21, KDK^{+}16]. Focused [CYK^{+}21]. Foot [JSCR^{+}15, SHBK05]. Force [ARAP^{+}18, CWT^{+}05, PTP14, BSH^{+}06, DFZ^{+}05, KL06, LKTH06, VGBF10]. force-feedback [KL06, LKTH06]. Forest [MNG^{+}25]. forward [BSPB10]. foundations [FRC10]. foveal [TGT^{+}09]. Foveated [APLK17, SWT^{+}23, WRHS18]. Frame [WAH^{+}15]. Frame-Rate [WAH^{+}15]. framework [Bar05a, Bar05b, FSG09, MMS06, YBC13]. FrankenFolk [OEMO16]. Frequency [BFCP^{+}25, DRT07, LPEP12, ZC06]. Friction [IOYK19, OOAY22]. frontal [SB12]. frowning [SII04]. Full [GFD^{+}15, NP15, PMS17]. Full-Body [GFD^{+}15]. Full-Face [NP15]. Function [DBR21]. Functional [TB24, WNW^{+}07]. functions [FCH09]. fused [DCN^{+}06]. Fusion [ZLQ^{+}19].

gain [LPO09]. Gains [RKP22]. Game [ALN^{+}21, GBTB25, KDCM15, PMS17]. Games [FKM17, YNYK25, BSW10]. GAN [SWT^{+}23]. GAN-Based [SWT^{+}23]. Gap [HDF^{+}23, JSCR^{+}15]. Gaze [BSHW14, BH17, FvdMV25, KDS^{+}15, WPDH14, WMA12, BSW10, ENC^{+}08, FCH09, HU11, KPAA10, MBG09, MDT09, NVW13, PI08, WWA11]. gaze-based [ENC^{+}08]. Gaze-contingent [WMA12, MDT09]. Gaze-to-Object [BSHW14]. GazeFusion [ZWL^{+}24]. Gazing [SII04]. Gender [FCH^{+}07, ZHRM15, ZNO^{+}20]. Generate [FvdMV25]. Generated [BRS^{+}25, FWN^{+}14, HBF16, SGHL^{+}19, JOZ^{+}21, WBC^{+}07]. Generating [SDBRC13]. Generation [ZWL^{+}24]. Generative [KVDE19]. Geometries [VVC^{+}22]. Gestural [CBB^{+}14]. Gesture [BDR^{+}21, LMM^{+}22, MSHLR16, SDG25, AZ10, MM13]. Gesture-Sound [BDR^{+}21]. gestures [RM12]. get [HU11]. Glare [MGVM16]. glideslope [KPAA10]. Global [KPDP19, WL21, HFJS09, VSWB07]. Glossiness [GSW^{+}21]. Glyphs [MW15, Ste15]. Grade [VCA16]. gradient [RTPG11]. Graph [SMS13, WMA12]. Graph-based [SMS13]. Graphics [FL09, NFD^{+}21, TUG^{+}20]. graphs [NW08, WM08]. Grasping [ALN^{+}21]. Gröhn [GLT05a]. grounding [YB04]. Guest [BO09, CRM09, FL09, Int06, Rus05, Tho07]. Guidance [GWFL22, GATM18]. Guided [ZWL^{+}24, HCS10, ZLWZ24]. Guidelines [WC22, CST^{+}10].

Hair [RTSW18]. Hand [Fer19, KPR22, RVH^{+}19, SHHF^{+}22, WTWN16, AAM08, VGBF10]. hand-arm [VGBF10]. hand-held [AAM08]. Handles [SCM18]. Handling [MO09a]. Handwriting [VVC^{+}22]. Haptic [BSH^{+}06, BCD15, BB21, CYK^{+}21, EBPJ16, MTÆ^{+}24, RTJ^{+}21, SHHF^{+}22, SXCS15, AJML13, BTDB20, CWT^{+}05, CKWB06, CWB10, DKR^{+}05, GEMA13, HDH10, JSHG08, KSM^{+}05, LBT08, PDZ05, RFR09, SHBK05]. HapticWalker [SHBK05]. Harmony [MP20]. HDR [APK15, AK16, KYL^{+}07, SDBRC13]. Head [ATSD23, BBE22, CEN^{+}23, EKL^{+}21, JWA19, JKB17, LCC15, LLBM15, MD05, NMVRB20, WRHS18, JWB12, LAE09, RPH10, WCCRT09, ZNWK12]. Head-Mounted [ATSD23, EKL^{+}21, LLBM15, MD05, WRHS18, BBE22, JWA19, NMVRB20, WCCRT09, ZNWK12]. Head-Worn [JKB17]. Heading [APP07, KBL^{+}06]. Heads [CEN^{+}23, CMR^{+}05]. Headsets [DKM21]. Health [FWS^{+}23, FGWdS^{+}25]. Height [LLBM15]. held [AAM08]. Help [ASG21]. Hermann [HR05b]. Hero [KHW^{+}15]. heuristics [BSW10]. High [APLR17, HNT^{+}22, KDCM15, EML13, HCS10, HR05a, MMS06, NCVW10]. high-dimensional [HR05a]. High-Dynamic-Range [APLR17, EML13]. high-fidelity [HCS10]. High-Level [KDCM15]. Higher [WAH^{+}15]. Highlighting [KM17]. Hint [ZLWZ24]. Hint-guided [ZLWZ24]. History [DBR21]. HMD [BYB18, KTCR09, LRB15]. HMD-Based [BYB18, LRB15]. HMDs [LWK18]. holes [LBT08, VVHV10]. holistic [FHC04]. Horizon [MD05]. HTC [KCS17, KKC19]. HTTP [BMB19]. HTTP-based [BMB19]. Human [ASG^{+}18, BLKD24, BDW^{+}23, DK19, EDAM^{+}24, FWN^{+}14, FGWdS^{+}25, HBK^{+}21, KMH^{+}19, KM17, KVDE19, LDDR18, NOSS17, SK18, SNW16, SZ22, TVR^{+}11, UHWT21, WZA^{+}23, DKR^{+}05, ECOG11, HJO^{+}10, JSG09, KVJG10, LZG^{+}13, SII04, TJL^{+}11, VGBF10]. Human-inspired [TVR^{+}11]. Human-like [WZA^{+}23]. Human-Machine [KM17]. Human-Perceived [LDDR18]. Human-Robot [SK18]. Humans [BRS^{+}25, FWS^{+}23, MVL25, PSB^{+}23, ZHRM15, ZMM19, OAD^{+}12, SB12]. Hybrid [MDT09]. Hypergravity [PUC^{+}24]. Hypothesis [CXZ14].

ICAD [Bar05a, BM05, Bre05, BS05a, EM05a, Fer05, FTB05, GLT05a, HR05b, MPC05, SCS05, Vic05, WK05b, vdD05a]. Icon [CR22]. Icons [SJ18]. Ideal [CEN^{+}23]. Identification [BFSV16, HBF16, PSB^{+}23, RTJ^{+}21, HJ07, NW08, TSRD07]. Identifying [BOK10, SF23, TGJ08, MP09]. Identity [HBK^{+}21]. Identity-Masking [HBK^{+}21]. II [LKTH06]. illumination [DCR06, HFJS09, LXXB10, YCK^{+}09]. Illusion [Kaw19, MTÆ^{+}24, RVH^{+}19, SHHF^{+}22, AR08, RVSP09]. Illustrations [SGHL^{+}19]. Image [AW15, BPFP16, FB05, FO23, NG06, PMS17, RNLH16, SWT^{+}23, ZWL^{+}24, LAE09, MDT09, RLH^{+}08, RLV^{+}10, SDBRC13, SLW^{+}11, TGT^{+}09, WMS08]. image-processing [RLV^{+}10]. Image-Quality [RNLH16]. image/model [MDT09]. image/model-based [MDT09]. Imagery [WDLC24, McN06, ONS12]. Images [ABK^{+}15, BRS^{+}25, CSUN05, DK19, FWN^{+}14, GBA17, Kaw19, MMSO15, TCGC19, WBHP20, AJML13, DCN^{+}06, MI07, MMS06, MO09a, NCVW10, SDBRC13, SMI06, WP10]. Imaging [APLR17, FHC04]. Immersive [CCV^{+}25, EBPJ16, GWFL22, GATM18, HNT^{+}22, JOY^{+}18, KSLM15, JWB12, KS12, KCRT08, LSRS10, MBCW10, MC05]. Impact [BFCP^{+}25, KHKP15, MNG^{+}25, MP20, NMVRB20, WTWN16, YON25, YNYK25, MM13]. Impaired [RL17, DKR^{+}05]. Impairments [BPFP16]. Impede [AAZMF21]. Implications [GMT09]. implicit [HMS09]. importance [BAMB13]. Important [SWT^{+}23, ZMM19]. Impressionism [SMO^{+}10]. Impressions [MSS^{+}22]. Improving [DKR^{+}05, HBF16, KCK^{+}18, KM17, KMHO13, MHPM23, PTP14]. Inconsistencies [TLS^{+}15]. Inconsistency [KMH^{+}19]. Index [LZL^{+}18]. indicators [SCRTW05, ZCRTW12]. indirect [YCK^{+}09]. Individual [DCRS15, RNLH16, ZCRTW12]. Indoor [RL17]. Induced [LZL^{+}18, SHHF^{+}22, RVSP09]. Industry [ASG^{+}18]. Inertial [SK18]. Influence [CKAD18, EBA^{+}21, FKM17, FvdMV25, LZL17, MJP^{+}24, RSM^{+}15, SS19, VSKL17, BSH^{+}06, NTKA12, RSPA^{+}06, YCK^{+}09]. Influences [HCKH16, MGM16, BB13]. Information [GHS^{+}20, KT21, YON25, BSH^{+}06, FSG09, KPSL10, VCR08, VSWB07, YBC13]. information-theoretic [FSG09]. infrared [KWI09]. Initial [KVE16]. Ink [ZLWZ24]. Inking [VVC^{+}22]. Input [VVC^{+}22]. Insights [AMR06]. inspired [TVR^{+}11]. Integrating [AC11]. integration [FCH^{+}07]. Intelligence [CCV^{+}25]. Intelligent [WKM^{+}17]. Intensity [CYK^{+}21, MTÆ^{+}24, NP15]. Intentions [FGWdS^{+}25, KHW^{+}15, MVL25, MDR10]. Inter [RTJ^{+}21]. Inter-Phoneme [RTJ^{+}21]. Inter-Word [RTJ^{+}21]. Interaction [DKM21, EBPJ16, MSHLR16, VSKL17, SII04]. Interactions [KBL14]. Interactive [FZL20, HHNOP19, KHH17, Ste15, AASH^{+}12, PI08, TSC13, WB04, ZZ13]. Interface [BDR^{+}21, PD17, AAM08, LBWP07, YB04]. Interfaces [EKL^{+}21, VCA16, AZ10, GMT09]. interference [HBW11]. Interferences [PD17]. Interlacing [HOH15]. Interleaved [HAHG17]. internal [RDLTS04]. international [KW05]. Interocular [WDLC24]. interpolation [SLW^{+}11]. Interpretation [RLH^{+}08]. Interpretation-based [RLH^{+}08]. Interpreting [SF23]. Interprets [CR22]. Interval [RTJ^{+}21]. Interventions [ABK^{+}15]. Intrinsic [SZ22]. Introduction [BT16, CR23, DB25, GS13, GGFE20, HC19, JO21, KMS15, LP24, MS12, RK18, SBC22]. Intuitiveness [LSRR13]. Invasive [LZL^{+}18]. inverse [WH08]. Investigating [AONB17, HHL10, MJM^{+}09, VVC^{+}22, WKM^{+}15, WIA24, ENC^{+}08]. investigation [HFJS09]. isotropic [HVM06]. Issue [BT16, CR23, DB25, GGFE20, HC19, JO21, KMS15, LP24, RK18, SBC22, FL09, GS13, HE05, MB10, MS12]. ITMOs [APK15].

Jigsaw [CCV^{+}25]. Jitter [BFCP^{+}25, JDKN18]. Joint [JOY^{+}18, SK18, SMS13, TSRD07]. joint-angle [TSRD07]. Jointly [ZYJ^{+}22]. Joystick [NPKR23]. Joystick-Based [NPKR23]. JPEG [SMI06]. Judging [DCRS15]. Judgments [LWK18, LRB15, PR24, KTCR09, KMHO13, SAB07, WCCRT09, ZNWK12]. Jumps [SF23].

Keep [JWA19]. Kinematic [YON25]. kinematics [WMVO05]. Kramer [WK05b].

Lamps [KFSN16]. language [TVR^{+}11, YB04]. Latency [APLK17, BFCP^{+}25, BBE16]. lateral [DFZ^{+}05, LPHL05]. Laughter [NP15]. Law [BTDB20]. LDR [SDBRC13]. Lead [AAM08]. Lead-me [AAM08]. Learning [AAZMF21, LMM^{+}22, PK20, PVK20, RVB13, SZ22, SWT^{+}23, TB24, DKR^{+}05, GBLR10, KS12, KWSS08, YB04, ZAAC12]. Learning-based [SZ22]. learns [KS12]. Ledge [LRB15]. less [KBP^{+}13]. Level [FB05, KDCM15, PDZ05]. level-of-detail [PDZ05]. Leveraging [KT21]. Light [KPDP19, LWK18, VSKL17]. Light-Material [VSKL17]. Lighting [TLS^{+}15, BAMB13, NCVW10, TGJ08]. Lightness [APLR17]. Like [OMN25, WZA^{+}23]. limb [BOK10]. limits [HS12, LPR06]. Line [TUG^{+}20]. Linear [KHJK13, MW15, SB12]. Linearities [TCP^{+}14]. Lines [SF23, TG19]. Linguistics [PK20]. link [WB04]. liquid [vdD05b]. Listener [CR22]. listeners [RPH10]. Listening [PD17]. LMA [KNL22]. Load [LZL^{+}18, ZLQ^{+}19, HMS09]. Local [VSWB07, AJML13, Lav09]. Localization [GBTB25, WPDH14]. Location [RL17]. Locomotion [ASG^{+}18, KLL24, NPKR23, KCRT08, LBWP07, MTCR^{+}07]. locomotor [WWA11]. LOD [DBS^{+}09]. Logos [ZYZ^{+}17]. Lokki [GLT05a]. Long [SWA14]. Long-Term [SWA14]. Longitudinal [RKP22]. look [HU11]. Looking [PHRE15]. Low [CYK^{+}21, FB05]. Low-Level [FB05]. LSB [WMS08]. LSB-encoded [WMS08]. lumigraphs [MO09b]. Luminance [YO25].

Machine [KM17, SNW16, SZ22, TB24]. Magnitude [HCKH16, VGBF10]. Make [KFSN16]. Making [SGF^{+}10]. Malformation [ABK^{+}15]. management [DBS^{+}09]. manipulate [KS12]. Manipulated [JLS^{+}17]. Manipulating [CKWB05]. Manipulation [FMM21, HHB^{+}24, SK18, FM05, NVW13]. Manipulations [BDR^{+}21, LLBM15]. Manual [USA20]. Manufacturing [ASG^{+}18]. map [ZAAC12]. Mapped [GBA17]. Mapping [BSHW14, GWFL22, AG06, GB08]. Mappings [WK05a]. Maps [AK16]. Marking [KW05]. Masking [HBK^{+}21, Lav09]. Mass [SK18]. Material [BCS17, VSKL17, BSVDD10, HJ07]. Materials [FK19, FB05, DFJ^{+}20]. mathematical [KHJK13]. Matters [RKP22]. maximized [LZG^{+}13]. McGurk [CMR^{+}05]. McNamara [Fer05]. Me [ASG21, AAM08]. Mean [BFCP^{+}25]. Meaning [CR22]. Means [PTP14]. Measure [LDDR18, HMS09, Lav09]. Measurement [LZL^{+}18, GNP^{+}10, KBL^{+}06, KWI09]. Measurements [ÆÁP^{+}22]. Measures [TNE20, McN06]. Measuring [ZLO13]. mechanical [WCCRT09]. mechanically [VVHV10]. mechanics [WH08]. Mechanism [ZLWZ24]. Mechanisms [WL21]. Media [OMN25, SGA^{+}07]. mediation [KWSS08]. Memory [JCM^{+}25, MRT^{+}10]. Mental [BPIC24, FWS^{+}23, FGWdS^{+}25]. Mesh [KVJG10, FSG09]. Meshes [GVC^{+}17, Lav09]. Message [IG15]. metaphors [WK05a, WK05b]. Method [KLL24, MMSO15, USA20, BB13, GNP^{+}10]. Methodology [AK16, EM05b]. Methods [CMR^{+}05, DCN^{+}06, GATM18, HBK^{+}21, KCK^{+}18, NFD^{+}21, JDR08]. Metric [GBA17, RKC^{+}22]. Metrics [PMS17, YO25]. Mid [HDF^{+}23, MHPM23, SHHF^{+}22]. Mid-air [HDF^{+}23, MHPM23, SHHF^{+}22]. Mimebot [AONB17]. Miner [MPC05]. Minification [ZNWK12]. mirrors [AC11]. Mismatch [BTDB20, BB21]. Mitsopoulos [EM05a]. Mixed [IOYK19]. Mobile [DKM21, KBL14, MSHLR16, ONS12, WWA11]. Modalities [PD17, TMM17, VVC^{+}22]. Mode [EKL^{+}21, NP15]. Model [ATSD23, FvdMV25, HR05b, KHKP15, SJ18, WL21, ZLQ^{+}19, HVM06, KHJK13, LPT^{+}06, SB12, SVHS06]. Model-Based [SJ18, HR05b, MDT09]. Modeling [APLR17, BDW^{+}23, KPL^{+}19, MVL25, SBR07, SZ22, TJL^{+}11, TCMH11, WWA11, YWL^{+}24, PW10]. Models [APK15, CMR^{+}05, HHO05, KVDE19, NVW13, VSKL17, KBP^{+}13, RDLTS04, vdD05b].

Moderate [OOAY22]. Modes [ALN^{+}21]. modification [LBT08]. modulated [MHPM23]. mono [WP10]. monocular [EML13]. Moral [FKM17]. Morphing [KT21, MJP^{+}24]. Morse [PVK20]. Motion [AWR18, HCKH16, JAA^{+}16, JLS^{+}17, KHW^{+}15, MSHLR16, NOSS17, OEMO16, VHBO14, WTWN16, WB04, ZHRM15, ZNO^{+}20, BSPB10, BOK10, CLR12, DRT07, FLKB07, JWB12, LPR06, LAE09, MJH^{+}09, MAYKM13, MTCR^{+}07, NCSG11, NGJT13, RSPA^{+}06, RFR09, RVSP09, TSRD07]. motion-field [LPR06]. motivated [CST^{+}10, SLW^{+}11]. Motivation [YNYK25]. Motor [BB21, LMM^{+}22]. Mounted [ATSD23, EKL^{+}21, LLBM15, MD05, WRHS18, BBE22, JWA19, NMVRB20, WCCRT09, ZNWK12]. Mouse [BFSV16, KL06]. MovEcho [BDR^{+}21]. Movement [GFD^{+}15, KVDE19, LMM^{+}22, HCS10, LAE09]. Movements [BFSV16, RKS16, NTKA12]. Moving [RVSP09, RBCK12]. MR [ABK^{+}15]. Multi [ZLQ^{+}19]. Multi-Feature [ZLQ^{+}19]. Multidimensional [CWB10]. multimedia [BJK13]. Multimodal [BWG12, LCC15, RM16, TG19, TUG^{+}20, GMT09, YB04]. multiple [AC11, LSRS10]. multiscale [LPT^{+}06]. multivariate [RLV^{+}10]. multiview [HHL10]. muscles [NJS06]. Music [DCRS15, MP20, SZ22, WKM^{+}15, WKM^{+}17, AZ10, MDR10]. Musical [LMM^{+}22, MGM16, PD17, VA05, WKM^{+}17]. My [ASG21].

Natural [RM16, SVHS06, TGT^{+}09,

VSWB07, WMS08, WP10]. natural-image [TGT^{+}09]. Naturalness [EBA^{+}21, KVDE19, MVL25]. Nature [MNG^{+}25]. Navigating [KLL24]. Navigation [NPKR23, USA20, GBLR10, GLT05b, MLK^{+}06, VVJD05]. Near [BTDB20, BB21, NAB^{+}11, KWSS08]. Near-Field [BB21, BTDB20, NAB^{+}11].

Negative [LPO09, KWI09]. Network [HBK^{+}21, PSB^{+}23, TCGC19]. Networks [WZA^{+}23, NJS06]. Neural [HBK^{+}21, PSB^{+}23, PW10, TCGC19, WZA^{+}23, NJS06]. Night [KRV^{+}14]. Nighttime [MGVM16]. NMF [ZLO13]. node [WB04]. Non [AONB17, FKM17, LZL^{+}18, TG19]. Non-Invasive [LZL^{+}18]. Non-player [FKM17]. Non-Verbal [AONB17]. Non-visual [TG19]. nonvisual [EM05b, JSHG08]. normal [DKR^{+}05]. North [FWS^{+}23]. novel [AJML13, DFZ^{+}05, SHBK05, WMVO05]. Numerical [UHWT21]. Numerically [RDLTS04]. numerosity [GDBP13].

Object [BSHW14, CKWB06, HU11, SK18]. Objective [GVC^{+}17]. Objects [KFSN16, LDDR18, CA13, NGJT13, RDLTS04]. O’Brien [AR08]. observers [ECOG11]. Obstacle [FFW07]. occluders [MO09a]. ocular [WMA12]. Oculomotor [KHKP15, KHJK13]. off [LRB15]. Offsets [KPR22]. Olfactory [NMVRB20, RBC14]. omnidirectional [SGS^{+}11]. omnistereoscopic [CLR10]. On-Air [VVC^{+}22]. Online [WPDH14, YNYK25]. Onset [BCB20]. Open [SYF^{+}25]. operator [GB08]. operators [AR08, AG06]. Optical [EKL^{+}21]. optically [VVHV10]. Optimal [ONS12]. Optimized [FO23]. Optimizing [BS05b, BS05a, WC22]. Order [MTÆ^{+}24]. organization [MDR10]. orientation [RPH10, ZCRTW12]. Other [MJP^{+}24, PJN^{+}11]. Other-Race [MJP^{+}24, PJN^{+}11]. Outdoor [TLS^{+}15]. Overlapping [PR24].

pace [TSC13]. painters [SMO^{+}10]. Painting [ZLWZ24, ZZ13]. Paintings [ZYJ^{+}22]. Pairs [HDF^{+}23, SDBRC13]. Palaces [JCM^{+}25]. palette [BC05]. panoramas [CLR12]. panoramic [MO09a]. Parallax [BB21]. Parallel [ÆÁP^{+}22].

Parameter [Fau17]. Parameters [KHKP15]. Parametric [LPEP12]. part [FBT05]. participating [SGA^{+}07]. Participatory [FWS^{+}23]. Partners [HHB^{+}24]. Passability [GSCRB23]. Password [ASG21]. path [FCH^{+}07, HHL10, KBL^{+}06]. path-searching [HHL10]. patients [APP07]. Pedestrian [JOY^{+}18, EPO11, SAB07]. Pedestrians [RL17]. people [KS12]. Perceivably [JAA^{+}16]. Perceived [BLKD24, BMB19, CKAD18, EBA^{+}21, KBL14, KCS17, KVDE19, KSLM15, LDDR18, MVL25, RKC^{+}22, SMI06, WKM^{+}15, YO25, ZOH^{+}15, BB13, DRT07, KBL^{+}06, KL06, LKTH06, MJH^{+}09, WP10]. Perceiving [AJML13]. Perception [ARAP^{+}18, AWR18, BRM23, BFSV16, BB21, CBB^{+}14, CXZ14, CPVC19, CLR12, FWN^{+}14, FZL20, FB05, FL09, GSCRB23, GSW^{+}21, GFD^{+}15, HBK^{+}21, JOZ^{+}21, KPDP19, KCK^{+}18, LAE09, MI07, MD05, MHPM23, NP15, PR24, RM12, SK18, SNW16, SLW^{+}11, TLS^{+}15, TMM17, TG19, TGT^{+}09, VVHV10, WZA^{+}23, WIA24, YON25, ZHRM15, ZMM19, AASH^{+}12, BSVDD10, Can09, CA13, CWT^{+}05, ENC^{+}08, FR08, KBP^{+}13, MJM^{+}09, MMS13, NAB^{+}11, NZG^{+}11, OR04, PKCR05, RBCK12, RSPA^{+}06, RDF11, SCRTW05, TJL^{+}11, YBC13, BRS^{+}25]. Perception-Action [BB21, JOZ^{+}21]. Perception-Based [BFSV16, KPDP19, MI07]. Perception-motivated [SLW^{+}11]. Perceptions [HCKH16, YB04]. perceptive [BBD^{+}09]. Perceptual [APLR17, AR08, AW15, BSH18, BAMB13, CEN^{+}23, EM05a, EPO11, FK19, HBM^{+}14, JDKN18, KMH^{+}19, KNL22, LPR06, MM13, MAYKM13, MDR10, MR18, NCSG11, SCSG05, SGF^{+}10, SGA^{+}07, TGG^{+}20, VCR08, VHBO14, WC22, YCK^{+}09, ZB17, Bar05b, BMGC05, JSG09, MMS06, McN06, MO09a, MO09b, NG06, PCK08, RO09, WBCB08, WH08, WMVO05, ZZ13]. Perceptually [CST^{+}10, CMR^{+}05, Fau17, FO23, HCS10, KFSN16, LCC15, SWT^{+}23, HVM06]. Perfect [KBL14]. Performance [BFCP^{+}25, BTDB20, DBR21, DK19, DKR^{+}05, GTAE04, HHL10, KBL^{+}06, MBG09, MDR10, RLV^{+}10]. Performances [DCRS15]. Peripheral [BDW^{+}23, LWK18, LVV^{+}20, DBS^{+}09, TGT^{+}09]. Periphery [BCS17]. Persistent [JOZ^{+}21]. Personal [WAEG06]. Personality [GFD^{+}15, HCKH16, SDG25, WTWN16]. Perspective [NOSS17]. phenomena [BAMB13]. Phoneme [RTJ^{+}21]. Phonemic [RTJ^{+}21]. Phonemic-Based [RTJ^{+}21]. Photo [FWN^{+}14, KPL^{+}19]. Photoplethysmogram [ZLQ^{+}19]. Photoplethysmogram-based [ZLQ^{+}19]. Photorealism [ZMM19]. Phrases [RTJ^{+}21]. Physical [EBPJ16]. Physicality [YON25]. Physically [vdD05b]. physics [vdD05a]. physiological [VSCM12]. Picking [KVE16]. pictorial [WP10]. Pilot [KPAA10]. Ping [DFJ^{+}20]. pipe [FR08]. pipeline [BSW10]. place [WBN^{+}11]. places [TKK^{+}13]. planar [MMS13, RM12]. Planning [ABK^{+}15]. Plant [KHKP15, KHJK13]. plasticity [SCSG05]. platform [BSPB10]. plausible [SVHS06]. Player [YNYK25, FKM17]. Point [AWR18, NW08]. point-estimation [NW08]. Pointers [MHPM23]. Pointing [WP10]. polar [ZC06]. Polarized [HAHG17]. Pole [LRB15]. Polygonal [HHO05]. Polyhedral [CPVC19]. Pong [DFJ^{+}20]. poorer [TGT^{+}09]. populations [DKR^{+}05]. portable [MP09]. Portraits [HBF16]. Portrayal [FMM21]. Posed [TCP^{+}14]. Position [KLL24, KS12, TSRD07]. positive [KWI09]. possibility [RFR09]. Possible [LLBM15]. potential [HBW11]. Practical [ATSD23]. Precedence [WL21]. Predict [KHW^{+}15, PI08]. predicted [KBP^{+}13]. Predicting [HHO05, KVDE19, PK20]. Prediction [ATSD23, SZ22, ZLWZ24, BSW10, LPT^{+}06]. predictions [RDF11]. Predictive [BCD15]. Predictor [FO23, KDCM15]. Prefer [WAH^{+}15]. Preference [PJGE21, TKK^{+}13, YWL^{+}24, FCH^{+}07]. preferences [GTAE04]. Presence [KLL24, LVV^{+}20, MBG09]. Presentation [AL15, BCB20, NP15, GNP^{+}10, WMA12]. Presents [IOYK19]. Press [KBL14]. pretended [KWI09]. Primary [WL21]. Primitive [KVDE19]. principled [EM05b]. principles [MB04, YBC13]. Printed [Kaw19]. private [BCB20]. problems [WH08]. procedures [CWB10]. processing [MMS06, RLV^{+}10]. Product [SS19]. profile [SB12]. Profiles [TUG^{+}20]. Program [BGK17, Vic05, VA05]. Projection [KFSN16, KBP^{+}13]. Promote [FGWdS^{+}25]. Propagation [RRM^{+}16]. Properties [HNT^{+}22, BSVDD10, WCCRT09]. Proposed [AK16]. proprioception [BOK10]. Prosody [EBA^{+}21]. prosthesis [BOK10]. protocol [GNP^{+}10]. Providing [BCB20]. Proxemics [LSRS10, KS12]. proximal [RTPG11]. proximal-distal [RTPG11]. Proximity [ZNO^{+}20]. pseudo [AJML13, LBT08]. pseudo-haptic [AJML13, LBT08]. Psychoacoustic [RRM^{+}16]. psychophysical [HFJS09]. psychophysically [SVHS06]. Psychophysics [Fer19, TGJ08]. PTRM [RKC^{+}22]. Public [BCB20]. Pulfrich [LGS25]. pulling [AAM08]. Pupil [JDKN18]. purposes [MAYKM13]. Puzzle [CCV^{+}25].

Qualities [GFD^{+}15, BMGC05]. Quality [BPFP16, BMB19, DK19, GBA17, GVC^{+}17, KBL14, MNG^{+}25, NFD^{+}21, PMS17, RNLH16, SGHL^{+}19, VSKL17, NCVW10, RLH^{+}08, RLV^{+}10, SMI06]. Quantification [CYK^{+}21]. Quantifying [MRT^{+}10, SGHL^{+}19, WBHP20, WP10]. Quantity [MW15]. queries [WB04].

Race [MJP^{+}24, PJN^{+}11]. radiologist [AMR06]. Range [APLR17, HNT^{+}22, EML13, MMS06, NCVW10]. ranking [SVHS06]. rapid [WB04]. Rate [WAH^{+}15]. ratings [CKWB06]. Ratio [IOYK19]. Reaches [RSM^{+}15]. Reaching [EBPJ16]. Reactions [BLKD24]. Reading [SF23, BWG12, GTAE04]. reading-related [GTAE04]. Real [ASG^{+}18, BRS^{+}25, Can09, EBPJ16, FK19, LFM12, FFW07, GNP^{+}10, LZG^{+}13, McN06, NAB^{+}11, NCVW10, ONS12, PK07, PKCR05, SCRTW05, SGF^{+}10, WBC^{+}07, WNW^{+}07]. Real-time [LFM12]. Real-world [Can09, FK19, WBC^{+}07]. Realism [FWN^{+}14, MNG^{+}25, MVL25, OOAY22, RKC^{+}22, SD22, BB13, ENC^{+}08, WBCB08]. Realistic [CMR^{+}05, FMM21, VHBO14]. Reality [ASG^{+}18, APLK17, AL15, BPIC24, BTDB20, CEN^{+}23, GRM^{+}21, GSCRB23, GBTB25, JLS^{+}17, KCK^{+}18, KKC19, KSLM15, LVV^{+}20, LGS25, MNG^{+}25, NFD^{+}21, PUC^{+}24, RKP22, RVH^{+}19, RM16, SHHF^{+}22, SYF^{+}25, VVC^{+}22, WC22, WDLC24, ZMM19, ZNO^{+}20, APP07, AG06, HFJS09, KS12, KWSS08, KMHO13, MBCW10, ONS12, RSPA^{+}06, RVSP09, RVB13, SDW05, WAEG06]. Reality-Based [SYF^{+}25]. Recalibration [KCRT08]. Recall [JCM^{+}25]. Recognition [RKS16, TB24, VCA16, BS06, BWG12, JSHG08, MB04, OAD^{+}12, ONS12, PJN^{+}11, SMO^{+}10, TVR^{+}11, ZC06]. Recognizing [HHB^{+}24]. Reconstructing [BFSV16]. Reconstruction [SWT^{+}23]. Recover [SWT^{+}23]. Recovery [ARAP^{+}18]. Redirected [HBW11]. Reduce [AAZMF21, LLBM15]. reducing [LFM12, LPO09]. Reduction [MGVM16]. redundancy [TVR^{+}11]. Reference [PMS17]. Reflected [LZL^{+}18]. Reflection [TGG^{+}20]. Reflections [Fer05]. reflexlike [RVB05]. Region [BS06]. Region-based [BS06]. Regularities [NOSS17]. regulate [VSCM12]. regulation [LSRR13]. Reinforcement [KS12]. Related [KPR22, GTAE04]. Relating [YO25]. Relationship [SZ22]. Relative [KLL24, LZL17, KMHO13, WP10]. Remember [ASG21]. Remote [RL17]. Rendered [FMM21]. Rendering [APLK17, JWA19, RTSW18, BGW11, BAMB13, BBD^{+}09, HCS10, HDH10, JSG09, KYL^{+}07, MDT09, NCSG11, PDZ05, PW10, SGA^{+}07, KL06]. replication [RM12]. representation [LPEP12]. representations [BS06, WNW^{+}07]. Representing [MW15]. Reproducing [RM16]. Reproduction [BSH18, AR08]. Requirements [APLK17, GHS^{+}20]. resampling [FCH09]. resistive [KL06, LKTH06]. resolution [GMT09, HS12]. Resonance [CEM24]. resonators [OR04]. Response [BB13, LZL^{+}18]. Responses [OOAY22, NAB^{+}11]. Restorative [MNG^{+}25]. Restriction [AAZMF21]. resulting [MTCR^{+}07]. Retargeting [KRV^{+}14]. Retrieval [FZL20, NG06]. reversals [NW08]. Review [TNE20]. revisited [HR05b]. Rhythmic [WKM^{+}15]. Ritter [HR05b]. Road [JOY^{+}18, GB08]. Roads [JOZ^{+}21]. Robot [SK18]. Role [CBB^{+}14, DCRS15, GSW^{+}21, BOK10, MJM^{+}09, VSWB07]. Rotation [BPIC24, JDKN18]. Rotational [RKP22, USA20, KCRT08]. Roughness [IOYK19, KL06, Lav09, LKTH06]. Row [HAHG17]. Row-Interleaved [HAHG17]. Rubber [SHHF^{+}22]. Rules [RNLH16].

Saccade [ATSD23, WMA12]. Saccadic [RKS16]. Salience [SS19, NTKA12]. Saliency [HZD23, HHO05, KDCM15, LZG^{+}13, ZWL^{+}24, FSG09, HJO^{+}10, KVJG10]. Saliency-Guided [ZWL^{+}24]. Saliency-maximized [LZG^{+}13]. Salient [HZD23]. Salient-Centeredness [HZD23]. sample [GDBP13]. Sampling [BCD15, HAHG17, ONS12]. SAP [BT16, CR23, DB25, GS13, GGFE20, HC19, JO21, KMS15, LP24, MS12, RK18, SBC22]. Scale [JSCR^{+}15, NZG^{+}11]. Scales [CEN^{+}23]. scaling [CWB10]. Scattering [GSW^{+}21]. Scatterplots [Ste15]. scenario [BGW11, RO09]. Scene [JWB12, MGM16, RM16, TCGC19, AC11, EPO11, MB04]. Scene-motion [JWB12]. Scenes [TLS^{+}15, ENC^{+}08, NCVW10, SVHS06, VSWB07, WP10]. Scents [ASG21]. schema [MRT^{+}10]. schools [SMO^{+}10]. Science [FWS^{+}23, KW05]. scientific [Bar05b]. Scotopic [KRV^{+}14]. Screen [BCB20, BB21, CKAD18, JSG09]. Screen-space [JSG09]. Screens [ZB17]. Search [BSHW14, KM17, MBG09, AMR06, BJK13, TVR^{+}11, VCR08]. Searches [SS19]. searching [HHL10]. See [DBR21, EKL^{+}21]. See-Through [EKL^{+}21]. Seeing [PD17, SB12]. seen [WH08]. Segmentation [ECOG11]. Selection [KPL^{+}19, FSG09, HU11]. Selective [Can09]. Self [JLS^{+}17, OMN25, PR24, DRT07, RSPA^{+}06, RFR09, RVSP09]. Self-Awareness [OMN25]. Self-Motion [JLS^{+}17, DRT07, RSPA^{+}06, RFR09, RVSP09]. Self-Overlapping [PR24]. Selfie [DKM21]. Semantic [LZL17]. Semantics [ZYJ^{+}22]. Semi [BCB20]. Semi-private [BCB20]. Sensation [CYK^{+}21, AAM08]. Sensing [RL17]. Sensitivity [KPR22, MEDO09, RKP22, SWA14, DBS^{+}09, RO09]. sensorial [MDR10]. Sensorimotor [PD17]. sensors [BWG12]. Sensory [ARAP^{+}18, BTDB20, BB21, YB04]. Sensory-Motor [BB21]. separation [SMI06]. Sequences [CKWB05, SLW^{+}11]. Server [SXCS15]. Sets [WKM^{+}17]. seven [BS05b]. seven-talker [BS05b]. sex [MJH^{+}09]. Shadow [HBM^{+}14, Kaw19]. Shadow-based [Kaw19]. Shape [CA13, DBR21, AASH^{+}12, MJH^{+}09, MJM^{+}09, WBNF06]. Shapes [CPVC19, KPDP19, MHPM23, JDR08]. Shared [BCB20, SXCS15, LPO09]. shear [DFZ^{+}05]. Shine [RTSW18]. Shinn [SCS05]. Shinn-Cunningham [SCS05]. shiny [WBNF06]. Should [FBT05]. Sickness [AAZMF21]. sight [MLK^{+}06]. Sighted [RL17, RSM^{+}15]. Sign [LZL17, TVR^{+}11]. Signals [BCD15, NJS06]. significant [AČMS10]. similar [SCRTW05]. similarities [WNW^{+}07]. Similarity [FWS^{+}23, FGWdS^{+}25, CKWB06, NG06]. Similarity-Attraction [FWS^{+}23]. Simple [JWA19]. Simplified [HHO05]. simplify [MP09]. Simpson [BS05a]. Simulated [FvdMV25, PUC^{+}24, ENC^{+}08, McN06, VVHV10]. Simulating [BSPB10, MGVM16, PHRE15, LBT08, TSC13]. simulation [RFR09]. Simulations [UHWT21, MRT^{+}10]. Simulator [BBE22]. Simultaneity [KBL14]. Simultaneous [KBL^{+}06]. Single [MGM12]. Single-trial [MGM12]. Situations [MGVM16]. Size [CKAD18, HZD23, JDKN18, JSCR^{+}15, LVV^{+}20, MSS^{+}22, LBT08, OR04]. Sketched [CPVC19]. Sketching [WBNF06]. skin [JSG09, KWI09, LPHL05]. Slant [WZA^{+}23]. slip [WMVO05]. Smart [KDCM15]. Smile [OMN25]. Smiles [TCP^{+}14]. Sneaking [KHW^{+}15]. social [VSCM12]. Soft [HBM^{+}14, KM17]. Soft-Shadow [HBM^{+}14]. Softness [LDDR18]. Software [SYF^{+}25, FBT05]. Soldier [GHS^{+}20]. Solid [MW15, JDR08]. Solution [LLBM15]. Solving [CCV^{+}25]. sonic [Fer05, RM12]. Sonically [Bre05, BC05]. Sonically-enhanced [Bre05]. Sonification [GWFL22, LMM^{+}22, WK05b, AASH^{+}12, FM05, FTB05, FBT05, HR05a, HR05b]. Sound [BGK17, BDR^{+}21, CBB^{+}14, GDBP13, GBTB25, KW05, MR18, BGW11, FBT05, LPEP12, NGJT13, vdD05a]. sounds [MAYKM13, MC05, RVSP09, TSC13, vdD05b]. Source [CBB^{+}14, SYF^{+}25]. Space [Fau17, GRM^{+}21, KBP^{+}13, KCS17, KSLM15, YO25, GLT05b, JSG09, KWSS08, WP10, WAEG06]. Spaces [PR24]. Spatial [AAZMF21, GWFL22, GBLR10, HOH15, MGM16, SCS05, TCP^{+}14, BS05b, GMT09, HBW11, LPEP12, MBCW10, MLK^{+}06, RVB05, RTPG11, SCSG05, WNW^{+}07]. spatialization [MM13]. spatialized [BGW11]. Spatio [MHPM23]. Spatio-temporally-modulated [MHPM23]. spatiochromatic [DBS^{+}09]. spatiotemporal [KPSL10]. speakers [RPH10]. Special [BT16, CR23, DB25, FL09, GGFE20, HC19, JO21, KMS15, LP24, RK18, SBC22, GS13, HE05, MB10, MS12]. specification [EM05b]. spectral [HVM06]. Specular [TGG^{+}20, WBNF06]. Speech [EBA^{+}21, SDG25, BS05b, BS05a]. speed [LBT08, LSRR13]. spoken [YB04]. Spontaneous [TCP^{+}14]. Sport [GFD^{+}15]. Spot [UHWT21]. Stabilization [KKC19]. standard [FBT05]. standing [APP07]. states [KWI09, MBCW10]. Static [BGK17, KFSN16, LVV^{+}20]. Statistics [BCS17]. Stealth [KT21]. steering [KBL^{+}06]. steganography [WMS08]. Step [DRT07]. Stepping [LRB15, TSC13]. Stereo [KRV^{+}14, RSM^{+}15, CLR12, WP10]. Stereographs [ZOH^{+}15]. Stereoscopic [AW15, AWR18, HOH15, MMSO15, SWA14, WPDH14, WC22, CA13, LFM12, SDBRC13, SMI06]. Stewart [BSPB10]. Stiffness [CEM24]. Stimulation [BRM23, LWK18, SHHF^{+}22, VSCM12]. Stimuli [HDF^{+}23, IOYK19, NMVRB20, OOAY22, BMGC05, TGT^{+}09]. Stochastic [CEM24, MC05]. stochastic-based [MC05]. strain [LFM12]. Strategies [BMB19, GWFL22, RNLH16]. strategy [LXXB10]. Streaming [BMB19]. streams [ECOG11]. Stress [LZL^{+}18]. Stress-Induced [LZL^{+}18]. Structural [MVL25]. structure [JSHG08, WP10]. Structures [AWR18]. Strutting [KHW^{+}15]. Students [FWS^{+}23, FGWdS^{+}25]. Studies [BOK10, MMSO15, SF23, VA05]. Study [ARAP^{+}18, DK19, GFD^{+}15, HBM^{+}14, ENC^{+}08, LBT08, LPHL05, MO09a, PDZ05]. Style [ZLWZ24]. Stylization [CSUN05]. stylized [WBC^{+}07]. Subconscious [VCA16]. Subjective [GVC^{+}17, MSS^{+}22, NFD^{+}21, RTSW18, VSCM12]. subjects [APP07, RBCK12]. subpixel [GTAE04]. Substantial [OOAY22]. Substitution [ARAP^{+}18]. Subsurface [GSW^{+}21]. Subthreshold [BRM23]. subtle [MBG09]. sufficient [RVB05]. suggesting [RFR09]. Summary [BCS17]. Supernumerary [RVH^{+}19]. supervised [LXXB10]. Support [FWS^{+}23, MSHLR16, WB04]. Surface [BSH18, MMS13, KSM^{+}05]. surfaces [CWT^{+}05, WBNF06]. Surpasses [PSB^{+}23]. surrealism [SMO^{+}10]. survey [FRC10]. Susceptibility [MJP^{+}24]. symmetric [SMI06]. synchronization [CST^{+}10]. Synthesis [SD22, MAYKM13, WH08]. synthesize [JDR08, MC05]. Synthesized [MR18]. synthetic [OR04]. System [BDW^{+}23, KHH17, UHWT21, WKM^{+}17, VGBF10]. Systematic [TNE20]. systems [FRC10, HU11].

Tablet [PD17, TG19]. Tactile [BSH18, CYK^{+}21, IOYK19, MHPM23, OOAY22, PTP14, PJGE21, DFJ^{+}20, TUG^{+}20, WH08, DFZ^{+}05, RTPG11, WMVO05]. tagging [MP09]. Takala [GLT05a]. talker [BS05b]. Talking [CMR^{+}05, LCC15, MEDO09].

Target [BFCP^{+}25, GHS^{+}20]. targeted [BOK10]. Task [BTDB20, CCV^{+}25, HHNOP19, PUC^{+}24, BGW11, Can09, MBG09]. task-facilitation [BGW11]. Tasks [USA20, AMR06, GTAE04, HHL10, NW08, NVW13, OAD^{+}12]. Technique [KFSN16, FHC04, SII04, WMA12]. Techniques [WIA24, BMGC05, BBD^{+}09]. Teleoperation [PTP14]. Teleportation [NPKR23]. Tell [BRS^{+}25]. temperature [KWI09]. Temporal [EDAM^{+}24, TCP^{+}14]. Temporally [KBL14, MHPM23]. ten [KW05]. Term [SWA14]. Terrain [RKC^{+}22, SD22]. terrains [TSC13]. Test [BPIC24]. Testing [SYF^{+}25]. Text [SF23]. Textual [YO25]. Texture [IOYK19, KHH17, OOAY22, WZA^{+}23, FCH09]. Textured [GVC^{+}17]. Textures [BSH18, IOYK19, PJGE21, HVM06, JDR08, KL06, LKTH06, TJL^{+}11]. their [FRC10]. Theme [YWL^{+}24]. theoretic [FSG09]. Thermal [SJ18, HJ07]. thin [CA13]. those [NCVW10]. Three [FvdMV25, TMM17, TGG^{+}20, WM08]. Three-Dimensional [FvdMV25]. Threshold [ÆÁP^{+}22]. Thresholds [CEN^{+}23, JWB12, VGBF10]. Throwing [PUC^{+}24, SCRTW05, VHBO14]. Time [KPR22, NPKR23, LPEP12, LFM12, LZG^{+}13, SAB07]. time-frequency [LPEP12]. time-to-contact [SAB07]. Together [JOY^{+}18, HHB^{+}24]. Tolerance [AW15]. Tone [GBA17, AR08, AG06, GB08]. tone-mapping [AG06, GB08]. tone-reproduction [AR08]. Tool [Fer19, BC05]. tools [PI08]. Top [MGM16]. Top-Down [MGM16]. torque [VGBF10]. Touch [KBL14]. Touch-Feedback [KBL14]. Touchscreen [KBL14, TG19]. TRAC [USA20]. trackers [MP09]. Tracking [DKM21, PK20, SF23, AMR06, FvdMV25, LME10]. Traditional [JCM^{+}25]. traffic [BGW11, PCK08]. Training [GBTB25, LPO09]. Trait [FMM21]. Traits [BLKD24]. Transfer [BTDB20, MBCW10]. Transfers [GSCRB23]. transit [BWG12]. Transition [KDS^{+}15]. transitions [MO09a, TKK^{+}13]. Translational [USA20]. Translucent [FB05]. Transparency [Fau17, Kaw19, PTP14]. transparent [CA13]. Transport [KPDP19]. travel [BB13, FLKB07]. treadmill [LBWP07, MTCR^{+}07, SGF^{+}10]. treadmill-based [MTCR^{+}07]. Treatment [SYF^{+}25]. trend [NW08]. trend-identification [NW08]. trial [MGM12]. triggering [RVB05]. Triggers [VVC^{+}22]. trimming [MO09b]. Tuning [JOZ^{+}21, SCM18, MO09b]. tunnel [APP07]. Twin [PSB^{+}23]. Two [CEM24, KCK^{+}18, WMVO05]. two-dimensional [WMVO05]. Two-finger [CEM24]. Type [ZHRM15, CYK^{+}21]. typicality [SVHS06].

Ultrasonic [SHHF^{+}22]. Ultrasound [CYK^{+}21, HDF^{+}23]. unattended [DKR^{+}05]. unconstrained [SGS^{+}11]. Underestimation [LLBM15]. Understanding [CCV^{+}25, KPL^{+}19, MB04, YON25, YNYK25, AC11]. undulation [KSM^{+}05]. unified [FSG09]. Uniform [Fau17, FCH09]. Unmodified [DKM21]. Unobtrusive [GATM18]. unstructured [MO09b]. Updates [KT21]. updating [RVB05]. upper [BOK10]. upper-limb [BOK10]. Urban [AL15]. Use [BGK17, PD17, TUG^{+}20, AZ10]. Used [FvdMV25, HFJS09, JDR08]. User [BLKD24, BMB19, BFSV16, BTDB20, EKL^{+}21, FZL20, FGWdS^{+}25, HDH10, KPL^{+}19, YWL^{+}24, GTAE04]. User-Agent [FGWdS^{+}25]. User-based [HDH10]. User-Designer [FGWdS^{+}25]. Using [ÆÁP^{+}22, BBE22, DKM21, FvdMV25, JSHG08, JSCR^{+}15, KKC19, MW15, MC05, MSHLR16, NMVRB20, PK20, PMS17, RTSW18, SAB07, SDG25, Ste15, VCA16, ZLQ^{+}19, AR08, BWG12, CKWB06, GDBP13, GBLR10, KWI09, LME10, MBG09, NAB^{+}11, SDBRC13, TSRD07]. utilizes [KS12]. Utilizing [KHW^{+}15].

Validation [BPIC24, CKWB06]. Value [LZL17]. Variability [LMM^{+}22]. variations [TGJ08]. varying [LKTH06]. Vascular [LZL^{+}18]. vection [RSPA^{+}06, RVSP09]. Vehicle [BBE16]. Verbal [AONB17, GBLR10, NAB^{+}11, ZNWK12]. verbal- [ZNWK12]. Verification [KHJK13]. Versus [MW15, KWSS08, SCRTW05, VSWB07]. Vertical [JKB17, SF23]. Vertices [CPVC19]. via [DFZ^{+}05, KHKP15, PCK08, RKS16, UHWT21]. Vibration [ÆÁP^{+}22, TUG^{+}20]. Vibration-Based [TUG^{+}20]. vibrational [RFR09]. Vibrations [SCM18]. Vibrotactile [ÆÁP^{+}22, HDF^{+}23, IOYK19, OOAY22, PVK20, LSRR13, VVJD05]. Vickers [Vic05]. Video [CKWB05, FKM17, HNT^{+}22, HAHG17, NOSS17, TKK^{+}13]. Videos [NMVRB20, TMM17, WIA24, MGM12]. View [AAZMF21, CKAD18, JKB17, SDW05, WC22, SB12, WCCRT09]. Viewers [WAH^{+}15]. Viewing [MSS^{+}22, RNLH16, CA13, TGT^{+}09]. Viewpoint [PSB^{+}23, EPO11, FSG09, WH08]. viewpoints [SAB07]. views [AC11]. Vigor [RKS16]. Villain [KHW^{+}15]. Virtual [ASG^{+}18, APLK17, BPIC24, BLKD24, BSHW14, BB13, BSH18, BTDB20, BYB18, CCV^{+}25, CEN^{+}23, EBPJ16, FWS^{+}23, FGWdS^{+}25, GSCRB23, GWFL22, GBTB25, HHNOP19, JLS^{+}17, JCM^{+}25, JOY^{+}18, JKB17, JDKN18, JSCR^{+}15, KBL14, KCK^{+}18, KKC19, KLL24, KPR22, KSLM15, LDDR18, LRB15, LVV^{+}20, LGS25, MNG^{+}25, MVL25, NFD^{+}21, PUC^{+}24, RBC14, RSM^{+}15, RKP22, RRM^{+}16, SHHF^{+}22, SYF^{+}25, SXCS15, VVC^{+}22, WTWN16, ZHRM15, ZMM19, ZNO^{+}20, APP07, AC11, BGW11, BSVDD10, BS05a, CWT^{+}05, EPO11, FFW07, FCH^{+}07, FLKB07, GBLR10, GNP^{+}10, HBW11, JWB12, KS12, KBP^{+}13, KL06, KCRT08, LPEP12, LKTH06, LPHL05, LPO09, LBWP07, LSRS10, MBCW10, MJH^{+}09, MC05, MTCR^{+}07, MRT^{+}10, NAB^{+}11, NZG^{+}11, PK07, PKCR05, RBCK12, RSPA^{+}06, RVSP09, RPH10, RDLTS04, RVB13, SCRTW05, SAB07, SGF^{+}10, SGS^{+}11, VSCM12, WAEG06, WCCRT09, WNW^{+}07]. virtual [WBN^{+}11]. viscosity [LKTH06]. Visibility [FO23, GB08, YCK^{+}09]. Vision [EDAM^{+}24, KRV^{+}14, LWK18, SYF^{+}25, APP07, Can09]. Visual [ABK^{+}15, BCS17, BSHW14, BBE16, BDW^{+}23, DCR06, DBR21, FWN^{+}14, GHS^{+}20, GATM18, GVC^{+}17, HHNOP19, JWA19, JLS^{+}17, KT21, KKC19, KM17, LZL17, MD05, MGVM16, NTKA12, PHRE15, RVB05, SGHL^{+}19, Ste15, SS19, UHWT21, VSKL17, WMS08, WL21, YO25, YON25, AMR06, BSH^{+}06, BSVDD10, CKWB06, DBS^{+}09, ENC^{+}08, FLKB07, FRC10, GLT05b, HMS09, HCS10, HS12, Lav09, LZG^{+}13, LPT^{+}06, McN06, MTCR^{+}07, PI08, RLV^{+}10, RBCK12, RDF11, RTPG11, SBR07, SMS13, TG19, TJL^{+}11, VCR08, WMA12]. Visualization [FL09, EML13, HS12, KWSS08, LPR06, LZG^{+}13, VCR08, YBC13]. Visualizations [ALN^{+}21, KPDP19, KPSL10]. Visualizing [WM08]. Visually [AČMS10, RL17, RVSP09]. visually-induced [RVSP09]. Visuo [BTDB20, EBPJ16, PJGE21]. Visuo-Haptic [EBPJ16, BTDB20]. Visuo-Tactile [PJGE21]. Visuomotor [KSLM15]. visuospatial [ZAAC12]. Vive [KCS17, KKC19]. Voice [FWS^{+}23, MVL25, OEMO16, CST^{+}10].

Volume [LME10, MW15, BBD^{+}09]. Voluntary [VSCM12]. VR [AAZMF21, ALN^{+}21, BB21, DKM21]. vs [SNW16].

waist [VVJD05]. walk [RVB13]. Walker [WK05b]. Walking [JLS^{+}17, JKB17, PK07, TSC13, APP07, FFW07, HBW11, SCRTW05, SGF^{+}10, SGS^{+}11, WBN^{+}11]. Wash [ZLWZ24]. water [BAMB13]. wavelets [MC05]. Waypoint [VVJD05]. Web [BFSV16]. Webpage [SS19]. Weighted [PK20]. Welcome [IG15]. Which [SB12]. Who [HHB^{+}24]. widgets [Bre05, EM05b]. Width [JSCR^{+}15]. Wii [WBN^{+}11]. Wind [NMVRB20]. without [MLK^{+}06]. Women [AAZMF21]. Word [RTJ^{+}21, ZYJ^{+}22]. Words [RTJ^{+}21]. Workload [DBR21, TNE20]. workstations [AMR06]. World [ASG^{+}18, Can09, FK19, WBC^{+}07]. worlds [LPEP12]. Worn [JKB17, BWG12]. Wrinkles [NP15]. Wrist [ÆÁP^{+}22]. Writing [AZ10].

yaw [JWB12].

References

Amemiya:2008:LMI

[AAM08] Tomohiro Amemiya, Hideyuki Ando, and Taro Maeda. Leadme interface for a pulling sensation from hand-held devices. ACM Transactions on Applied Perception, 5(3):15:1– 15:??, August 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Alonso-Arevalo:2012:CSC

[AASH^{+}12] Miguel A. Alonso-Arevalo, Simon Shelley, Dik Hermes, Jacqueline Hollowood, Michael Pettitt, Sarah Sharples, and Armin Kohlrausch. Curve shape and curvature perception through interactive sonification. ACM Transactions on Applied Perception, 9(4):17:1– 17:??, October 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Adhanom:2021:FVR

[AAZMF21] Isayas Berhe Adhanom, Majed Al-Zayer, Paul Macneilage, and Eelke Folmer. Field-of-view restriction to reduce VR sickness does not impede spatial learning in women. ACM Transactions on Applied Perception, 18(2):5:1–5:17, June 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3448304.

Abhari:2015:VEM

[ABK^{+}15] Kamyar Abhari, John S. H. Baxter, Ali R. Khan, Terry M. Peters, Sandrine De Ribaupierre, and Roy Eagleson. Visual enhancement of MR angiography images to facilitate planning of arteriovenous malformation interventions. ACM Transactions on Applied Perception, 12(1):4:1–4:??, March 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Au:2011:IMV

[AC11] Carmen E. Au and James J. Clark. Integrating multiple views with virtual mirrors to facilitate scene understanding. ACM Transactions on Applied Perception, 8(4):28:1– 28:??, November 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Aydin:2010:VSE

[AČMS10] Tunç Ozan Aydin, Martin Čadík, Karol Myszkowski, and Hans-Peter Seidel. Visually significant edges. ACM Transactions on Applied Perception, 7 (4):27:1–27:??, July 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Aevarsson:2022:VTM

[ÆÁP^{+}22] Elvar Atli Ævarsson, Thórhildur Ásgeirsdóttir, Finnur Pind, Árni Kristjánsson, and Runar Unnthorsson. Vibrotactile threshold measurements at the wrist using parallel vibration actuators. ACM Transactions on Applied Perception, 19(3): 10:1–10:??, July 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3529259.

Ashikhmin:2006:RCT

[AG06] Michael Ashikhmin and Jay Goyal. A reality check for tone-mapping operators. ACM Transactions on Applied Perception, 3(4):399–411, October 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Argelaguet:2013:EIP

[AJML13] Ferran Argelaguet, David Antonio Gómez Jáuregui, Maud Marchal, and Anatole Lécuyer. Elastic images: Perceiving local elasticity of images through a novel pseudo-haptic deformation effect. ACM Transactions on Applied Perception, 10(3): 17:1–17:??, August 2013. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Akyuz:2016:PME

[AK16] Ahmet Oguz Akyüz and Osman Kaya. A proposed methodology for evaluating HDR false color maps. ACM Transactions on Applied Perception, 14(1):2:1– 2:??, August 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Albrecht:2015:ADP

[AL15] Robert Albrecht and Tapio Lokki. Auditory distance presentation in an urban augmented reality environment. ACM Transactions on Applied Perception, 12(2):5:1–5:??, April 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Adkins:2021:EGV

[ALN^{+}21] Alex Adkins, Lorraine Lin, Aline Normoyle, Ryan Canales, Yuting Ye, and Sophie Jörg. Evaluating grasping visualizations and control modes in a VR game. ACM Transactions on Applied Perception, 18 (4):19:1–19:14, October 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3486582.

Atkins:2006:AET

[AMR06] M. Stella Atkins, Adrian Moise, and Robert Rohling. An application of eyegaze tracking for designing radiologists’ workstations: Insights for comparative visual search tasks. ACM Transactions on Applied Perception, 3(2):136–151, April 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Alexanderson:2017:MIE

[AONB17] Simon Alexanderson, Carol O’sullivan, Michael Neff, and Jonas Beskow. Mimebot — investigating the expressibility of non-verbal communication across agent embodiments. ACM Transactions on Applied Perception, 14(4):24:1–24:??, September 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Abebe:2015:ECF

[APK15] Mekides Assefa Abebe, Tania Pouli, and Jonathan Kervec. Evaluating the color fidelity of ITMOs and HDR color appearance models. ACM Transactions on Applied Perception, 12 (4):14:1–14:??, September 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Albert:2017:LRF

[APLK17] Rachel Albert, Anjul Patney, David Luebke, and Joohwan Kim. Latency requirements for foveated rendering in virtual reality. ACM Transactions on Applied Perception, 14(4):25:1– 25:??, September 2017. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Abebe:2017:PLM

[APLR17] Mekides Assefa Abebe, Tania Pouli, Mohamed-Chaker Larabi, and Erik Reinhard. Perceptual lightness modeling for highdynamic-range imaging. ACM Transactions on Applied Perception, 15(1):1:1–1:??, November 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Apfelbaum:2007:HAT

[APP07] Henry Apfelbaum, Adar Pelah, and Eli Peli. Heading assessment by “tunnel vision” patients and control subjects standing or walking in a virtual reality environment. ACM Transactions on Applied Perception, 4(1):??, January 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Akyuz:2008:PET

[AR08] Ahmet Oğuz Akyüz and Erik Reinhard. Perceptual evaluation of tone-reproduction operators using the cornsweet– craik–O’brien illusion. ACM Transactions on Applied Perception, 4(4):1:1–1:??, January 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Aviles-Rivero:2018:SSF

[ARAP^{+}18] Angelica I. Aviles-Rivero, Samar M.

Alsaleh, John Philbeck, Stella P. Raventos, Naji Younes, James K. Hahn, and Alicia Casals. Sensory substitution for force feedback recovery: a perception experimental study. ACM Transactions on Applied Perception, 15(3):16:1–16:??, August 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Agethen:2018:BAH

[ASG^{+}18] Philipp Agethen, Viswa Subramanian Sekar, Felix Gaisbauer, Thies Pfeiffer, Michael Otto, and Enrico Rukzio. Behavior analysis of human locomotion in the real world and virtual reality for the manufacturing industry. ACM Transactions on Applied Perception, 15(3):20:1– 20:??, August 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Alkasasbeh:2021:WSH

[ASG21] Anas Ali Alkasasbeh, Fotios Spyridonis, and Gheorghita Ghinea. When scents help me remember my password. ACM Transactions on Applied Perception, 18(3):16:1–16:18, July 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3469889.

Arabadzhiyska:2023:PSP

[ATSD23] Elena Arabadzhiyska, Cara Tursun, Hans-Peter Seidel, and Piotr Didyk. Practical saccade prediction for head-mounted displays: Towards a comprehensive model. ACM Transactions on Applied Perception, 20(1):2:1–2:??, January 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3568311.

Allison:2015:PTS

[AW15] Robert S. Allison and Laurie M. Wilcox. Perceptual tolerance to stereoscopic 3D image distortion. ACM Transactions on Applied Perception, 12(3):10:1– 10:??, July 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Aygar:2018:CSM

[AWR18] Erol Aygar, Colin Ware, and David Rogers. The contribution of stereoscopic and motion depth cues to the perception of structures in 3D point clouds. ACM Transactions on Applied Perception, 15(2):9:1– 9:??, April 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Andersen:2010:WME

[AZ10] Tue Haste Andersen and Shumin Zhai. “writing with music”: Exploring the use of auditory feedback in gesture interfaces. ACM Transactions on Applied Perception, 7(3):17:1– 17:??, June 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bojrab:2013:PIL

[BAMB13] Micah Bojrab, Michel Abdul-Massih, and Bedrich Benes. Perceptual importance of lighting phenomena in rendering of animated water. ACM Transactions on Applied Perception, 10(1):2:1–2:??, February 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Barrass:2005:CFA

[Bar05a] Stephen Barrass. A comprehensive framework for auditory display: Comments on Barrass, ICAD 1994. ACM Transactions on Applied Perception, 2 (4):403–406, October 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [Bar05b].

Barrass:2005:PFA

[Bar05b] Stephen Barrass. A perceptual framework for the auditory display of scientific data. ACM Transactions on Applied Perception, 2(4):389–402, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [Bar05a].

Blom:2013:VTC

[BB13] Kristopher J. Blom and Steffi Beckhaus. Virtual travel collisions: Response method influences perceived realism of virtual environments. ACM Transactions on Applied Perception, 10(4):25:1–25:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Brickler:2021:ESP

[BB21] David Brickler and Sabarish V. Babu. An evaluation of screen parallax, haptic feedback, and sensory-motor mismatch on near-field perception-action coordination in VR. ACM Transactions on Applied Perception, 18(4):20:1–20:16, October 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3486583.

Boucheny:2009:PEV

[BBD^{+}09] Christian Boucheny, Georges-Pierre Bonneau, Jacques Droulez, Guillaume Thibault, and Stephane Ploix. A perceptive evaluation of volume rendering techniques. ACM Transactions on Applied Perception, 5(4):23:1– 23:??, January 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Blissing:2016:EVL

[BBE16] Björn Blissing, Fredrik Bruzelius, and Olle Eriksson. Effects of visual latency on vehicle driving behavior. ACM Transactions on Applied Perception, 14(1): 5:1–5:??, August 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Blissing:2022:EDB

[BBE22] Björn Blissing, Fredrik Bruzelius, and Olle Eriksson. The effects on driving behavior when using a head-mounted display in a dynamic driving simulator. ACM Transactions on Applied Perception, 19(1):4:1– 4:18, January 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3483793.

Brewster:2005:DES

[BC05] Stephen A. Brewster and Catherine V. Clarke. The design and evaluation of a sonically enhanced tool palette. ACM Transactions on Applied Perception, 2(4):455–461, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [Bre05].

Beshai:2020:PSP

[BCB20] Peter Beshai, Ricardo Caceffo, and Kellogg S. Booth. Providing semi-private feedback on a shared public screen by controlling presentation onset. ACM Transactions on Applied Perception, 17(3):11:1–11:32, November 2020. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3419983.

Bhardwaj:2015:DAP

[BCD15] Amit Bhardwaj, Subhasis Chaudhuri, and Onkar Dabeer.

Design and analysis of predictive sampling of haptic signals. ACM Transactions on Applied Perception, 11(4):16:1– 16:??, January 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Balas:2017:SSM

[BCS17] Benjamin Balas, Catherine Conlin, and Dylan Shipman. Summary statistics and material categorization in the visual periphery. ACM Transactions on Applied Perception, 14(2): 8:1–8:??, February 2017. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bressolette:2021:MGS

[BDR^{+}21] Benjamin Bressolette, Sébastien Denjean, Vincent Roussarie, Mitsuko Aramaki, Sølvi Ystad, and Richard Kronland-Martinet. MovEcho: a gesturesound interface allowing blind manipulations in a driving context. ACM Transactions on Applied Perception, 18(3):15:1– 15:19, July 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3464692.

Brown:2023:EDM

[BDW^{+}23] Rachel Brown, Vasha Dutell, Bruce Walter, Ruth Rosenholtz, Peter Shirley, Morgan McGuire, and David Luebke. Efficient dataflow modeling of peripheral encoding in the human visual system. ACM Transactions on Applied Perception, 20(1):1:1–1:??, January 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3564605.

Beech:2025:HCM

[BFCP^{+}25] Sam Beech, Danaë Stanton Fraser, Andy Corston-Petrie, Andy P. Gower, and Iain D. Gilchrist. How changes in the mean latency, jitter amplitude, and jitter frequency impact target acquisition performance. ACM Transactions on Applied Perception, 22(2): 8:1–8:??, April 2025. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3701984.

Boi:2016:RUA

[BFSV16] Paolo Boi, Gianni Fenu, Lucio Davide Spano, and Valentino Vargiu. Reconstructing user’s attention on the Web through mouse movements and perception-based content identification. ACM Transactions on Applied Perception, 13(3): 15:1–15:??, May 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Berman:2017:EUS

[BGK17] Lewis Berman, Keith Gallagher, and Suzanne Kozaitis. Evaluating the use of sound in static program comprehension. ACM Transactions on Applied Perception, 15(1):7:1–7:??, November 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bicego:2008:DFC

[BGL^{+}08] Manuele Bicego, Enrico Grosso, Andrea Lagorio, Gavin Brelstaff, Linda Brodo, and Massimo Tistarelli. Distinctiveness of faces: a computational approach. ACM Transactions on Applied Perception, 5(2):11:1– 11:??, May 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bernhard:2011:BTF

[BGW11] Matthias Bernhard, Karl Grosse, and Michael Wimmer. Bimodal task-facilitation in a virtual traffic scenario through spatialized sound rendering. ACM Transactions on Applied Perception, 8(4):24:1– 24:??, November 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Breeden:2017:GDA

[BH17] Katherine Breeden and Pat Hanrahan. Gaze data for the analysis of attention in feature films. ACM Transactions on Applied Perception, 14(4):23:1– 23:??, September 2017. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bouchara:2013:CMS

[BJK13] Tifanie Bouchara, Christian Jacquemin, and Brian F. G. Katz. Cueing multimedia search with audiovisual blur.

ACM Transactions on Applied Perception, 10(2):7:1–7:??, May 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bendiksen:2024:AHR

[BLKD24] Bennie Bendiksen, Nana Lin, Jiehyun Kim, and Funda Durupinar. Assessing human reactions in a virtual crowd based on crowd disposition, perceived agency, and user traits. ACM Transactions on Applied Perception, 21(3):9:1–9:??, July 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3658670.

Bonebright:2005:EAD

[BM05] Terri L. Bonebright and Nadine E. Miner. Evaluation of auditory displays: Comments on Bonebright et al., ICAD 1998. ACM Transactions on Applied Perception, 2(4):517–520, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [BMGC05].

Bhargava:2019:CEU

[BMB19] Ayush Bhargava, James Martin, and Sabarish V. Babu. Comparative evaluation of user perceived quality assessment of design strategies for HTTP-based adaptive streaming. ACM Transactions on Applied Perception, 16(4):22:1– 22:??, September 2019. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3345313.

Bonebright:2005:DCA

[BMGC05] Terri L. Bonebright, Nadine E. Miner, Timothy E. Goldsmith, and Thomas P. Caudell. Data collection and analysis techniques for evaluating the perceptual qualities of auditory stimuli. ACM Transactions on Applied Perception, 2(4):505– 516, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [BM05].

Bodenheimer:2009:GE

[BO09] Bobby Bodenheimer and Carol O’Sullivan. Guest editorial. ACM Transactions on Applied Perception, 6(4):21:1– 21:??, September 2009. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bodenheimer:2021:E

[Bod21] Bobby Bodenheimer. Editorial. ACM Transactions on Applied Perception, 18(4):17:1– 17:2, October 2021. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3486957.

Blank:2010:IRP

[BOK10] Amy Blank, Allison M. Okamura, and Katherine J. Kuchenbecker. Identifying the role of proprioception in upperlimb prosthesis control: Studies on targeted motion. ACM Transactions on Applied Perception, 7(3):15:1–15:??, June 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bernardo:2016:IQU

[BPFP16] Marco V. Bernardo, António M. G. Pinheiro, Paulo T. Fiadeiro, and Manuela Pereira. Image quality under chromatic impairments. ACM Transactions on Applied Perception, 14 (1):6:1–6:??, August 2016. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bartlett:2024:DVV

[BPIC24] Kristin A. Bartlett, Almudena Palacios-Ibáñez, and Jorge Dorribo Camba. Design and validation of a virtual reality mental rotation test. ACM Transactions on Applied Perception, 21(2):5:1–5:??, April 2024. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3626238.

Brewster:2005:SEW

[Bre05] Stephen A. Brewster. Sonicallyenhanced widgets: Comments on Brewster and Clarke, ICAD 1997. ACM Transactions on Applied Perception, 2(4):462– 466, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [BC05].

Balasubramanian:2023:ESE

[BRM23] Jagan Krishnasamy Balasubramanian, Rahul Kumar Ray, and Manivannan Muniyandi. Effect of subthreshold electrotactile stimulation on the perception of electrovibration. ACM Transactions on Applied Perception, 20(3):9:1–9:??, July 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3599970.

Bozkir:2025:CYT

[BRS^{+}25] Efe Bozkir, Clara Riedmiller, Athanassios N. Skodras, Gjergji Kasneci, and Enkelejda Kasneci. Can you tell real from fake face images? Perception of computer-generated faces by humans. ACM Transactions on Applied Perception, 22(2): 6:1–6:??, April 2025. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3696667.

Brungart:2005:OVS

[BS05a] Douglas S. Brungart and Brian D. Simpson. Optimizing a virtual speech display: Comments on Brungart and Simpson, ICAD 2003. ACM Transactions on Applied Perception, 2(4):437–441, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [BS05b].

Brungart:2005:OSC

[BS05b] Douglas S. Brungart and Brian D. Simpson. Optimizing the spatial configuration of a seven-talker speech display. ACM Transactions on Applied Perception, 2(4):430–436, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [BS05a].

Balas:2006:RBR

[BS06] Benjamin J. Balas and Pawan Sinha. Region-based representations for face recognition. ACM Transactions on Applied Perception, 3(4):354–375, October 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Barbagli:2006:HDF

[BSH^{+}06] Federico Barbagli, Ken Salisbury, Cristy Ho, Charles Spence, and Hong Z. Tan. Haptic discrimination of force direction and the influence of visual information. ACM Transactions on Applied Perception, 3(2):125–135, April 2006. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bochereau:2018:PCR

[BSH18] Séréna Bochereau, Stephen Sinclair, and Vincent Hayward. Perceptual constancy in the reproduction of virtual tactile textures with surface displays. ACM Transactions on Applied Perception, 15(2):10:1– 10:??, April 2018. CODEN ????

ISSN 1544-3558 (print), 1544-3965 (electronic).

Bernhard:2014:GOM

[BSHW14] Matthias Bernhard, Efstathios Stavrakis, Michael Hecher, and Michael Wimmer. Gaze-to-object mapping during visual search in 3D virtual environments. ACM Transactions on Applied Perception, 11(3):14:1– 14:??, August 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Berger:2010:SBF

[BSPB10] Daniel R. Berger, Jörg Schulte-Pelkum, and Heinrich H. Bülthoff. Simulating believable forward accelerations on a Stewart motion platform. ACM Transactions on Applied Perception, 7(1):5:1–5:??, January 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bonneel:2010:BPA

[BSVDD10] Nicolas Bonneel, Clara Suied, Isabelle Viaud-Delmon, and George Drettakis. Bimodal perception of audio-visual material properties for virtual environments. ACM Transactions on Applied Perception, 7(1):1:1– 1:??, January 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bernhard:2010:EPD

[BSW10] Matthias Bernhard, Efstathios Stavrakis, and Michael Wimmer. An empirical pipeline to derive gaze prediction heuristics for 3D action games. ACM Transactions on Applied Perception, 8(1):4:1–4:??, October 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Bailey:2016:ISI

[BT16] Reynold Bailey and Laura Trutoiu. Introduction to special issue SAP 2016. ACM Transactions on Applied Perception, 13 (4):18:1–18:??, July 2016. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Brickler:2020:FLE

[BTDB20] David Brickler, Robert J.

Teather, Andrew T. Duchowski, and Sabarish V. Babu. A Fitts’ Law evaluation of visuo-haptic fidelity and sensory mismatch on user performance in a nearfield disc transfer task in virtual reality. ACM Transactions on Applied Perception, 17(4):15:1– 15:20, December 2020. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3419986.

Bulling:2012:MRR

[BWG12] Andreas Bulling, Jamie A.

Ward, and Hans Gellersen. Multimodal recognition of reading activity in transit using body-worn sensors. ACM Transactions on Applied Perception, 9(1):2:1–2:??, March 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Buck:2018:CDE

[BYB18] Lauren E. Buck, Mary K. Young, and Bobby Bodenheimer. A comparison of distance estimation in HMD-based virtual environments with different HMD-based conditions. ACM Transactions on Applied Perception, 15(3):21:1–21:??, August 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Chen:2013:SPT

[CA13] Jianhui Chen and Robert S. Allison. Shape perception of thin transparent objects with stereoscopic viewing. ACM Transactions on Applied Perception, 10(3):15:1–15:??, August 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Canosa:2009:RWV

[Can09] Roxanne L. Canosa. Real-world vision: Selective perception and task. ACM Transactions on Applied Perception, 6(2):11:1– 11:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Caramiaux:2014:RSS

[CBB^{+}14] B. Caramiaux, F. Bevilacqua, T. Bianco, N. Schnell, O. Houix, and P. Susini. The role of sound source perception in gestural sound description. ACM Transactions on Applied Perception, 11(1):1:1– 1:??, April 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Choi:2025:TUE

[CCV^{+}25] Minsoo Choi, Dixuan Cui, Matias Volonte, Alexandros Koilias, Dominic Kao, and Christos Mousas. Toward understanding the effects of intelligence of a virtual character during an immersive jigsaw puzzle co-solving task. ACM Transactions on Applied Perception, 22 (2):9:1–9:??, April 2025. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3700822.

Chamnongthai:2024:TFS

[CEM24] Komi Chamnongthai, Takahiro Endo, and Fumitoshi Matsuno. Two-finger stiffness discrimination with the stochastic resonance effect. ACM Transactions on Applied Perception, 21 (2):6:1–6:??, April 2024. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3630254.

Choudhary:2023:VBH

[CEN^{+}23] Zubin Choudhary, Austin Erickson, Nahal Norouzi, Kangsoo Kim, Gerd Bruder, and Gregory Welch. Virtual big heads in extended reality: Estimation of ideal head scales and perceptual thresholds for comfort and facial cues. ACM Transactions on Applied Perception, 20(1):4:1–4:??, January 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3571074.

Chapiro:2018:ISS

[CKAD18] Alexandre Chapiro, Timo Kunkel, Robin Atkins, and Scott Daly. Influence of screen size and field of view on perceived brightness. ACM Transactions on Applied Perception, 15(3):18:1–18:??, August 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Cunningham:2005:MVS

[CKWB05] Douglas W. Cunningham, Mario Kleiner, Christian Wallraven, and Heinrich H. Bülthoff. Manipulating video sequences to determine the components of conversational facial expressions. ACM Transactions on Applied Perception, 2(3):251– 269, July 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Cooke:2006:OFV

[CKWB06] Theresa Cooke, Sebastian Kannengiesser, Christian Wallraven, and Heinrich H. Bülthoff. Object feature validation using visual and haptic similarity ratings. ACM Transactions on Applied Perception, 3(3):239– 261, July 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Couture:2010:ADD

[CLR10] Vincent Couture, Michael S.

Langer, and Sébastien Roy. Analysis of disparity distortions in omnistereoscopic displays. ACM Transactions on Applied Perception, 7(4):25:1– 25:??, July 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Couture:2012:PBS

[CLR12] Vincent Couture, Michael S.

Langer, and Sébastien Roy. Perception of blending in stereo motion panoramas. ACM Transactions on Applied Perception, 9(3):15:1–15:??, July 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Cosker:2005:TPR

[CMR^{+}05] Darren Cosker, David Marshall, Paul L. Rosin, Susan Paddock, and Simon Rushton. Toward perceptually realistic talking heads: Models, methods, and McGurk. ACM Transactions on Applied Perception, 2 (3):270–285, July 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Company:2019:APV

[CPVC19] Pedro Company, Raquel Plumed, Peter A. C. Varley, and Jorge D. Camba. Algorithmic perception of vertices in sketched drawings of polyhedral shapes. ACM Transactions on Applied Perception, 16(3):18:1– 18:??, September 2019. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3345507.

Cabral:2022:DAI

[CR22] João P. Cabral and Gerard B. Remijn. The duration of an auditory icon can affect how the listener interprets its meaning. ACM Transactions on Applied Perception, 19(2): 8:1–8:16, April 2022. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3527269.

Chapiro:2023:ISS

[CR23] Alexandre Chapiro and Andrew Robb. Introduction to the SAP 2023 special issue. ACM Transactions on Applied Perception, 20(4):12:1–12:??, October 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3629977.

Creem-Regehr:2009:GE

[CRM09] Sarah Creem-Regehr and Karol Myszkowski. Guest editorial. ACM Transactions on Applied Perception, 6(3):13:1– 13:??, August 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Carter:2010:PMG

[CST^{+}10] Elizabeth J. Carter, Lavanya Sharan, Laura Trutoiu, Iain Matthews, and Jessica K. Hodgins. Perceptually motivated guidelines for voice synchronization in film. ACM Transactions on Applied Perception, 7 (4):23:1–23:??, July 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Chang:2005:EBC

[CSUN05] Youngha Chang, Suguru Saito, Keiji Uchikawa, and Masayuki Nakajima. Example-based color stylization of images. ACM Transactions on Applied Perception, 2(3):322–345, July 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Cooke:2010:MSA

[CWB10] Theresa Cooke, Christian Wallraven, and Heinrich H. Bülthoff. Multidimensional scaling analysis of haptic exploratory procedures. ACM Transactions on Applied Perception, 7(1): 7:1–7:??, January 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Choi:2005:FCE

[CWT^{+}05] Seungmoon Choi, Laron Walker, Hong Z. Tan, Scott Crittenden, and Ron Reifenberger. Force constancy and its effect on haptic perception of virtual surfaces. ACM Transactions on Applied Perception, 2(2):89– 105, April 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Chen:2014:NHF

[CXZ14] Fangmei Chen, Yong Xu, and David Zhang. A new hypothesis on facial beauty perception. ACM Transactions on Applied Perception, 11(2):8:1–8:??, July 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Choi:2021:QDT

[CYK^{+}21] Jeongbong Choi, Soonhyun Yook, In Young Kim, Mok Kun Jeong, and Dong Pyo Jang. Quantification of displacement for tactile sensation in a contact-type low intensity focused ultrasound haptic device. ACM Transactions on Applied Perception, 18(1):1:1– 1:8, January 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3422820.

Durupinar:2025:ISI

[DB25] Funda Durupinar and Sabarish Babu. Introduction to the special issue on SAP 2025. ACM Transactions on Applied Perception, 22(4):13:1–13:2, October 2025. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Devlin:2021:DWW

[DBR21] Shannon P. Devlin, Jennifer K. Byham, and Sara Lu Riggs. Does what we see shape history? Examining workload history as a function of performance and ambient/focal visual attention. ACM Transactions on Applied Perception, 18 (2):8:1–8:17, June 2021. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3449066.

Duchowski:2009:SVS

[DBS^{+}09] Andrew T. Duchowski, David Bate, Paris Stringfellow, Kaveri Thakur, Brian J. Melloy, and Anand K. Gramopadhye. On spatiochromatic visual sensitivity and peripheral color LOD management. ACM Transactions on Applied Perception, 6(2):9:1–9:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Dixon:2006:MAF

[DCN^{+}06] T. D. Dixon, E. F. Canga, J. M. Noyes, T. Troscianko, S. G. Nikolov, D. R. Bull, and C. N. Canagarajah. Methods for the assessment of fused images. ACM Transactions on Applied Perception, 3(3):309– 332, July 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Devlin:2006:VCC

[DCR06] Kate Devlin, Alan Chalmers, and Erik Reinhard. Visual calibration and correction for ambient illumination. ACM Transactions on Applied Perception, 3(4):429–452, October 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

DePoli:2015:RID

[DCRS15] Giovanni De Poli, Sergio Canazza, Antonio Rodà, and Emery Schubert. The role of individual difference in judging expressiveness of computerassisted music performances by experts. ACM Transactions on Applied Perception, 11(4):22:1– 22:??, January 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Pra:2020:DIP

[DFJ^{+}20] Yuri De Pra, Federico Fontana, Hanna Järveläinen, Stefano Papetti, and Michele Simonato. Does it ping or pong? Auditory and tactile classification of materials by bouncing events. ACM Transactions on Applied Perception, 17 (2):8:1–8:17, May 2020. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3393898.

Drewing:2005:FEN

[DFZ^{+}05] Knut Drewing, Michael Fritschi, Regine Zopf, Marc O. Ernst, and Martin Buss. First evaluation of a novel tactile display exerting shear force via lateral displacement. ACM Transactions on Applied Perception, 2 (2):118–131, April 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Dodge:2019:HDC

[DK19] Samuel Dodge and Lina Karam.

Human and DNN classification performance on images with quality distortions: a comparative study. ACM Transactions on Applied Perception, 16(2):7:1–7:??, August 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3306241.

Drakopoulos:2021:ETI

[DKM21] Panagiotis Drakopoulos, George-Alex Koulieris, and Katerina Mania. Eye tracking interaction on unmodified mobile VR headsets using the selfie camera. ACM Transactions on Applied Perception, 18(3):11:1– 11:20, July 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3456875.

Dinse:2005:IHH

[DKR^{+}05] Hubert R. Dinse, Tobias Kalisch, Patrick Ragert, Burkhard Pleger, Peter Schwenkreis, and Martin Tegenthoff. Improving human haptic performance in normal and impaired human populations through unattended activation-based learning. ACM Transactions on Applied Perception, 2(2):71–88, April 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Durgin:2007:SFP

[DRT07] Frank H. Durgin, Catherine Reed, and Cara Tigue. Step frequency and perceived selfmotion. ACM Transactions on Applied Perception, 4(1):??, January 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ehret:2021:DPE

[EBA^{+}21] Jonathan Ehret, Andrea Bönsch, Lukas Aspöck, Christine T. Röhr, Stefan Baumann, Martine Grice, Janina Fels, and Torsten W. Kuhlen. Do prosody and embodiment influence the perceived naturalness of conversational agents’ speech? ACM Transactions on Applied Perception, 18(4):21:1–21:15, October 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3486580.

Ebrahimi:2016:EEV

[EBPJ16] Elham Ebrahimi, Sabarish V. Babu, Christopher C. Pagano, and Sophie Jörg. An empirical evaluation of visuo-haptic feedback on physical reaching behaviors during 3D interaction in real and immersive virtual environments. ACM Transactions on Applied Perception, 13(4): 19:1–19:??, July 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Endres:2011:EHO

[ECOG11] Dominik Endres, Andrea Christensen, Lars Omlor, and Martin A. Giese. Emulating human observers with Bayesian binning: Segmentation of action streams. ACM Transactions on Applied Perception, 8(3):16:1– 16:??, August 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ebelin:2024:ETE

[EDAM^{+}24] Pontus Ebelin, Gyorgy Denes, Tomas Akenine-Möller, Kalle Åström, Magnus Oskarsson, and William H. McIlhagga. Estimates of temporal edge detection filters in human vision. ACM Transactions on Applied Perception, 21(2):7:1– 7:??, April 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639052.

Erickson:2021:EAB

[EKL^{+}21] Austin Erickson, Kangsoo Kim, Alexis Lambert, Gerd Bruder, Michael P. Browne, and Gregory F. Welch. An extended analysis on the benefits of dark mode user interfaces in optical see-through head-mounted displays. ACM Transactions on Applied Perception, 18(3): 12:1–12:22, July 2021. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3456874.

Edwards:2005:PAD

[EM05a] Alistair D. N. Edwards and Evangelos Mitsopoulos. Perceptual auditory design: Comments on Edwards and Mitsopoulos, ICAD 1998. ACM Transactions on Applied Perception, 2(4):450–454, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [EM05b].

Edwards:2005:PMS

[EM05b] Alistair D. N. Edwards and Evangelos Mitsopoulos. A principled methodology for the specification and design of nonvisual widgets. ACM Transactions on Applied Perception, 2 (4):442–449, October 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [EM05a].

Easa:2013:EMD

[EML13] Haider K. Easa, Rafal K. Mantiuk, and Ik Soo Lim. Evaluation of monocular depth cues on a high-dynamic-range display for visualization. ACM Transactions on Applied Perception, 10(3):16:1–16:??, August 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Elhelw:2008:GBS

[ENC^{+}08] Mohamed Elhelw, Marios Nicolaou, Adrian Chung, Guang-Zhong Yang, and M. Stella Atkins. A gaze-based study for investigating the perception of visual realism in simulated scenes. ACM Transactions on Applied Perception, 5 (1):3:1–3:??, January 2008. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ennis:2011:PES

[EPO11] Cathy Ennis, Christopher Peters, and Carol O’Sullivan. Perceptual effects of scene context and viewpoint for virtual pedestrian crowds. ACM Transactions on Applied Perception, 8(2):10:1–10:??, January 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Faul:2017:TPU

[Fau17] Franz Faul. Toward a perceptually uniform parameter space for filter transparency. ACM Transactions on Applied Perception, 14(2):13:1–13:??, February 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Fleming:2005:LLI

[FB05] Roland W. Fleming and Heinrich H. Bülthoff. Low-level image cues in the perception of translucent materials. ACM Transactions on Applied Perception, 2(3):346–382, July 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Flowers:2005:DSD

[FBT05] John H. Flowers, Dion C. Buhman, and Kimberly D. Turnage. Data sonification from the desktop: Should sound be part of standard data analysis software? ACM Transactions on Applied Perception, 2 (4):467–472, October 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [FTB05].

Fortenbaugh:2007:GDC

[FCH^{+}07] Francesca C. Fortenbaugh, Sidhartha Chaudhury, John C. Hicks, Lei Hao, and Kathleen A. Turano. Gender differences in cue preference during path integration in virtual environments. ACM Transactions on Applied Perception, 4 (1):??, January 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Filip:2009:URG

[FCH09] Jiří Filip, Michael J. Chantler, and Michal Haindl. On uniform resampling and gaze analysis of bidirectional texture functions. ACM Transactions on Applied Perception, 6(3):18:1– 18:??, August 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Fernstrom:2005:RSB

[Fer05] Mikael Fernström. Reflections on sonic browsing: Comments on Fernström and McNamara, ICAD 1998. ACM Transactions on Applied Perception, 2 (4):500–504, October 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [FM05].

Ferwerda:2019:FHT

[Fer19] James Ferwerda. The FechDeck: a hand tool for exploring psychophysics. ACM Transactions on Applied Perception, 16(2):9:1–9:??, August 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3313186.

Fink:2007:OAD

[FFW07] Philip W. Fink, Patrick S. Foo, and William H. Warren. Obstacle avoidance during walking in real and virtual environments. ACM Transactions on Applied Perception, 4(1):??, January 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Feijoo-Garcia:2025:EEU

[FGWdS^{+}25] Pedro Guillermo Feijóo-García, Chase Wrenn, Alexandre Gomes de Siqueira, Rashi Ghosh, Jacob Stuart, Heng Yao, and Benjamin Lok. Exploring the effects of user-agent and userdesigner similarity in virtual human design to promote mental health intentions for college students. ACM Transactions on Applied Perception, 22(1):5:1– 5:??, January 2025. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3689822.

Frowd:2004:EHE

[FHC04] Charlie D. Frowd, Peter J. B.

Hancock, and Derek Carson. EvoFIT: a holistic, evolutionary facial imaging technique for creating composites. ACM Transactions on Applied Perception, 1(1):19–39, July 2004. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Filip:2019:PAA

[FK19] Jirí Filip and Martina Kolafová. Perceptual attributes analysis of real-world materials. ACM Transactions on Applied Perception, 16(1):1:1–1:??, February 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3301412.

Ferstl:2017:FFN

[FKM17] Ylva Ferstl, Elena Kokkinara, and Rachel Mcdonnell. Facial features of non-player creatures can influence moral decisions in video games. ACM Transactions on Applied Perception, 15(1):4:1–4:??, November 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Fleming:2009:GES

[FL09] Roland Fleming and Michael Langer. Guest editorial: Special issue on Applied Perception in Graphics and Visualization (APGV07). ACM Transactions on Applied Perception, 5(4):18:1–18:??, January 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Frenz:2007:ETD

[FLKB07] Harald Frenz, Markus Lappe, Marina Kolesnik, and Thomas Bührmann. Estimation of travel distance from visual motion in virtual environments. ACM Transactions on Applied Perception, 4(1):??, January 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Fernstrom:2005:ADM

[FM05] Mikael Fernström and Caolan McNamara. After direct manipulation—direct sonification. ACM Transactions on Applied Perception, 2(4):495–499, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [Fer05].

Ferstl:2021:FFM

[FMM21] Ylva Ferstl, Michael McKay, and Rachel McDonnell. Facial feature manipulation for trait portrayal in realistic and cartoon-rendered characters. ACM Transactions on Applied Perception, 18(4):22:1– 22:8, October 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3486579.

Fukiage:2023:CAV

[FO23] Taiki Fukiage and Takeshi Oishi. A content-adaptive visibility predictor for perceptually optimized image blending. ACM Transactions on Applied Perception, 20(1):3:1– 3:??, January 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3565972.

Fontana:2008:ADP

[FR08] Federico Fontana and Davide Rocchesso. Auditory distance perception in an acoustic pipe. ACM Transactions on Applied Perception, 5(3):16:1– 16:??, August 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Frintrop:2010:CVA

[FRC10] Simone Frintrop, Erich Rome, and Henrik I. Christensen. Computational visual attention systems and their cognitive foundations: a survey. ACM Transactions on Applied Perception, 7(1):6:1–6:??, January 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Feixas:2009:UIT

[FSG09] Miquel Feixas, Mateu Sbert, and Francisco González. A unified information-theoretic framework for viewpoint selection and mesh saliency. ACM Transactions on Applied Perception, 6(1):1:1–1:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Flowers:2005:DDS

[FTB05] John H. Flowers, Kimberly D. Turnage, and Dion C. Buhman. Desktop data sonification: Comments on Flowers et al., ICAD 1996. ACM Transactions on Applied Perception, 2 (4):473–476, October 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [FBT05].

Fischer:2025:ICT

[FvdMV25] Joshua David Fischer, Johan van der Merwe, and David Vandenheever. The influence that the complexity of the threedimensional eye model used to generate simulated eye-tracking data has on the gaze estimation errors achieved using the data. ACM Transactions on Applied Perception, 22(1): 2:1–2:??, January 2025. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3660637.

Fan:2014:HPV

[FWN^{+}14] Shaojing Fan, Rangding Wang, Tian-Tsong Ng, Cheston Y.- C. Tan, Jonathan S. Herberg, and Bryan L. Koenig. Human perception of visual realism for photo and computer-generated face images. ACM Transactions on Applied Perception, 11(2): 7:1–7:??, July 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Feijoo-Garcia:2023:PDV

[FWS^{+}23] Pedro Guillermo Feijóo-García, Chase Wrenn, Jacob Stuart, Alexandre Gomes De Siqueira, and Benjamin Lok. Participatory design of virtual humans for mental health support among North American computer science students: Voice, appearance, and the similarityattraction effect. ACM Transactions on Applied Perception, 20(3):11:1–11:??, July 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3613961.

Fang:2020:PAU

[FZL20] Yuchun Fang, Wei Zhang, and Ningjie Liu. On the perception analysis of user feedback for interactive face retrieval. ACM Transactions on Applied Perception, 17(3):10:1– 10:20, November 2020. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3403964.

Grogorick:2018:CUV

[GATM18] Steve Grogorick, Georgia Albuquerque, Jan-Philipp Tauscher, and Marcus Magnor. Comparison of unobtrusive visual guidance methods in an immersive dome environment. ACM Transactions on Applied Perception, 15(4):27:1–27:??, October 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3238303.

Grave:2008:TMO

[GB08] Justine Grave and Roland Bremond. A tone-mapping operator for road visibility experiments. ACM Transactions on Applied Perception, 5(2):12:1– 12:??, May 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Gao:2017:FBQ

[GBA17] Xihe Gao, Stephen Brooks, and Dirk V. Arnold. A featurebased quality metric for tone mapped images. ACM Transactions on Applied Perception, 14 (4):26:1–26:??, September 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Giudice:2010:SLN

[GBLR10] Nicholas A. Giudice, Jonathan Z.

Bakdash, Gordon E. Legge, and Rudrava Roy. Spatial learning and navigation using a virtual verbal display. ACM Transactions on Applied Perception, 7 (1):3:1–3:??, January 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Gilberto:2025:VRA

[GBTB25] Lucas G. Gilberto, Fernando Raúl Bermejo, Fabián C. Tommasini, and Cristian García Bauza. Virtual reality audio game for entertainment and sound localization training. ACM Transactions on Applied Perception, 22(1):4:1–4:??, January 2025. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3676557.

Gamper:2013:SSD

[GDBP13] Hannes Gamper, Christina Dicke, Mark Billinghurst, and Kai Puolamäki. Sound sample detection and numerosity estimation using auditory display. ACM Transactions on Applied Perception, 10(1):4:1–4:??, February 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Gaffary:2013:CAC

[GEMA13] Yoren Gaffary, Victoria Eyharabide, Jean-Claude Martin, and Mehdi Ammi. Clustering approach to characterize haptic expressions of emotions. ACM Transactions on Applied Perception, 10(4):21:1–21:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Giraud:2015:PEP

[GFD^{+}15] Tom Giraud, Florian Focone, Virginie Demulier, Jean Claude Martin, and Brice Isableu. Perception of emotion and personality through full-body movement qualities: a sport coach case study. ACM Transactions on Applied Perception, 13(1): 2:1–2:??, December 2015. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Grimm:2020:ISI

[GGFE20] Cindy Grimm, Mar Gonzalez-Franco, and Elham Ebrahimi. Introduction to the special issue on SAP 2020. ACM Transactions on Applied Perception, 17(4):13e:1–13e:2, December 2020. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3428144.

Glaholt:2020:VIR

[GHS^{+}20] Mackenzie G. Glaholt, Justin G.

Hollands, Grace Sim, Tzvi Spivak, and Beatrice Sacripanti. Visual information requirements for dismounted soldier target acquisition. ACM Transactions on Applied Perception, 17(1):2:1–2:20, March 2020. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3375000.

Grohn:2005:ACG

[GLT05a] Matti Gröhn, Tapio Lokki, and Tapio Takala. Author’s comments on Gröhn, Lokki, and Takala, ICAD 2003. ACM Transactions on Applied Perception, 2(4):571–573, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [GLT05b].

Grohn:2005:CAV

[GLT05b] Matti Gröhn, Tapio Lokki, and Tapio Takala. Comparison of auditory, visual, and audiovisual navigation in a 3D space. ACM Transactions on Applied Perception, 2(4):564–570, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [GLT05a].

Gray:2009:SRC

[GMT09] Rob Gray, Rayka Mohebbi, and Hong Z. Tan. The spatial resolution of crossmodal attention: Implications for the design of multimodal interfaces. ACM Transactions on Applied Perception, 6(1):4:1–4:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Grechkin:2010:HDP

[GNP^{+}10] Timofey Y. Grechkin, Tien Dat Nguyen, Jodie M. Plumert, James F. Cremer, and Joseph K. Kearney. How does presentation method and measurement protocol affect distance estimation in real and virtual environments? ACM Transactions on Applied Perception, 7(4):26:1– 26:??, July 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Gagnon:2021:EDA

[GRM^{+}21] Holly C. Gagnon, Carlos Salas Rosales, Ryan Mileris, Jeanine K. Stefanucci, Sarah H. Creem-Regehr, and Robert E. Bodenheimer. Estimating distances in action space in augmented reality. ACM Transactions on Applied Perception, 18 (2):7:1–7:16, June 2021. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3449067.

Geigel:2013:ISI

[GS13] Joe Geigel and Jeanine Stefanucci. Introduction to special issue SAP 2013. ACM Transactions on Applied Perception, 10(3):12:1–12:??, August 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Gagnon:2023:CPP

[GSCRB23] Holly Gagnon, Jeanine Stefanucci, Sarah Creem-Regehr, and Bobby Bodenheimer. Calibrated passability perception in virtual reality transfers to augmented reality. ACM Transactions on Applied Perception, 20(4):14:1–14:??, October 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3613450.

Gigilashvili:2021:RSS

[GSW^{+}21] Davit Gigilashvili, Weiqi Shi, Zeyu Wang, Marius Pedersen, Jon Yngve Hardeberg, and Holly Rushmeier. The role of subsurface scattering in glossiness perception. ACM Transactions on Applied Perception, 18 (3):10:1–10:26, July 2021. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3458438.

Gugerty:2004:ESA

[GTAE04] Leo Gugerty, Richard A.

Tyrrell, Thomas R. Aten, and K. Andy Edmonds. The effects of subpixel addressing on users’ performance and preferences during reading-related tasks. ACM Transactions on Applied Perception, 1(2):81– 101, October 2004. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Guo:2017:SOV

[GVC^{+}17] Jinjiang Guo, Vincent Vidal, Irene Cheng, Anup Basu, Atilla Baskurt, and Guillaume Lavoue. Subjective and objective visual quality assessment of textured 3D meshes. ACM Transactions on Applied Perception, 14(2):11:1–11:??, February 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Gao:2022:ESM

[GWFL22] Zihan Gao, Huiqiang Wang, Guangsheng Feng, and Hongwu Lv. Exploring sonification mapping strategies for spatial auditory guidance in immersive virtual environments. ACM Transactions on Applied Perception, 19(3):9:1–9:??, July 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3528171.

Homayouni:2017:RIS

[HAHG17] Maryam Homayouni, Payman Aflaki, Miska M. Hannuksela, and Moncef Gabbouj. Rowinterleaved sampling for depthenhanced 3D video coding for polarized displays. ACM Transactions on Applied Perception, 14(3):15:1–15:??, July 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Holmes:2016:AII

[HBF16] Olivia Holmes, Martin S.

Banks, and Hany Farid. Assessing and improving the identification of computer-generated portraits. ACM Transactions on Applied Perception, 13(2): 7:1–7:??, March 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hooge:2021:EAF

[HBK^{+}21] Kimberley D. Orsten Hooge, Asal Baragchizadeh, Thomas P. Karnowski, David S. Bolme, Regina Ferrell, Parisa R. Jesudasen, Carlos D. Castillo, and Alice J. O’Toole. Evaluating automated face identitymasking methods with human perception and a deep convolutional neural network. ACM Transactions on Applied Perception, 18(1):3:1–3:20, January 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3422988.

Hecher:2014:CPS

[HBM^{+}14] Michael Hecher, Matthias Bernhard, Oliver Mattausch, Daniel Scherzer, and Michael Wimmer. A comparative perceptual study of soft-shadow algorithms. ACM Transactions on Applied Perception, 11(2):5:1– 5:??, July 2014. CODEN ????

ISSN 1544-3558 (print), 1544-3965 (electronic).

Hodgson:2011:RWE

[HBW11] Eric Hodgson, Eric Bachmann, and David Waller. Redirected walking to explore virtual environments: Assessing the potential for spatial interference. ACM Transactions on Applied Perception, 8(4):22:1– 22:??, November 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hoyet:2019:ISI

[HC19] Ludovic Hoyet and Douglas W. Cunningham. Introduction to the special issue on SAP 2019. ACM Transactions on Applied Perception, 16(3):13:1– 13:??, September 2019. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3355996.

Hyde:2016:EAC

[HCKH16] Jennifer Hyde, Elizabeth J. Carter, Sara Kiesler, and Jessica K. Hodgins. Evaluating animated characters: Facial motion magnitude influences personality perceptions. ACM Transactions on Applied Perception, 13(2):8:1–8:??, March 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hasic:2010:PGH

[HCS10] Jasminka Hasic, Alan Chalmers, and Elena Sikudova. Perceptually guided high-fidelity rendering exploiting movement bias in visual attention. ACM Transactions on Applied Perception, 8 (1):6:1–6:??, October 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Howard:2023:GDP

[HDF^{+}23] Thomas Howard, Karina Driller, William Frier, Claudio Pacchierotti, Maud Marchal, and Jessica Hartcher-O’Brien. Gap detection in pairs of ultrasound mid-air vibrotactile stimuli. ACM Transactions on Applied Perception, 20(1):5:1– 5:??, January 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3570904.

Hover:2010:UBE

[HDH10] Raphael Höver, Massimiliano Di Luca, and Matthias Harders. User-based evaluation of datadriven haptic rendering. ACM Transactions on Applied Perception, 8(1):7:1–7:??, October 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Harders:2005:ESI

[HE05] Matthias Harders and Marc Ernst. EuroHaptics special issue editorial. ACM Transactions on Applied Perception, 2 (2):69–70, April 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hattenberger:2009:PIG

[HFJS09] Timothy J. Hattenberger, Mark D. Fairchild, Garrett M. Johnson, and Carl Salvaggio. A psychophysical investigation of global illumination algorithms used in augmented reality. ACM Transactions on Applied Perception, 6(1):2:1–2:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hashiura:2024:TWR

[HHB^{+}24] Kenta Hashiura, Takayoshi Hagiwara, Giulia Barbareschi, Sohei Wakisaka, and Kouta Minamizawa. “Together with who?” Recognizing partners during collaborative avatar manipulation. ACM Transactions on Applied Perception, 21 (4):17:1–17:??, October 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3698237.

Hassaine:2010:IPP

[HHL10] Djamel Hassaine, Nicolas S. Holliman, and Simon P. Liversedge. Investigating the performance of path-searching tasks in depth on multiview displays. ACM Transactions on Applied Perception, 8(1): 8:1–8:??, October 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hadnett-Hunter:2019:ETV

[HHNOP19] Jacob Hadnett-Hunter, George Nicolaou, Eamonn O’Neill, and Michael Proulx. The effect of task on visual attention in interactive virtual environments. ACM Transactions on Applied Perception, 16(3): 17:1–17:??, September 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3352763.

Howlett:2005:PES

[HHO05] Sarah Howlett, John Hamill, and Carol O’Sullivan. Predicting and evaluating saliency for simplified polygonal models. ACM Transactions on Applied Perception, 2(3):286–308, July 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ho:2007:DET

[HJ07] Hsin-Ni Ho and Lynette A.

Jones. Development and evaluation of a thermal display for material identification and discrimination. ACM Transactions on Applied Perception, 4 (2):13:1–13:??, July 2007. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hodgins:2010:SAA

[HJO^{+}10] Jessica Hodgins, Sophie Jörg, Carol O’Sullivan, Sang Il Park, and Moshe Mahler. The saliency of anomalies in animated human characters. ACM Transactions on Applied Perception, 7(4):22:1–22:??, July 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Harper:2009:TDV

[HMS09] Simon Harper, Eleni Michailidou, and Robert Stevens. Toward a definition of visual complexity as an implicit measure of cognitive load. ACM Transactions on Applied Perception, 6(2):10:1–10:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hinde:2022:IPH

[HNT^{+}22] Stephen J. Hinde, Katy C. Noland, Graham A. Thomas, David R. Bull, and Iain D. Gilchrist. On the immersive properties of high dynamic range video. ACM Transactions on Applied Perception, 19 (2):7:1–7:15, April 2022. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3524692.

Hakala:2015:DAC

[HOH15] Jussi H. Hakala, Pirkko Oittinen, and Jukka P. Häkkinen. Depth artifacts caused by spatial interlacing in stereoscopic 3D displays. ACM Transactions on Applied Perception, 12(1): 3:1–3:??, March 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hermann:2005:CSH

[HR05a] Thomas Hermann and Helge Ritter. Crystallization sonification of high-dimensional datasets. ACM Transactions on Applied Perception, 2(4): 550–558, October 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [HR05b].

Hermann:2005:MBS

[HR05b] Thomas Hermann and Helge Ritter. Model-based sonification revisited—authors’ comments on Hermann and Ritter, ICAD 2002. ACM Transactions on Applied Perception, 2 (4):559–563, October 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [HR05a].

Healey:2012:LRV

[HS12] Christopher G. Healey and Amit P. Sawant. On the limits of resolution and visual angle in visualization. ACM Transactions on Applied Perception, 9(4):20:1–20:??, October 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Huckauf:2011:OSG

[HU11] Anke Huckauf and Mario H.

Urbina. Object selection in gaze controlled systems: What you don’t look at is what you get. ACM Transactions on Applied Perception, 8(2):13:1– 13:??, January 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Holten:2006:PBS

[HVM06] Danny Holten, Jarke J. Van Wijk, and Jean-Bernard Martens.

A perceptually based spectral model for isotropic textures. ACM Transactions on Applied Perception, 3(4):376–398, October 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Hoh:2023:SCS

[HZD23] Weng Khuan Hoh, Fang-Lue Zhang, and Neil A. Dodgson. Salient-centeredness and saliency size in computational aesthetics. ACM Transactions on Applied Perception, 20(2): 8:1–8:??, April 2023. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3588317.

Interrante:2015:WMN

[IG15] Victoria Interrante and Diego Gutierrez. Welcome message from the new Editors-in-Chief. ACM Transactions on Applied Perception, 12(3):8:1–8:??, July 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Interrante:2006:GE

[Int06] Victoria Interrante. Guest editorial. ACM Transactions on Applied Perception, 3(3):153– 154, July 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ito:2019:TTD

[IOYK19] Ken Ito, Shogo Okamoto, Yoji Yamada, and Hiroyuki Kajimoto. Tactile texture display with vibrotactile and electrostatic friction stimuli mixed at appropriate ratio presents better roughness textures. ACM Transactions on Applied Perception, 16(4):20:1– 20:??, September 2019. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3340961.

Jain:2016:MCP

[JAA^{+}16] Eakta Jain, Lisa Anthony, Aishat Aloba, Amanda Castonguay, Isabella Cuba, Alex Shaw, and Julia Woodward. Is the motion of a child perceivably different from the motion of an adult? ACM Transactions on Applied Perception, 13(4): 22:1–22:??, July 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Jewst:2025:VTM

[JCM^{+}25] Anika Jewst, Susana Castillo, Marcus Magnor, Martin Eisemann, and Dagmar Meyer. Virtual and traditional memory palaces in recall with ADHD. ACM Transactions on Applied Perception, 22(4):15:1–15:15, October 2025. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Jorg:2018:PAE

[JDKN18] Sophie Jörg, Andrew Duchowski, Krzysztof Krejtz, and Anna Niedzielska. Perceptual adjustment of eyeball rotation and pupil size jitter for virtual characters. ACM Transactions on Applied Perception, 15(4):24:1– 24:??, October 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Jagnow:2008:EMA

[JDR08] Robert Jagnow, Julie Dorsey, and Holly Rushmeier. Evaluation of methods for approximating shapes used to synthesize 3D solid textures. ACM Transactions on Applied Perception, 4(4):5:1–5:??, January 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Jones:2017:VFV

[JKB17] J. Adam Jones, David M. Krum, and Mark T. Bolas. Vertical field-of-view extension and walking characteristics in head-worn virtual environments. ACM Transactions on Applied Perception, 14(2):9:1– 9:??, February 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Janeh:2017:WVR

[JLS^{+}17] Omar Janeh, Eike Langbehn, Frank Steinicke, Gerd Bruder, Alessandro Gulberti, and Monika Poetter-Nerger. Walking in virtual reality: Effects of manipulated visual selfmotion on walking biomechanics. ACM Transactions on Applied Perception, 14(2):12:1– 12:??, February 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Jain:2021:ISI

[JO21] Eakta Jain and Anne-Hélène Olivier. Introduction to the special issue on SAP 2021. ACM Transactions on Applied Perception, 18(4):18:1–18:2, October 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3486577.

Jiang:2018:ATJ

[JOY^{+}18] Yuanyuan Jiang, Elizabeth E.

O’neal, Junghum Paul Yon, Luke Franzen, Pooya Rahimian, Jodie M. Plumert, and Joseph K. Kearney. Acting together: Joint pedestrian road crossing in an immersive virtual environment. ACM Transactions on Applied Perception, 15(2):8:1– 8:??, April 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Jiang:2021:CRC

[JOZ^{+}21] Yuanyuan Jiang, Elizabeth E.

O’Neal, Shiwen Zhou, Jodie M. Plumert, and Joseph K. Kearney. Crossing roads with a computer-generated agent: Persistent effects on perceptionaction tuning. ACM Transactions on Applied Perception, 18(1):4:1–4:16, January 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3431923.

Jun:2015:BFU

[JSCR^{+}15] Eunice Jun, Jeanine K. Stefanucci, Sarah H. Creem-Regehr, Michael N. Geuss, and William B. Thompson. Big foot: Using the size of a virtual foot to scale gap width. ACM Transactions on Applied Perception, 12(4):16:1–16:??, September 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Jimenez:2009:SSP

[JSG09] Jorge Jimenez, Veronica Sundstedt, and Diego Gutierrez. Screen-space perceptual rendering of human skin. ACM Transactions on Applied Perception, 6(4):23:1–23:??, September 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Jay:2008:UHC

[JSHG08] Caroline Jay, Robert Stevens, Roger Hubbold, and Mashhuda Glencross. Using haptic cues to aid nonvisual structure recognition. ACM Transactions on Applied Perception, 5(2):8:1– 8:??, May 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Jacobs:2019:KIS

[JWA19] Jochen Jacobs, Xi Wang, and Marc Alexa. Keep it simple: Depth-based dynamic adjustment of rendering for headmounted displays decreases visual comfort. ACM Transactions on Applied Perception, 16 (3):16:1–16:??, September 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3353902.

Jerald:2012:SMT

[JWB12] Jason Jerald, Mary Whitton, and Frederick P. Brooks, Jr. Scene-motion thresholds during head yaw for immersive virtual environments. ACM Transactions on Applied Perception, 9 (1):4:1–4:??, March 2012. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kawabe:2019:SBI

[Kaw19] Takahiro Kawabe. Shadowbased illusion of depth and transparency in printed images. ACM Transactions on Applied Perception, 16(2):10:1– 10:??, August 2019. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3342350.

Kelly:2006:SMS

[KBL^{+}06] Jonathan W. Kelly, Andrew C.

Beall, Jack M. Loomis, Roy S. Smith, and Kristen L. Macuga. Simultaneous measurement of steering performance and perceived heading on a curving path. ACM Transactions on Applied Perception, 3(2):83–94, April 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kaaresoja:2014:TTP

[KBL14] Topi Kaaresoja, Stephen Brewster, and Vuokko Lantz. Towards the temporally perfect virtual button: Touchfeedback simultaneity and perceived quality in mobile touchscreen press interactions. ACM Transactions on Applied Perception, 11(2):9:1–9:??, July 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kelly:2013:SPV

[KBP^{+}13] Jonathan W. Kelly, Melissa Burton, Brice Pollock, Eduardo Rubio, Michael Curtis, Julio De La Cruz, Stephen Gilbert, and Eliot Winer. Space perception in virtual environments: Displacement from the center of projection causes less distortion than predicted by cuebased models. ACM Transactions on Applied Perception, 10(4):18:1–18:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kelly:2018:CTM

[KCK^{+}18] Jonathan W. Kelly, Lucia A. Cherep, Brenna Klesel, Zachary D. Siegel, and Seth George. Comparison of two methods for improving distance perception in virtual reality. ACM Transactions on Applied Perception, 15 (2):11:1–11:??, April 2018. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kuhl:2008:RRL

[KCRT08] Scott A. Kuhl, Sarah H.

Creem-Regehr, and William B. Thompson. Recalibration of rotational locomotion in immersive virtual environments. ACM Transactions on Applied Perception, 5(3):17:1–17:??, August 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kelly:2017:PSH

[KCS17] Jonathan W. Kelly, Lucia A.

Cherep, and Zachary D. Siegel. Perceived space in the HTC vive. ACM Transactions on Applied Perception, 15(1):2:1–2:??, November 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Koulieris:2015:AHL

[KDCM15] George Alex Koulieris, George Drettakis, Douglas Cunningham, and Katerina Mania. An automated high-level saliency predictor for smart game balancing. ACM Transactions on Applied Perception, 11(4):17:1– 17:??, January 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Krejtz:2016:DAF

[KDK^{+}16] Krzysztof Krejtz, Andrew Duchowski, Izabela Krejtz, Agnieszka Szarkowska, and Agata Kopacz. Discerning ambient/ focal attention with coefficient K. ACM Transactions on Applied Perception, 13(3):11:1– 11:??, May 2016. CODEN ????

ISSN 1544-3558 (print), 1544-3965 (electronic).

Krejtz:2015:GTE

[KDS^{+}15] Krzysztof Krejtz, Andrew Duchowski, Tomasz Szmidt, Izabela Krejtz, Fernando González Perilli, Ana Pires, Anna Vilaro, and Natalia Villalobos. Gaze transition entropy. ACM Transactions on Applied Perception, 13(1):4:1–4:??, December 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kawabe:2016:DLP

[KFSN16] Takahiro Kawabe, Taiki Fukiage, Masataka Sawayama, and Shin’ya Nishida. Deformation lamps: a projection technique to make static objects perceptually dynamic. ACM Transactions on Applied Perception, 13(2):10:1–10:??, March 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Katsunuma:2017:FAC

[KHH17] Takafumi Katsunuma, Keita Hirai, and Takahiko Horiuchi. Fabric appearance control system for example-based interactive texture and color design. ACM Transactions on Applied Perception, 14(3):16:1– 16:??, July 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Komogortsev:2013:LOP

[KHJK13] Oleg Komogortsev, Corey Holland, Sampath Jayarathna, and Alex Karpov. 2D linear oculomotor plant mathematical model: Verification and biometric applications. ACM Transactions on Applied Perception, 10(4):27:1–27:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Komogortsev:2015:BOP

[KHKP15] Oleg Komogortsev, Corey Holland, Alex Karpov, and Larry R. Price. Biometrics via oculomotor plant characteristics: Impact of parameters in oculomotor plant model. ACM Transactions on Applied Perception, 11(4):20:1–20:??, January 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kiiski:2015:SHS

[KHW^{+}15] Hanni Kiiski, Ludovic Hoyet, Andy T. Woods, Carol O’Sullivan, and Fiona N. Newell. Strutting hero, sneaking villain: Utilizing body motion cues to predict the intentions of others. ACM Transactions on Applied Perception, 13(1):1:1–1:??, December 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kelly:2019:VSB

[KKC19] Jonathan W. Kelly, Brenna C.

Klesel, and Lucia A. Cherep. Visual stabilization of balance in virtual reality using the HTC Vive. ACM Transactions on Applied Perception, 16(2):8:1–8:??, August 2019.

CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3313902.

Klatzky:2006:PRR

[KL06] Roberta L. Klatzky and Susan J. Lederman. The perceived roughness of resistive virtual textures: I. Rendering by a force-feedback mouse. ACM Transactions on Applied Perception, 3(1):1–14, January 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kim:2024:ERE

[KLL24] Aelee Kim, Jeong-Eun Lee, and Kyoung-Min Lee. Exploring the relative effects of body position and locomotion method on presence and cybersickness when navigating a virtual environment. ACM Transactions on Applied Perception, 21(1): 3:1–3:??, January 2024. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3627706.

Kneusel:2017:IHM

[KM17] Ronald T. Kneusel and Michael C. Mozer. Improving humanmachine cooperative visual search with soft highlighting. ACM Transactions on Applied Perception, 15(1):3:1–3:??, November 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kenny:2019:PEI

[KMH^{+}19] Sophie Kenny, Naureen Mahmood, Claire Honda, Michael J. Black, and Nikolaus F. Troje. Perceptual effects of inconsistency in human animations. ACM Transactions on Applied Perception, 16(1):2:1– 2:??, February 2019. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3301411.

Kyto:2013:IRD

[KMHO13] Mikko Kytö, Aleksi Mäkinen, Jukka Häkkinen, and Pirkko Oittinen. Improving relative depth judgments in augmented reality with auxiliary augmentations. ACM Transactions on Applied Perception, 10(1):6:1– 6:??, February 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kuhl:2015:ISI

[KMS15] Scott Kuhl, Rafal Mantiuk, and Betsy Sanders. Introduction to special issue SAP 2015. ACM Transactions on Applied Perception, 12(4):13:1– 13:??, September 2015. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kim:2022:PCA

[KNL22] Hye Ji Kim, Michael Neff, and Sung-Hee Lee. The perceptual consistency and association of the LMA effort elements. ACM Transactions on Applied Perception, 19(1): 1:1–1:17, January 2022. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3473041.

Kim:2010:PGG

[KPAA10] Juno Kim, Stephen A. Palmisano, April Ash, and Robert S. Allison. Pilot gaze and glideslope control. ACM Transactions on Applied Perception, 7(3):18:1– 18:??, June 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kartashova:2019:LSP

[KPDP19] Tatiana Kartashova, Susan F. Te Pas, Huib De Ridder, and Sylvia C. Pont. Light shapes: Perception-based visualizations of the global light transport. ACM Transactions on Applied Perception, 16(1):4:1–4:??, February 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3232851.

Kuzovkin:2019:CPA

[KPL^{+}19] Dmitry Kuzovkin, Tania Pouli, Olivier Le Meur, Rémi Cozot, Jonathan Kervec, and Kadi Bouatouch. Context in photo albums: Understanding and modeling user behavior in clustering and selection. ACM Transactions on Applied Perception, 16(2):11:1–11:??, August 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3333612.

Kohm:2022:SHO

[KPR22] Kristopher Kohm, John Porter, and Andrew Robb. Sensitivity to hand offsets and related behavior in virtual environments over time. ACM Transactions on Applied Perception, 19 (4):17:1–17:??, October 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561055.

Kjellin:2010:EVS

[KPSL10] Andreas Kjellin, Lars Winkler Pettersson, Stefan Seipel, and Mats Lind. Evaluating 2D and 3D visualizations of spatiotemporal information. ACM Transactions on Applied Perception, 7(3):19:1–19:??, June 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kellnhofer:2014:SDN

[KRV^{+}14] Petr Kellnhofer, Tobias Ritschel, Peter Vangorp, Karol Myszkowski, and Hans-Peter Seidel. Stereo day-for-night: Retargeting disparity for scotopic vision. ACM Transactions on Applied Perception, 11(3):15:1–15:??, August 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kastanis:2012:RLU

[KS12] Iason Kastanis and Mel Slater. Reinforcement learning utilizes proxemics: An avatar learns to manipulate the position of people in immersive virtual reality. ACM Transactions on Applied Perception, 9(1):3:1–3:??, March 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kokkinara:2015:EVC

[KSLM15] Elena Kokkinara, Mel Slater, and Joan López-Moliner. The effects of visuomotor calibration to the perceived space and body, through embodiment in immersive virtual reality. ACM Transactions on Applied Perception, 13(1):3:1–3:??, December 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kikuuwe:2005:EHD

[KSM^{+}05] Ryo Kikuuwe, Akihito Sano, Hiromi Mochiyama, Naoyuki Takesue, and Hideo Fujimoto. Enhancing haptic detection of surface undulation. ACM Transactions on Applied Perception, 2(1):46–67, January 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kasahara:2021:SUV

[KT21] Shunichi Kasahara and Kazuma Takada. Stealth updates of visual information by leveraging change blindness and computational visual morphing. ACM Transactions on Applied Perception, 18(4):23:1– 23:17, October 2021. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3486581.

Kuhl:2009:HCE

[KTCR09] Scott A. Kuhl, William B.

Thompson, and Sarah H. Creem-Regehr. HMD calibration and its effects on distance judgments. ACM Transactions on Applied Perception, 6(3): 19:1–19:??, August 2009. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Knopp:2019:PPN

[KVDE19] Benjamin Knopp, Dmytro Velychko, Johannes Dreibrodt, and Dominik Endres. Predicting perceived naturalness of human animations based on generative movement primitive models. ACM Transactions on Applied Perception, 16 (3):15:1–15:??, September 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3355401.

Koenderink:2016:CPI

[KVE16] Jan Koenderink, Andrea Van Doorn, and Vebjørn Ekroll. Color picking: The initial 20s. ACM Transactions on Applied Perception, 13(3):13:1– 13:??, May 2016. CODEN ????

ISSN 1544-3558 (print), 1544-3965 (electronic).

Kim:2010:MSH

[KVJG10] Youngmin Kim, Amitabh Varshney, David W. Jacobs, and François Guimbretière. Mesh saliency and human eye fixations. ACM Transactions on Applied Perception, 7(2): 12:1–12:??, February 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kramer:2005:SSM

[KW05] Gregory Kramer and Bruce N. Walker. Sound science: Marking ten international conferences on auditory display. ACM Transactions on Applied Perception, 2(4):383–388, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Khan:2009:CPE

[KWI09] Masood Mehmood Khan, Robert D. Ward, and Michael Ingleby. Classifying pretended and evoked facial expressions of positive and negative affective states using infrared measurement of skin temperature. ACM Transactions on Applied Perception, 6(1):6:1–6:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Klatzky:2008:EAR

[KWSS08] Roberta L. Klatzky, Bing Wu, Damion Shelton, and George Stetten. Effectiveness of augmented-reality visualization versus cognitive mediation for learning actions in near space. ACM Transactions on Applied Perception, 5(1):1:1–1:??, January 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Kuang:2007:EHR

[KYL^{+}07] Jiangtao Kuang, Hiroshi Yamaguchi, Changmeng Liu, Garrett M. Johnson, and Mark D. Fairchild. Evaluating HDR rendering algorithms. ACM Transactions on Applied Perception, 4(2):9:1–9:??, July 2007. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Li:2009:PIM

[LAE09] Li Li, Bernard D. Adelstein, and Stephen R. Ellis. Perception of image motion during head movement. ACM Transactions on Applied Perception, 6(1):5:1–5:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lavoue:2009:LRM

[Lav09]  Guillaume Lavoué.  A lo-

cal roughness measure for 3D meshes and its application to visual masking. ACM Transactions on Applied Perception, 5(4):21:1–21:??, January 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lecuyer:2008:SMS

[LBT08] Anatole Lécuyer, Jean-Marie Burkhardt, and Chee-Hian Tan.

A study of the modification of the speed and size of the cursor for simulating pseudohaptic bumps and holes. ACM Transactions on Applied Perception, 5(3):14:1–14:??, August 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lichtenstein:2007:FCI

[LBWP07] Lee Lichtenstein, James Barabas, Russell L. Woods, and Eli Peli. A feedback-controlled interface for treadmill locomotion in virtual environments. ACM Transactions on Applied Perception, 4(1):??, January 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Legde:2015:MAP

[LCC15] Katharina Legde, Susana Castillo, and Douglas W. Cunningham. Multimodal affect: Perceptually evaluating an affective talking head. ACM Transactions on Applied Perception, 12(4):17:1– 17:??, September 2015. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lau:2018:HPS

[LDDR18] Manfred Lau, Kapil Dev, Julie Dorsey, and Holly Rushmeier. A human-perceived softness measure of virtual 3D objects. ACM Transactions on Applied Perception, 15(3):19:1–19:??, August 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Leroy:2012:RTA

[LFM12] Laure Leroy, Philippe Fuchs, and Guillaume Moreau. Realtime adaptive blur for reducing eye strain in stereoscopic displays. ACM Transactions on Applied Perception, 9(2): 9:1–9:??, June 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

LoPrete:2025:PEV

[LGS25] Anthony LoPrete, Alexander Gokan, and Arthur G. Shapiro. The Pulfrich effect in virtual reality. ACM Transactions on Applied Perception, 22(3):12:1– 12:??, July 2025. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lederman:2006:PRR

[LKTH06] Susan J. Lederman, Roberta L.

Klatzky, Christine Tong, and Cheryl Hamilton. The perceived roughness of resistive virtual textures: II. effects of varying viscosity with a forcefeedback device. ACM Transactions on Applied Perception, 3(1):15–30, January 2006. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Leyrer:2015:EHM

[LLBM15] Markus Leyrer, Sally A. Linkenauger, Heinrich H. Bülthoff, and Betty J. Mohler. Eye height manipulations: a possible solution to reduce underestimation of egocentric distances in head-mounted displays. ACM Transactions on Applied Perception, 12(1):1:1–1:??, March 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lu:2010:VCE

[LME10] Aidong Lu, Ross Maciejewski, and David S. Ebert. Volume composition and evaluation using eye-tracking data. ACM Transactions on Applied Perception, 7(1):4:1–4:??, January 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Liu:2022:MVC

[LMM^{+}22] Wanyu Liu, Michelle Agnes Magalhaes, Wendy E. Mackay, Michel Beaudouin-Lafon, and Frédéric Bevilacqua. Motor variability in complex gesture learning: Effects of movement sonification and musical background. ACM Transactions on Applied Perception, 19(1): 2:1–2:21, January 2022. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3482967.

Lau:2024:ISI

[LP24] Manfred Lau and Julien Pettré. Introduction to the special issue on SAP 2024. ACM Transactions on Applied Perception, 21(4):13:1–13:??, October 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3702963.

Laitinen:2012:PTF

[LPEP12] Mikko-Ville Laitinen, Tapani Pihlajamäki, Cumhur Erkut, and Ville Pulkki. Parametric time-frequency representation of spatial sound in virtual worlds. ACM Transactions on Applied Perception, 9 (2):8:1–8:??, June 2012. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Levesque:2005:DVB

[LPHL05] Vincent Lévesque, Jérôme Pasquero, Vincent Hayward, and Maryse Legault. Display of virtual Braille dots by lateral skin deformation: feasibility study. ACM Transactions on Applied Perception, 2(2):132–149, April 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Li:2009:NEF

[LPO09] Yanfang Li, Volkan Patoglu, and Marcia K. O’Malley. Negative efficacy of fixed gain error reducing shared control for training in virtual environments. ACM Transactions on Applied Perception, 6(1):3:1– 3:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Langer:2006:PLM

[LPR06] Michael S. Langer, Javeen Pereira, and Dipinder Rekhi. Perceptual limits on 2D motionfield visualization. ACM Transactions on Applied Perception, 3(3):179–193, July 2006. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lovell:2006:EMC

[LPT^{+}06] P. George Lovell, C. Alejandro Párraga, Tom Troscianko, Caterina Ripamonti, and David J. Tolhurst. Evaluation of a multiscale color model for visual difference prediction. ACM Transactions on Applied Perception, 3(3):155–178, July 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lin:2015:AJH

[LRB15] Qiufeng Lin, John Rieser, and Bobby Bodenheimer. Affordance judgments in HMD-based virtual environments: Stepping over a pole and stepping off a ledge. ACM Transactions on Applied Perception, 12(2):6:1– 6:??, April 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lylykangas:2013:IVS

[LSRR13] Jani Lylykangas, Veikko Surakka, Jussi Rantala, and Roope Raisamo. Intuitiveness of vibrotactile speed regulation cues. ACM Transactions on Applied Perception, 10(4):24:1–24:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Llobera:2010:PMD

[LSRS10] Joan Llobera, Bernhard Spanlang, Giulio Ruffini, and Mel Slater. Proxemics with multiple dynamic characters in an immersive virtual environment. ACM Transactions on Applied Perception, 8(1):3:1–3:??, October 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lin:2020:HPS

[LVV^{+}20] Yun-Xuan Lin, Rohith Venkatakrishnan, Roshan Venkatakrishnan, Elham Ebrahimi, Wen-Chieh Lin, and Sabarish V. Babu. How the presence and size of static peripheral blur affects cybersickness in virtual reality. ACM Transactions on Applied Perception, 17(4):16:1– 16:18, December 2020. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3419984.

Li:2018:EPV

[LWK18] Bochao Li, James Walker, and Scott A. Kuhl. The effects of peripheral vision and light stimulation on distance judgments through HMDs. ACM Transactions on Applied Perception, 15 (2):12:1–12:??, April 2018. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Li:2010:SCS

[LXXB10] Bing Li, Weihua Xiong, De Xu, and Hong Bao. A supervised combination strategy for illumination chromaticity estimation. ACM Transactions on Applied Perception, 8(1):5:1–5:??, October 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lin:2013:SMA

[LZG^{+}13] Kai-Hsiang Lin, Xiaodan Zhuang, Camille Goudeseune, Sarah King, Mark Hasegawa-Johnson, and Thomas S. Huang. Saliencymaximized audio visualization and efficient audio-visual browsing for faster-than-real-time human acoustic event detection. ACM Transactions on Applied Perception, 10(4):26:1– 26:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Li:2017:HDS

[LZL17] Yi-Na Li, Kang Zhang, and Dong-Jin Li. How dimensional and semantic attributes of visual sign influence relative value estimation. ACM Transactions on Applied Perception, 14(3): 18:1–18:??, July 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Lyu:2018:NIM

[LZL^{+}18] Yongqiang Lyu, Xiao Zhang, Xiaomin Luo, Ziyue Hu, Jingyu Zhang, and Yuanchun Shi. Noninvasive measurement of cognitive load and stress based on the reflected stress-induced vascular response index. ACM Transactions on Applied Perception, 15(3):17:1–17:??, August 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Merer:2013:PCM

[MAYKM13] Adrien Merer, Mitsuko Aramaki, Sølvi Ystad, and Richard Kronland-Martinet. Perceptual characterization of motion evoked by sounds for synthesis control purposes. ACM Transactions on Applied Perception, 10(1):1:1–1:??, February 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

McGookin:2004:UCE

[MB04] David K. McGookin and Stephen A. Brewster. Understanding concurrent earcons: Applying auditory scene analysis principles to concurrent earcon recognition. ACM Transactions on Applied Perception, 1(2):130–155, October 2004. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mania:2010:EAS

[MB10] Katerina Mania and Martin S.

Banks. Editorial – APGV 2010 special issue. ACM Transactions on Applied Perception, 7 (4):21:1–21:??, July 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mania:2010:CTS

[MBCW10] Katerina Mania, Shahrul Badariah, Matthew Coxon, and Phil Watten. Cognitive transfer of spatial awareness states from immersive virtual environments to reality. ACM Transactions on Applied Perception, 7(2):9:1– 9:??, February 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

McNamara:2009:STP

[MBG09] Ann McNamara, Reynold Bailey, and Cindy Grimm. Search task performance using subtle gaze direction with the presence of distractions. ACM Transactions on Applied Perception, 6(3):17:1–17:??, August 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Miner:2005:UWS

[MC05] Nadine E. Miner and Thomas P. Caudell. Using wavelets to synthesize stochastic-based sounds for immersive virtual environments. ACM Transactions on Applied Perception, 2(4):521– 528, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [MPC05].

McNamara:2006:EVA

[McN06] Ann McNamara. Exploring visual and automatic measures of perceptual fidelity in real and simulated imagery. ACM Transactions on Applied Perception, 3(3):217–238, July 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Messing:2005:DPV

[MD05] Ross Messing and Frank H. Durgin. Distance perception and the visual horizon in headmounted displays. ACM Transactions on Applied Perception, 2(3):234–250, July 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mion:2010:POA

[MDR10] Luca Mion, Giovanni De Poli, and Ennio Rapanà. Perceptual organization of affective and sensorial expressive intentions in music performance. ACM Transactions on Applied Perception, 7(2):14:1–14:??, February 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Murphy:2009:HIM

[MDT09] Hunter A. Murphy, Andrew T.

Duchowski, and Richard A. Tyrrell. Hybrid image/modelbased gaze-contingent rendering. ACM Transactions on Applied Perception, 5(4):22:1– 22:??, January 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

McDonnell:2009:TBS

[MEDO09] Rachel McDonnell, Cathy Ennis, Simon Dobbyn, and Carol O’Sullivan. Talking bodies: Sensitivity to desynchronization of conversations. ACM Transactions on Applied Perception, 6(4):22:1– 22:??, September 2009. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mustafa:2012:STE

[MGM12] Maryam Mustafa, Stefan Guthe, and Marcus Magnor. Singletrial EEG classification of artifacts in videos. ACM Transactions on Applied Perception, 9(3):12:1–12:??, July 2012. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Marentakis:2016:TID

[MGM16] Georgios Marentakis, Cathryn Griffiths, and Stephen Mcadams. Top-down influences in the detection of spatial displacement in a musical scene. ACM Transactions on Applied Perception, 14(1):3:1–3:??, August 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Meyer:2016:SVC

[MGVM16] Benjamin Meyer, Steve Grogorick, Mark Vollrath, and Marcus Magnor. Simulating visual contrast reduction during nighttime glare situations on conventional displays. ACM Transactions on Applied Perception, 14(1):4:1–4:??, August 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mulot:2023:IPM

[MHPM23] Lendy Mulot, Thomas Howard, Claudio Pacchierotti, and Maud Marchal. Improving the perception of mid-air tactile shapes with spatio-temporally-modulated tactile pointers. ACM Transactions on Applied Perception, 20(4):13:1–13:??, October 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3611388.

Majumder:2007:PBC

[MI07] Aditi Majumder and Sandy Irani. Perception-based contrast enhancement of images. ACM Transactions on Applied Perception, 4(3):17:1– 17:??, November 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

McDonnell:2009:EEM

[MJH^{+}09] Rachel McDonnell, Sophie Jörg, Jessica K. Hodgins, Fiona Newell, and Carol O’Sullivan. Evaluating the effect of motion and body shape on the perceived sex of virtual characters. ACM Transactions on Applied Perception, 5(4):20:1– 20:??, January 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

McDonnell:2009:IRB

[MJM^{+}09] Rachel McDonnell, Sophie Jörg, Joanna McHugh, Fiona N. Newell, and Carol O’Sullivan. Investigating the role of body shape on the perception of emotion. ACM Transactions on Applied Perception, 6(3):14:1– 14:??, August 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mallick:2024:IOR

[MJP^{+}24] Snipta Mallick, Géraldine Jeckeln, Connor J. Parde, Carlos D. Castillo, and Alice J. O’Toole. The influence of the other-race effect on susceptibility to face morphing attacks. ACM Transactions on Applied Perception, 21(1):2:1–2:??, January 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3618113.

Marston:2006:ESD

[MLK^{+}06] James R. Marston, Jack M. Loomis, Roberta L. Klatzky, Reginald G. Golledge, and Ethan L. Smith. Evaluation of spatial displays for navigation without sight. ACM Transactions on Applied Perception, 3 (2):110–124, April 2006. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Marentakis:2013:PIG

[MM13] G. Marentakis and S. Mcadams. Perceptual impact of gesture control of spatialization. ACM Transactions on Applied Perception, 10(4):22:1–22:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mantiuk:2006:PFC

[MMS06] Rafal Mantiuk, Karol Myszkowski, and Hans-Peter Seidel. A perceptual framework for contrast processing of high dynamic range images. ACM Transactions on Applied Perception, 3 (3):286–308, July 2006. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mccrae:2013:SPP

[MMS13] James Mccrae, Niloy J. Mitra, and Karan Singh. Surface perception of planar abstractions. ACM Transactions on Applied Perception, 10(3):14:1– 14:??, August 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Moscoso:2015:ASI

[MMSO15] Claudia Moscoso, Barbara Matusiak, U. Peter Svensson, and Krzysztof Orleanski. Analysis of stereoscopic images as a new method for daylighting studies. ACM Transactions on Applied Perception, 11(4):21:1– 21:??, January 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Masters:2025:INR

[MNG^{+}25] Rachel Masters, Jalynn Nicoly, Vidya Gaddy, Victoria Interrante, and Francisco Ortega. The impact of nature realism on the restorative quality of virtual reality forest bathing. ACM Transactions on Applied Perception, 22(1):3:1–3:??, January 2025. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3670406.

Morvan:2009:HOT

[MO09a] Yann Morvan and Carol O’Sullivan. Handling occluders in transitions from panoramic images: a perceptual study. ACM Transactions on Applied Perception, 6(4):25:1– 25:??, September 2009. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Morvan:2009:PAT

[MO09b] Yann Morvan and Carol O’Sullivan. A perceptual approach to trimming and tuning unstructured lumigraphs. ACM Transactions on Applied Perception, 5(4):19:1–19:??, January 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Munn:2009:FAI

[MP09] Susan M. Munn and Jeff B. Pelz. FixTag: An algorithm for identifying and tagging fixations to simplify the analysis of data collected by portable eye trackers. ACM Transactions on Applied Perception, 6(3):16:1– 16:??, August 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mihelac:2020:ICH

[MP20] Lorena Mihelac and Janez Povh. The impact of the complexity of harmony on the acceptability of music. ACM Transactions on Applied Perception, 17(1):3:1–3:27, March 2020. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3375014.

Miner:2005:ACM

[MPC05] Nadine E. Miner, Victor E. Vergara Panaiotis, and Thomas Preston Caudell. Authors’ comments on Miner and Caudell, ICAD 1997. ACM Transactions on Applied Perception, 2(4):529–533, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [MC05].

Moffat:2018:PES

[MR18] David Moffat and Joshua D.

Reiss. Perceptual evaluation of synthesized sound effects. ACM Transactions on Applied Perception, 15(2):13:1–13:??, April 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mourkoussis:2010:QFV

[MRT^{+}10] Nicholaos Mourkoussis, Fiona M.

Rivera, Tom Troscianko, Tim Dixon, Rycharde Hawkes, and Katerina Mania. Quantifying fidelity for virtual environment simulations employing memory schema assumptions. ACM Transactions on Applied Perception, 8(1):2:1–2:??, October 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

McDonnell:2012:ISI

[MS12] Rachel McDonnell and Veronica Sundstedt. Introduction to special issue SAP 2012. ACM Transactions on Applied Perception, 9(3):10:1–10:??, July 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Morrison-Smith:2016:UAC

[MSHLR16] Sarah Morrison-Smith, Megan Hofmann, Yang Li, and Jaime Ruiz. Using audio cues to support motion gesture interaction on mobile devices. ACM Transactions on Applied Perception, 13(3):16:1–16:??, May 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Miyashita:2022:DSD

[MSS^{+}22] Yamato Miyashita, Yasuhito Sawahata, Akihiro Sakai, Masamitsu Harasawa, Kazuhiro Hara, Toshiya Morita, and Kazuteru Komine. Display-size dependent effects of 3D viewing on subjective impressions. ACM Transactions on Applied Perception, 19(2):5:1–5:15, April 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3510461.

Makarov:2024:HIO

[MTÆ^{+}24] Ivan Makarov, Snorri Steinn Stefánsson Thors, Elvar Atli Ævarsson, Finnur Kári Pind Jörgensson, Nashmin Yeganeh, Árni Kristjánsson, and Runar Unnthorsson. The haptic intensity order illusion is caused by amplitude changes. ACM Transactions on Applied Perception, 21(1):4:1–4:??, January 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3626237.

Mohler:2007:CLR

[MTCR^{+}07] Betty J. Mohler, William B.

Thompson, Sarah H. Creem-Regehr, Peter Willemsen, Herbert L. Pick, Jr., and John J. Rieser. Calibration of locomotion resulting from visual motion in a treadmill-based virtual environment. ACM Transactions on Applied Perception, 4(1):??, January 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Maxim:2025:PRV

[MVL25] Andrew Maxim, Roshan Venkatakrishnan, and Benjamin Lok. Perceived realism and voice naturalness of virtual humans: Structural equation modeling of behavioral intentions among Black American adults. ACM Transactions on Applied Perception, 22(4):17:1–17:16, October 2025. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Mihtsentu:2015:DVS

[MW15] Mezgeb Tesfayesus Mihtsentu and Colin Ware. Discrete versus solid: Representing quantity using linear, area, and volume glyphs. ACM Transactions on Applied Perception, 12(3): 12:1–12:??, July 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Napieralski:2011:NFD

[NAB^{+}11] Phillip E. Napieralski, Bliss M.

Altenhoff, Jeffrey W. Bertrand, Lindsay O. Long, Sabarish V. Babu, Christopher C. Pagano, Justin Kern, and Timothy A. Davis. Near-field distance perception in real and virtual environments using both verbal and action responses. ACM Transactions on Applied Perception, 8(3):18:1–18:??, August 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Navarro:2011:PCM

[NCSG11] Fernando Navarro, Susana Castillo, Francisco J. Serón, and Diego Gutierrez. Perceptual considerations for motion blur rendering. ACM Transactions on Applied Perception, 8(3):20:1–20:??, August 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Newsham:2010:CLQ

[NCVW10] Guy R. Newsham, Duygu Cetegen, Jennifer A. Veitch, and Lorne Whitehead. Comparing lighting quality evaluations of real scenes with those from high dynamic range and conventional images. ACM Transactions on Applied Perception, 7(2):13:1–13:??, February 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Nehme:2021:CSM

[NFD^{+}21] Yana Nehmé, Jean-Philippe Farrugia, Florent Dupont, Patrick Le Callet, and Guillaume Lavoué. Comparison of subjective methods for quality assessment of 3D graphics in virtual reality. ACM Transactions on Applied Perception, 18(1):2:1–2:23, January 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3427931.

Neumann:2006:IRP

[NG06] Dirk Neumann and Karl R.

Gegenfurtner. Image retrieval and perceptual similarity. ACM Transactions on Applied Perception, 3(1):31–47, January 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Nymoen:2013:ACB

[NGJT13] Kristian Nymoen, Rolf Inge Godøy, Alexander Refsum Jensenius, and Jim Torresen. Analyzing correspondence between sound objects and body motion. ACM Transactions on Applied Perception, 10(2):9:1– 9:??, May 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Niemenlehto:2006:DES

[NJS06]  Pekka-Henrik  Niemenlehto,

Martti Juhola, and Veikko Surakka. Detection of electromyographic signals from facial muscles with neural networks. ACM Transactions on Applied Perception, 3(1):48–61, January 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Narciso:2020:IOW

[NMVRB20] David Narciso, Miguel Melo, José Vasconcelos-Raposo, and Maximino Bessa. The impact of olfactory and wind stimuli on 360 videos using head-mounted displays. ACM Transactions on Applied Perception, 17(1): 4:1–4:13, March 2020. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3380903.

Noceti:2017:EBM

[NOSS17] Nicoletta Noceti, Francesca Odone, Alessandra Sciutti, and Giulio Sandini. Exploring biological motion regularities of human actions: a new perspective on video analysis. ACM Transactions on Applied Perception, 14(3):21:1–21:??, July 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Niewiadomski:2015:EWP

[NP15] Radoslaw Niewiadomski and Catherine Pelachaud. The effect of wrinkles, presentation mode, and intensity on the perception of facial actions and full-face expressions of laughter. ACM Transactions on Applied Perception, 12(1):2:1–2:??, March 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Nasiri:2023:CNT

[NPKR23] Moloud Nasiri, John Porter, Kristopher Kohm, and Andrew Robb. Changes in navigation over time: a comparison of teleportation and joystickbased locomotion. ACM Transactions on Applied Perception, 20(4):16:1–16:??, October 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3613902.

Niu:2012:VES

[NTKA12] Yaqing Niu, Rebecca M. Todd, Matthew Kyan, and Adam K. Anderson. Visual and emotional salience influence eye movements. ACM Transactions on Applied Perception, 9 (3):13:1–13:??, July 2012. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Nunez-Varela:2013:MGC

[NVW13] Jose Nunez-Varela and Jeremy L.

Wyatt. Models of gaze control for manipulation tasks. ACM Transactions on Applied Perception, 10(4):20:1–20:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Nees:2008:DDT

[NW08] Michael A. Nees and Bruce N.

Walker. Data density and trend reversals in auditory graphs: Effects on point-estimation and trend-identification tasks. ACM Transactions on Applied Perception, 5(3):13:1–13:??, August 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Nguyen:2011:ESC

[NZG^{+}11] Tien Dat Nguyen, Christine J. Ziemer, Timofey Grechkin, Benjamin Chihak, Jodie M. Plumert, James F. Cremer, and Joseph K. Kearney. Effects of scale change on distance perception in virtual environments. ACM Transactions on Applied Perception, 8(4):26:1– 26:??, November 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

OToole:2012:CFR

[OAD^{+}12] Alice J. O’Toole, Xaiobo An, Joseph Dunlop, Vaidehi Natu, and P. Jonathon Phillips. Comparing face recognition algorithms to humans on challenging tasks. ACM Transactions on Applied Perception, 9(4): 16:1–16:??, October 2012. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ondrej:2016:FDA

[OEMO16] Jan Ondrej, Cathy Ennis, Niamh A. Merriman, and Carol O’sullivan. FrankenFolk: Distinctiveness and attractiveness of voice and motion. ACM Transactions on Applied Perception, 13(4):20:1–20:??, July 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ohara:2025:SLE

[OMN25] Ryo Ohara, Fuyuki Matsubara, and Takuji Narumi. Smile to like: Effects of facial expressionbased evaluation on media engagement and emotional selfawareness. ACM Transactions on Applied Perception, 22(4): 14:1–14:16, October 2025. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Oulasvirta:2012:HRR

[ONS12] Antti Oulasvirta, Antti Nurminen, and Tiia Suomalainen. How real is real enough? Optimal reality sampling for fast recognition of mobile imagery. ACM Transactions on Applied Perception, 9(4):21:1– 21:??, October 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Otake:2022:TTD

[OOAY22] Kazuya Otake, Shogo Okamoto, Yasuhiro Akiyama, and Yoji Yamada. Tactile texture display combining vibrotactile and electrostatic-friction stimuli: Substantial effects on realism and moderate effects on behavioral responses. ACM Transactions on Applied Perception, 19(4):18:1–18:??, October 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3539733.

Ottaviani:2004:APS

[OR04] Laura Ottaviani and Davide Rocchesso. Auditory perception of 3D size: Experiments with synthetic resonators. ACM Transactions on Applied Perception, 1(2):118–129, October 2004. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

OSullivan:2005:CA

[O’S05] Carol O’Sullivan. Collisions and attention. ACM Transactions on Applied Perception, 2 (3):309–321, July 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Palmer:2008:EAT

[PCK08] Evan M. Palmer, Timothy C. Clausner, and Philip J. Kellman. Enhancing air traffic displays via perceptual cues. ACM Transactions on Applied Perception, 5(1):4:1–4:??, January 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Perrotin:2017:SLD

[PD17] Olivier Perrotin and Christophe D’Alessandro. Seeing, listening, drawing: Interferences between sensorimotor modalities in the use of a tablet musical interface. ACM Transactions on Applied Perception, 14(2):10:1– 10:??, February 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Payandeh:2005:SLD

[PDZ05] Shahram Payandeh, John Dill, and Jian Zhang. A study of level-of-detail in haptic rendering. ACM Transactions on Applied Perception, 2(1):15–34, January 2005. CODEN ????

ISSN 1544-3558 (print), 1544-3965 (electronic).

Pfluger:2015:SFW

[PHRE15] Hermann Pflüger, Benjamin Höferlin, Michael Raschke, and Thomas Ertl. Simulating fixations when looking at visual arts. ACM Transactions on Applied Perception, 12(3):9:1– 9:??, July 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Peters:2008:ACT

[PI08] Robert J. Peters and Laurent Itti. Applying computational tools to predict gaze direction in interactive visual environments. ACM Transactions on Applied Perception, 5(2):9:1–9:??, May 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Park:2021:ETT

[PJGE21] Wanjoo Park, Muhammad Hassan Jamil, Ruth Ghidey Gebremedhin, and Mohamad Eid. Effects of tactile textures on preference in visuo-tactile exploration. ACM Transactions on Applied Perception, 18(2): 9:1–9:13, June 2021. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3449065.

Phillips:2011:ORE

[PJN^{+}11] P. Jonathon Phillips, Fang Jiang, Abhijit Narvekar, Julianne Ayyad, and Alice J.

O’Toole. An other-race effect for face recognition algorithms. ACM Transactions on Applied Perception, 8(2):14:1– 14:??, January 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Pelah:2007:EWR

[PK07] Adar Pelah and Jan J. Koenderink. Editorial: Walking in real and virtual environments. ACM Transactions on Applied Perception, 4(1):??, January 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Parikh:2020:FWL

[PK20] Saurin S. Parikh and Hari Kalva. Feature weighted linguistics classifier for predicting learning difficulty using eye tracking. ACM Transactions on Applied Perception, 17 (2):5:1–5:25, May 2020. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3380877.

Plumert:2005:DPR

[PKCR05] Jodie M. Plumert, Joseph K. Kearney, James F. Cremer, and Kara Recker. Distance perception in real and virtual environments. ACM Transactions on Applied Perception, 2(3):216– 233, July 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Piorkowski:2017:ADG

[PMS17] Rafal Piórkowski, Radoslaw Mantiuk, and Adam Siekawa. Automatic detection of game engine artifacts using full reference image quality metrics. ACM Transactions on Applied Perception, 14(3):14:1–14:??, July 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Payne:2024:EES

[PR24] Jordan Payne and Andrew Robb. Exploring the effects of self-overlapping spaces on distance perception and action judgments. ACM Transactions on Applied Perception, 21(4):15:1–15:??, October 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3695632.

Parde:2023:TIV

[PSB^{+}23] Connor J. Parde, Virginia E.

Strehle, Vivekjyoti Banerjee, Ying Hu, Jacqueline G. Cavazos, Carlos D. Castillo, and Alice J. O’Toole. Twin identification over viewpoint change: a deep convolutional neural network surpasses humans. ACM Transactions on Applied Perception, 20(3):10:1–10:??, July 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3609224.

Pacchierotti:2014:ITT

[PTP14] Claudio Pacchierotti, Asad Tirmizi, and Domenico Prattichizzo. Improving transparency in teleoperation by means of cutaneous tactile force feedback. ACM Transactions on Applied Perception, 11(1): 4:1–4:??, April 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Pouke:2024:ASH

[PUC^{+}24] Matti Pouke, Elmeri Uotila, Evan G. Center, Kalle G. Timperi, Alexis P. Chambers, Timo Ojala, and Steven M. Lavalle. Adaptation to simulated hypergravity in a virtual reality throwing task. ACM Transactions on Applied Perception, 21 (2):8:1–8:??, April 2024. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3643849.

Plaisier:2020:LVM

[PVK20] Myrthe A. Plaisier, Daphne S. Vermeer, and Astrid M. L. Kappers. Learning the vibrotactile Morse code alphabet. ACM Transactions on Applied Perception, 17(3):9:1– 9:10, November 2020. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3402935.

Pineo:2010:NMF

[PW10] Daniel Pineo and Colin Ware. Neural modeling of flow rendering effectiveness. ACM Transactions on Applied Perception, 7(3):20:1–20:??, June 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Reinhard:2004:E

[RB04] Erik Reinhard and Heinrich Bülthoff. Editorial. ACM Transactions on Applied Perception, 1(1):1–2, July 2004. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Reinhard:2008:E

[RB08] Erik Reinhard and Heinrich Bülthoff. Editorial. ACM Transactions on Applied Perception, 5(2):6:1–6:??, May 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ramic-Brkic:2014:OAV

[RBC14] Belma Ramic-Brkic and Alan Chalmers. Olfactory adaptation in virtual environments. ACM Transactions on Applied Perception, 11(2):6:1–6:??, July 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Rebillat:2012:AVA

[RBCK12] Marc Rébillat, Xavier Boutillon, Étienne Corteel, and Brian F. G. Katz. Audio, visual, and audio-visual egocentric distance perception by moving subjects in virtual environments. ACM Transactions on Applied Perception, 9(4):19:1–19:??, October 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Rosenholtz:2011:DPV

[RDF11] Ruth Rosenholtz, Amal Dorai, and Rosalind Freeman. Do predictions of visual perception aid design? ACM Transactions on Applied Perception, 8(2):12:1– 12:??, January 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Robles-De-La-Torre:2004:NEI

[RDLTS04] G. Robles-De-La-Torre and R. Sekuler. Numerically estimating internal models of dynamic virtual objects. ACM Transactions on Applied Perception, 1(2):102–117, October 2004. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Riecke:2009:ASM

[RFR09] Bernhard E. Riecke, Daniel Feuereissen, and John J. Rieser. Auditory self-motion simulation is facilitated by haptic and vibrational cues suggesting the possibility of actual motion. ACM Transactions on Applied Perception, 6(3):20:1– 20:??, August 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Riecke:2018:ISI

[RK18] Bernhard Riecke and Joseph Kearny. Introduction to special issue SAP 2018. ACM Transactions on Applied Perception, 15(4):23:1–23:??, October 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3274477.

Rajasekaran:2022:PPT

[RKC^{+}22] Suren Deepak Rajasekaran, Hao Kang, Martin Cadík, Eric Galin, Eric Guérin, Adrien Peytavie, Pavel Slavík, and Bedrich Benes. PTRM: Perceived terrain realism metric. ACM Transactions on Applied Perception, 19(2):6:1–6:22, April 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3514244.

Robb:2022:EML

[RKP22] Andrew Robb, Kristopher Kohm, and John Porter. Experience matters: Longitudinal changes in sensitivity to rotational gains in virtual reality. ACM Transactions on Applied Perception, 19(4):16:1– 16:??, October 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3560818.

Rigas:2016:BRE

[RKS16] Ioannis Rigas, Oleg Komogortsev, and Reza Shadmehr. Biometric recognition via eye movements: Saccadic vigor and acceleration cues. ACM Transactions on Applied Perception, 13(2):6:1–6:??, March 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Rafian:2017:RSA

[RL17] Paymon Rafian and Gordon E. Legge. Remote sighted assistants for indoor location sensing of visually impaired pedestrians. ACM Transactions on Applied Perception, 14(3):19:1– 19:??, July 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Radun:2008:CQI

[RLH^{+}08] Jenni Radun, Tuomas Leisti, Jukka Häkkinen, Harri Ojanen, Jean-Luc Olives, Tero Vuori, and Göte Nyman. Content and quality: Interpretationbased estimation of image quality. ACM Transactions on Applied Perception, 4(4):2:1–2:??, January 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Radun:2010:EMV

[RLV^{+}10] Jenni Radun, Tuomas Leisti, Toni Virtanen, Jukka Häkkinen, Tero Vuori, and Göte Nyman. Evaluating the multivariate visual quality performance of image-processing components. ACM Transactions on Applied Perception, 7 (3):16:1–16:??, June 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Rocchesso:2012:PRP

[RM12] Davide Rocchesso and Stefano Delle Monache. Perception and replication of planar sonic gestures. ACM Transactions on Applied Perception, 9(4):18:1– 18:??, October 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Rummukainen:2016:RRM

[RM16] Olli Rummukainen and Catarina Mendonça. Reproducing reality: Multimodal contributions in natural scene discrimination. ACM Transactions on Applied Perception, 14(1):1:1– 1:??, August 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Radun:2016:IDI

[RNLH16] Jenni Radun, Mikko Nuutinen, Tuomas Leisti, and Jukka Häkkinen. Individual differences in image-quality estimations: Estimation rules and viewing strategies. ACM Transactions on Applied Perception, 13(3):14:1–14:??, May 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Reitsma:2009:ESP

[RO09] Paul S. A. Reitsma and Carol O’Sullivan. Effect of scenario on perceptual sensitivity to errors in animation. ACM Transactions on Applied Perception, 6(3):15:1–15:??, August 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Rienks:2010:DHO

[RPH10] Rutger Rienks, Ronald Poppe, and Dirk Heylen. Differences in head orientation behavior for speakers and listeners: An experiment in a virtual environment. ACM Transactions on Applied Perception, 7(1):2:1– 2:??, January 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Rungta:2016:PCP

[RRM^{+}16] Atul Rungta, Sarah Rust, Nicolas Morales, Roberta Klatzky, Ming Lin, and Dinesh Manocha. Psychoacoustic characterization of propagation effects in virtual environments. ACM Transactions on Applied Perception, 13(4):21:1–21:??, July 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Renner:2015:ISB

[RSM^{+}15] Rebekka S. Renner, Erik Steindecker, Mathias MüLler, Boris M. Velichkovsky, Ralph Stelzer, Sebastian Pannasch, and Jens R. Helmert. The influence of the stereo base on blind and sighted reaches in a virtual environment. ACM Transactions on Applied Perception, 12 (2):7:1–7:??, April 2015. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Riecke:2006:CFC

[RSPA^{+}06] Bernhard E. Riecke, Jörg Schulte-Pelkum, Marios N.

Avraamides, Markus Von Der Heyde, and Heinrich H. Bülthoff. Cognitive factors can influence self-motion perception (vection) in virtual reality. ACM Transactions on Applied Perception, 3(3):194–216, July 2006. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Reed:2021:IWP

[RTJ^{+}21] Charlotte M. Reed, Hong Z.

Tan, Yang Jiao, Zachary D. Perez, and E. Courtenay Wilson. Identification of words and phrases through a phonemicbased haptic display: Effects of inter-phoneme and interword interval durations. ACM Transactions on Applied Perception, 18(3):13:1–13:22, July 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3458725.

Rosli:2011:AGC

[RTPG11] Roslizawaty Mohd Rosli, Hong Z.

Tan, Robert W. Proctor, and Rob Gray. Attentional gradient for crossmodal proximal-distal tactile cueing of visual spatial attention. ACM Transactions on Applied Perception, 8 (4):23:1–23:??, November 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ramesh:2018:AHS

[RTSW18] Girish Ramesh, Martin Turner, Björn Schröder, and Franz Wortmann. Analysis of hair shine using rendering and subjective evaluation. ACM Transactions on Applied Perception, 15(4):25:1–25:??, October 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3274478.

Rushmeier:2005:GE

[Rus05] Holly Rushmeier. Guest editorial. ACM Transactions on Applied Perception, 2(3):181– 182, July 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Riecke:2005:VCC

[RVB05] Bernhard E. Riecke, Markus Von Der Heyde, and Heinrich H. Bülthoff. Visual cues can be sufficient for triggering automatic, reflexlike spatial updating. ACM Transactions on Applied Perception, 2(3):183– 215, July 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ruddle:2013:LWV

[RVB13] Roy A. Ruddle, Ekaterina Volkova, and Heinrich H. Bülthoff. Learning to walk in virtual reality. ACM Transactions on Applied Perception, 10 (2):11:1–11:??, May 2013. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Rosa:2019:SHI

[RVH^{+}19] Nina Rosa, Remco C. Veltkamp, Wolfgang Hürst, Tanja Nijboer, Carolien Gilbers, and Peter Werkhoven. The supernumerary hand illusion in augmented reality. ACM Transactions on Applied Perception, 16(2):12:1–12:??, August 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3341225.

Riecke:2009:MSE

[RVSP09] Bernhard E. Riecke, Aleksander Väljamäe, and Jörg Schulte-Pelkum. Moving sounds enhance the visually-induced selfmotion illusion (circular vection) in virtual reality. ACM Transactions on Applied Perception, 6(2):7:1–7:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Seward:2007:UVE

[SAB07] A. Elizabeth Seward, Daniel H.

Ashmead, and Bobby Bodenheimer. Using virtual environments to assess time-to-contact judgments from pedestrian viewpoints. ACM Transactions on Applied Perception, 4(3):18:1–18:??, November 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Schumacher:2012:WFP

[SB12] Matthaeus Schumacher and Volker Blanz. Which facial profile do humans expect after seeing a frontal view? a comparison with a linear face model. ACM Transactions on Applied Perception, 9(3):11:1– 11:??, July 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Serrano:2022:ISI

[SBC22] Ana Serrano and Michael Barnett-Cowan. Introduction to the special issue on SAP 2022. ACM Transactions on Applied Perception, 19(4):14:1– 14:??, October 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563136.

Sprague:2007:MEV

[SBR07] Nathan Sprague, Dana Ballard, and Al Robinson. Modeling embodied visual behaviors. ACM Transactions on Applied Perception, 4(2):11:1–11:??, July 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Seifi:2018:TAH

[SCM18] Hasti Seifi, Mattew Chun, and Karon E. Maclean. Toward affective handles for tuning vibrations. ACM Transactions on Applied Perception, 15(3):22:1– 22:??, August 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Sahm:2005:TVW

[SCRTW05] Cynthia S. Sahm, Sarah H. Creem-Regehr, William B. Thompson, and Peter Willemsen. Throwing versus walking as indicators of distance perception in similar real and virtual environments. ACM Transactions on Applied Perception, 2 (1):35–45, January 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Shinn-Cunningham:2005:SAD

[SCS05] Barbara G. Shinn-Cunningham and Timothy Streeter. Spatial auditory display: Comments on Shinn-Cunningham et al., ICAD 2001. ACM Transactions on Applied Perception, 2 (4):426–429, October 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [SCSG05].

Shinn-Cunningham:2005:PPS

[SCSG05] Barbara G. Shinn-Cunningham, Timothy Streeter, and Jean-François Gyss. Perceptual plasticity in spatial auditory displays. ACM Transactions on Applied Perception, 2(4):418– 425, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [SCS05].

Scott:2022:ERE

[SD22] Joshua J. Scott and Neil A.

Dodgson. Evaluating realism in example-based terrain synthesis. ACM Transactions on Applied Perception, 19(3): 11:1–11:??, July 2022. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3531526.

Selmanovic:2013:GSH

[SDBRC13] Elmedin Selmanović, Kurt Debattista, Thomas Bashford-Rogers, and Alan Chalmers. Generating stereoscopic HDR images using HDR–LDR image pairs. ACM Transactions on Applied Perception, 10(1):3:1– 3:??, February 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Sonlu:2025:PEU

[SDG25] Sinan Sonlu, Halil Özgür Demir, and Ugur Güdükbay. Personality expression using cospeech gesture. ACM Transactions on Applied Perception, 22 (2):7:1–7:??, April 2025. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3694905.

Shin:2005:VCA

[SDW05] Do Hyoung Shin, Phillip S. Dunston, and Xiangyu Wang. View changes in augmented reality computer-aided-drawing. ACM Transactions on Applied Perception, 2(1):1–14, January 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Shamy:2023:ILI

[SF23] Mor Shamy and Dror G. Feitelson. Identifying lines and interpreting vertical jumps in eye tracking studies of reading text and code. ACM Transactions on Applied Perception, 20(2):6:1–6:??, April 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579357.

Sundstedt:2007:PRP

[SGA^{+}07] Veronica Sundstedt, Diego Gutierrez, Oscar Anson, Francesco Banterle, and Alan Chalmers. Perceptual rendering of participating media. ACM Transactions on Applied Perception, 4 (3):15:1–15:??, November 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Souman:2010:MVW

[SGF^{+}10] Jan L. Souman, Paolo Robuffo Giordano, Ilja Frissen, Alessandro De Luca, and Marc O. Ernst. Making virtual walking real: Perceptual evaluation of a new treadmill control algorithm. ACM Transactions on Applied Perception, 7(2):11:1– 11:??, February 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Spicker:2019:QVA

[SGHL^{+}19] Marc Spicker, Franz Götz-Hahn, Thomas Lindemeier, Dietmar Saupe, and Oliver Deussen. Quantifying visual abstraction quality for computergenerated illustrations. ACM Transactions on Applied Perception, 16(1):5:1–5:??, February 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3301414.

Souman:2011:CEU

[SGS^{+}11] J. L. Souman, P. Robuffo Giordano, M. Schwaiger, I. Frissen, T. Thümmel, H. Ulbrich, A. De Luca, H. H. Bülthoff, and M. O. Ernst. CyberWalk: Enabling unconstrained omnidirectional walking through virtual environments. ACM Transactions on Applied Perception, 8 (4):25:1–25:??, November 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Schmidt:2005:HNH

[SHBK05] Henning Schmidt, Stefan Hesse, Rolf Bernhardt, and Jörg Krüger. HapticWalker—a novel haptic foot device. ACM Transactions on Applied Perception, 2(2):166–180, April 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Salagean:2022:VRA

[SHHF^{+}22] Anca Salagean, Jacob Hadnett-Hunter, Daniel J. Finnegan, Alexandra A. De Sousa, and Michael J. Proulx. A virtual reality application of the rubber hand illusion induced by ultrasonic mid-air haptic stimulation. ACM Transactions on Applied Perception, 19(1): 3:1–3:19, January 2022. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3487563.

Surakka:2004:GFN

[SII04] Veikko Surakka, Marko Illi, and Poika Isokoski. Gazing and frowning as a new human– computer interaction technique. ACM Transactions on Applied Perception, 1(1):40–56, July 2004. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Singhal:2018:CTI

[SJ18] Anshul Singhal and Lynette A.

Jones. Creating thermal icons — a model-based approach. ACM Transactions on Applied Perception, 15(2):14:1–14:??, April 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Subedar:2016:BD

[SK16] Mahesh M. Subedar and Lina J.

Karam. 3D blur discrimination. ACM Transactions on Applied Perception, 13(3):12:1– 12:??, May 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Schmidtler:2018:HPI

[SK18] Jonas Schmidtler and Moritz Körber. Human perception of inertial mass for joint humanrobot object manipulation. ACM Transactions on Applied Perception, 15(3):15:1–15:??, August 2018. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Stich:2011:PMI

[SLW^{+}11] Timo Stich, Christian Linz, Christian Wallraven, Douglas Cunningham, and Marcus Magnor. Perception-motivated interpolation of image sequences. ACM Transactions on Applied Perception, 8(2):11:1– 11:??, January 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Seuntiens:2006:PQC

[SMI06] Pieter Seuntiens, Lydia Meesters, and Wijnand Ijsselsteijn. Perceived quality of compressed stereoscopic images: Effects of symmetric and asymmetric JPEG coding and camera separation. ACM Transactions on Applied Perception, 3(2):95– 109, April 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Shamir:2010:IES

[SMO^{+}10] Lior Shamir, Tomasz Macura, Nikita Orlov, D. Mark Eckley, and Ilya G. Goldberg. Impressionism, expressionism, surrealism: Automated recognition of painters and schools of art. ACM Transactions on Applied Perception, 7(2):8:1–8:??, February 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Sugano:2013:GBJ

[SMS13] Yusuke Sugano, Yasuyuki Matsushita, and Yoichi Sato. Graph-based joint clustering of fixations and visual entities. ACM Transactions on Applied Perception, 10(2):10:1– 10:??, May 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Shamir:2016:DBA

[SNW16] Lior Shamir, Jenny Nissel, and Ellen Winner. Distinguishing between abstract art by artists vs. children and animals: Comparison between human and machine perception. ACM Transactions on Applied Perception, 13(3):17:1–17:??, May 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Still:2019:IVS

[SS19] Jeremiah Still and Mary Still.

Influence of visual salience on webpage product searches. ACM Transactions on Applied Perception, 16(1):3:1– 3:??, February 2019. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3301413.

Stenholt:2015:BUC

[Ste15] Rasmus Stenholt. On the benefits of using constant visual angle glyphs in interactive exploration of 3D scatterplots. ACM Transactions on Applied Perception, 11(4):19:1– 19:??, January 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Schwaninger:2006:PPM

[SVHS06] Adrian Schwaninger, Julia Vogel, Franziska Hofer, and Bernt Schiele. A psychophysically plausible model for typicality ranking of natural scenes. ACM Transactions on Applied Perception, 3(4):333–353, October 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Stransky:2014:ELT

[SWA14] Debi Stransky, Laurie M. Wilcox, and Robert S. Allison. Effects of long-term exposure on sensitivity and comfort with stereoscopic displays. ACM Transactions on Applied Perception, 11(1):2:1–2:??, April 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Surace:2023:LGB

[SWT^{+}23] Luca Surace, Marek Wernikowski, Cara Tursun, Karol Myszkowski, Radosław Mantiuk, and Piotr Didyk. Learning GANbased foveated reconstruction to recover perceptually important image features. ACM Transactions on Applied Perception, 20(2):7:1–7:??, April 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3583072.

Schuwerk:2015:CEC

[SXCS15] Clemens Schuwerk, Xiao Xu, Rahul Chaudhari, and Eckehard Steinbach. Compensating the effect of communication delay in client-server–based shared haptic virtual environments. ACM Transactions on Applied Perception, 13(1):5:1– 5:??, December 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Schneider:2025:EOS

[SYF^{+}25] Johann Schneider, Yu Yi Yang, Maria Fronius, + 2, and ... An extensible open source software designed for virtual realitybased testing and treatment of vision disorders. ACM Transactions on Applied Perception, 22 (3):11:1–11:??, July 2025. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Su:2022:MLB

[SZ22] Jun Su and Peng Zhou. Machine learning-based modeling and prediction of the intrinsic relationship between human emotion and music. ACM Transactions on Applied Perception, 19(3):12:1–12:??, July 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3534966.

Tulay:2024:DFB

[TB24] Emine Elif Tülay and Tugçe Balli. Decoding functional brain data for emotion recognition: a machine learning approach. ACM Transactions on Applied Perception, 21(3): 11:1–11:??, July 2024. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3657638.

Tadros:2019:ANN

[TCGC19] Timothy Tadros, Nicholas C. Cullen, Michelle R. Greene, and Emily A. Cooper. Assessing neural network scene classification from degraded images. ACM Transactions on Applied Perception, 16(4):21:1– 21:??, September 2019. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3342349.

Trutoiu:2011:MAE

[TCMH11] Laura C. Trutoiu, Elizabeth J. Carter, Iain Matthews, and Jessica K. Hodgins. Modeling and animating eye blinks. ACM Transactions on Applied Perception, 8(3):17:1–17:??, August 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Trutoiu:2014:STL

[TCP^{+}14] Laura C. Trutoiu, Elizabeth J. Carter, Nancy Pollard, Jeffrey F. Cohn, and Jessica K. Hodgins. Spatial and temporal linearities in posed and spontaneous smiles. ACM Transactions on Applied Perception, 11(3):12:1–12:??, August 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Tennison:2019:NVP

[TG19] Jennifer L. Tennison and Jenna L. Gorlewicz. Non-visual perception of lines on a multimodal touchscreen tablet. ACM Transactions on Applied Perception, 16(1):6:1–6:??, February 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3301415.

Toscani:2020:TPD

[TGG^{+}20] Matteo Toscani, Dar’ya Guarnera, Giuseppe Claudio Guarnera, Jon Yngve Hardeberg, and Karl R. Gegenfurtner. Three perceptual dimensions for specular and diffuse reflection. ACM Transactions on Applied Perception, 17(2): 6:1–6:26, May 2020. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3380741.

Tarr:2008:IFA

[TGJ08] Michael J. Tarr, Athinodoros S.

Georghiades, and Cullen D. Jackson. Identifying faces across variations in lighting: Psychophysics and computation. ACM Transactions on Applied Perception, 5(2):10:1– 10:??, May 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

To:2009:PDN

[TGT^{+}09] M. P. S. To, I. D. Gilchrist, T. Troscianko, J. S. B. Kho, and D. J. Tolhurst. Perception of differences in naturalimage stimuli: Why is peripheral viewing poorer than foveal? ACM Transactions on Applied Perception, 6(4):26:1– 26:??, September 2009. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Thompson:2007:GE

[Tho07] William B. Thompson. Guest editorial. ACM Transactions on Applied Perception, 4(3):14:1– 14:??, November 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Thumfart:2011:MHA

[TJL^{+}11] Stefan Thumfart, Richard H. A. H. Jacobs, Edwin Lughofer, Christian Eitzinger, Frans W. Cornelissen, Werner Groissboeck, and Roland Richter. Modeling human aesthetic perception of visual textures. ACM Transactions on Applied Perception, 8(4):27:1– 27:??, November 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Tompkin:2013:PAA

[TKK^{+}13] James Tompkin, Min H. Kim, Kwang In Kim, Jan Kautz, and Christian Theobalt. Preference and artifact analysis for video transitions of places. ACM Transactions on Applied Perception, 10(3):13:1–13:??, August 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Tan:2015:PLI

[TLS^{+}15] Minghui Tan, Jean-François Lalonde, Lavanya Sharan, Holly Rushmeier, and Carol O’sullivan. The perception of lighting inconsistencies in composite outdoor scenes. ACM Transactions on Applied Perception, 12(4):18:1–18:??, September 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Tauscher:2017:CAT

[TMM17] Jan-Philipp Tauscher, Maryam Mustafa, and Marcus Magnor. Comparative analysis of three different modalities for perception of artifacts in videos. ACM Transactions on Applied Perception, 14(4):22:1– 22:??, September 2017. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Thorpe:2020:SRE

[TNE20] Alexander Thorpe, Keith Nesbitt, and Ami Eidels. A systematic review of empirical measures of workload capacity. ACM Transactions on Applied Perception, 17(3):12:1– 12:26, November 2020. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3422869.

Turchet:2013:WPA

[TSC13] Luca Turchet, Stefania Serafin, and Paola Cesari. Walking pace affected by interactive sounds simulating stepping on different terrains. ACM Transactions on Applied Perception, 10(4):23:1– 23:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Tan:2007:DIF

[TSRD07] Hong Z. Tan, Mandayam A. Srinivasan, Charlotte M. Reed, and Nathaniel I. Durlach. Discrimination and identification of finger joint-angle position using active motion. ACM Transactions on Applied Perception, 4(2):10:1–10:??, July 2007. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Tennison:2020:EVB

[TUG^{+}20] Jennifer L. Tennison, P. Merlin Uesbeck, Nicholas A. Giudice, Andreas Stefik, Derrick W. Smith, and Jenna L. Gorlewicz. Establishing vibrationbased tactile line profiles for use in multimodal graphics. ACM Transactions on Applied Perception, 17(2):7:1–7:14, May 2020. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3383457.

TenHolt:2011:HIS

[TVR^{+}11] Gineke A. Ten Holt, Andrea J. Van Doorn, Marcel J. T. Reinders, Emile A. Hendriks, and Huib De Ridder. Humaninspired search for redundancy in automatic sign language recognition. ACM Transactions on Applied Perception, 8(2): 15:1–15:??, January 2011. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Um:2021:SDA

[UHWT21] Kiwon Um, Xiangyu Hu, Bing Wang, and Nils Thuerey. Spot the difference: Accuracy of numerical simulations via the human visual system. ACM Transactions on Applied Perception, 18(2):6:1–6:15, June 2021. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3449064.

Usevitch:2020:TRA

[USA20] David E. Usevitch, Adam J.

Sperry, and Jake J. Abbott. Translational and Rotational Arrow Cues (TRAC) navigation method for manual alignment tasks. ACM Transactions on Applied Perception, 17 (1):1:1–1:19, March 2020. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3375001.

Vickers:2005:MPA

[VA05] Paul Vickers and James L.

Alty. Musical program auralization: Empirical studies. ACM Transactions on Applied Perception, 2(4):477–489, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [Vic05].

VargasMartin:2016:DSF

[VCA16] Miguel Vargas Martin, Victor Cho, and Gabriel Aversano. Detection of subconscious face recognition using consumergrade brain-computer interfaces. ACM Transactions on Applied Perception, 14(1):7:1– 7:??, August 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

VandenBerg:2008:PDI

[VCR08] Ronald Van den Berg, Frans W. Cornelissen, and Jos B. T. M. Roerdink. Perceptual dependencies in information visualization assessed by complex visual search. ACM Transactions on Applied Perception, 4 (4):3:1–3:??, January 2008. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

vandenDoel:2005:PSC

[vdD05a] Kees van den Doel. From physics to sound: Comments on van den Doel, ICAD 2004. ACM Transactions on Applied Perception, 2(4):547–549, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [vdD05b].

vandenDoel:2005:PBM

[vdD05b] Kees van den Doel. Physically based models for liquid sounds. ACM Transactions on Applied Perception, 2(4):534–546, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [vdD05a].

Vicentini:2010:EFT

[VGBF10] M. Vicentini, S. Galvan, D. Botturi, and P. Fiorini. Evaluation of force and torque magnitude discrimination thresholds on the human hand-arm system. ACM Transactions on Applied Perception, 8(1):1:1– 1:??, October 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Vicovaro:2014:PEM

[VHBO14] Michele Vicovaro, Ludovic Hoyet, Luigi Burigana, and Carol O’sullivan. Perceptual evaluation of motion editing for realistic throwing animations. ACM Transactions on Applied Perception, 11(2):10:1– 10:??, July 2014. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Vickers:2005:PAA

[Vic05] Paul Vickers. Program auralization: Author’s comments on Vickers and Alty, ICAD 2000. ACM Transactions on Applied Perception, 2(4):490–494, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [VA05].

Vanhala:2012:VFA

[VSCM12] Toni Vanhala, Veikko Surakka, Matthieu Courgeon, and Jean-Claude Martin. Voluntary facial activations regulate physiological arousal and subjective experiences during virtual social stimulation. ACM Transactions on Applied Perception, 9(1):1:1–1:??, March 2012. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Vanhoey:2017:VQA

[VSKL17] Kenneth Vanhoey, Basile Sauvage, Pierre Kraemer, and Guillaume Lavoué. Visual quality assessment of 3D models: On the influence of light-material interaction. ACM Transactions on Applied Perception, 15(1):5:1– 5:??, November 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Vogel:2007:CNS

[VSWB07] Julia Vogel, Adrian Schwaninger, Christian Wallraven, and Heinrich H. Bülthoff. Categorization of natural scenes: Local versus global information and the role of color. ACM Transactions on Applied Perception, 4(3):19:1– 19:??, November 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Venkatakrishnan:2022:ICI

[VVC^{+}22] Roshan Venkatakrishnan, Rohith Venkatakrishnan, Chih-Han Chung, Yu-Shuen Wang, and Sabarish Babu. Investigating a combination of input modalities, canvas geometries, and inking triggers on onair handwriting in virtual reality. ACM Transactions on Applied Perception, 19(4):15:1– 15:??, October 2022. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3560817.

VanMensvoort:2010:PMO

[VVHV10] Koert Van Mensvoort, Peter Vos, Dik J. Hermes, and Robert Van Liere. Perception of mechanically and optically simulated bumps and holes. ACM Transactions on Applied Perception, 7(2):10:1–10:??, February 2010. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

VanErp:2005:WNV

[VVJD05] Jan B. F. Van Erp, Hendrik A.

H. C. Van Veen, Chris Jansen, and Trevor Dobbins. Waypoint navigation with a vibrotactile waist belt. ACM Transactions on Applied Perception, 2(2):106–117, April 2005. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wilcox:2006:PSV

[WAEG06] Laurie M. Wilcox, Robert S. Allison, Samuel Elfassy, and Cynthia Grelik. Personal space in virtual reality. ACM Transactions on Applied Perception, 3 (4):412–428, October 2006. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wilcox:2015:EVP

[WAH^{+}15] Laurie M. Wilcox, Robert S. Allison, John Helliker, Bert Dunk, and Roy C. Anthony. Evidence that viewers prefer higher frame-rate film. ACM Transactions on Applied Perception, 12 (4):15:1–15:??, September 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ware:2004:MSR

[WB04] Colin Ware and Robert Bobrow. Motion to support rapid interactive queries on node–link diagrams. ACM Transactions on Applied Perception, 1(1):3– 18, July 2004. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wallraven:2007:ERW

[WBC^{+}07] Christian Wallraven, Heinrich H. Bülthoff, Douglas W. Cunningham, Jan Fischer, and Dirk Bartz. Evaluation of real-world and computergenerated stylized facial expressions. ACM Transactions on Applied Perception, 4(3):16:1– 16:??, November 2007. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wallraven:2008:EPR

[WBCB08] Christian Wallraven, Martin Breidt, Douglas W. Cunningham, and Heinrich H. Bülthoff. Evaluating the perceptual realism of animated facial expressions. ACM Transactions on Applied Perception, 4(4):4:1– 4:??, January 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wang:2020:TQA

[WBHP20] Xi Wang, Zoya Bylinskii, Aaron Hertzmann, and Robert Pepperell. Toward quantifying ambiguities in artistic images. ACM Transactions on Applied Perception, 17(4):13:1–13:10, December 2020. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418054.

Williams:2011:EWP

[WBN^{+}11] Betsy Williams, Stephen Bailey, Gayathri Narasimham, Muqun Li, and Bobby Bodenheimer. Evaluation of walking in place on a Wii balance board to explore a virtual environment. ACM Transactions on Applied Perception, 8(3):19:1– 19:??, August 2011. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Weidenbacher:2006:SSS

[WBNF06] Ulrich Weidenbacher, Pierre Bayerl, Heiko Neumann, and Roland Fleming. Sketching shiny surfaces: 3D shape extraction and depiction of specular surfaces. ACM Transactions on Applied Perception, 3 (3):262–285, July 2006. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wang:2022:PGO

[WC22] Minqi Wang and Emily A.

Cooper. Perceptual guidelines for optimizing field of view in stereoscopic augmented reality displays. ACM Transactions on Applied Perception, 19(4):19:1–19:??, October 2022. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3554921.

Willemsen:2009:EHM

[WCCRT09] Peter Willemsen, Mark B. Colton, Sarah H. Creem-Regehr, and William B. Thompson. The effects of headmounted display mechanical properties and field of view on distance judgments in virtual environments. ACM Transactions on Applied Perception, 6(2):8:1–8:??, February 2009. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wang:2024:EIC

[WDLC24] Minqi Wang, Jian Ding, Dennis M. Levi, and Emily A. Cooper. The effect of interocular contrast differences on the appearance of augmented reality imagery. ACM Transactions on Applied Perception, 21(1):1:1–1:??, January 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3617684.

Wang:2008:TSP

[WH08] Qi Wang and Vincent Hayward. Tactile synthesis and perceptual inverse problems seen from the viewpoint of contact mechanics. ACM Transactions on Applied Perception, 5(2): 7:1–7:??, May 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wohler:2024:IPF

[WIA24] Leslie Wöhler, Satoshi Ikehata, and Kiyoharu Aizawa. Investigating the perception of facial anonymization techniques in 360^{∘} videos. ACM Transactions on Applied Perception, 21(4):16:1–16:??, October 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3695254.

Walker:2005:MMA

[WK05a] Bruce N. Walker and Gregory Kramer. Mappings and metaphors in auditory displays: An experimental assessment. ACM Transactions on Applied Perception, 2(4):407–412, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See comments [WK05b].

Walker:2005:SDM

[WK05b] Bruce N. Walker and Gregory Kramer. Sonification design and metaphors: Comments on Walker and Kramer, ICAD 1996. ACM Transactions on Applied Perception, 2(4):413– 417, October 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). See [WK05a].

Williams:2015:IPE

[WKM^{+}15] Duncan Williams, Alexis Kirke, Eduardo Miranda, Ian Daly, James Hallowell, James Weaver, Asad Malik, Etienne Roesch, Faustina Hwang, and Slawomir Nasuto. Investigating perceived emotional correlates of rhythmic density in algorithmic music composition. ACM Transactions on Applied Perception, 12(3):8:1–8:??, July 2015. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Williams:2017:ACM

[WKM^{+}17] Duncan Williams, Alexis Kirke, Eduardo Miranda, Ian Daly, Faustina Hwang, James Weaver, and Slawomir Nasuto. Affective calibration of musical feature sets in an emotionally intelligent music composition system. ACM Transactions on Applied Perception, 14(3):17:1– 17:??, July 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wei:2021:CMG

[WL21] Hui Wei and Jingmeng Li. Computational model for global contour precedence based on primary visual cortex mechanisms. ACM Transactions on Applied Perception, 18(3): 14:1–14:21, July 2021. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3459999.

Ware:2008:VGT

[WM08] Colin Ware and Peter Mitchell. Visualizing graphs in three dimensions. ACM Transactions on Applied Perception, 5(1): 2:1–2:??, January 2008. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Watanabe:2012:GCV

[WMA12] Junji Watanabe, Taro Maeda, and Hideyuki Ando. Gazecontingent visual presentation technique with electro-ocular-graph-based saccade detection. ACM Transactions on Applied Perception, 9(2):6:1–6:??, June 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Watters:2008:VDL

[WMS08] Paul Watters, Frances Martin, and H. Steffen Stripf. Visual detection of LSB-encoded natural image steganography. ACM Transactions on Applied Perception, 5(1):5:1–5:??, January 2008. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Webster:2005:NTD

[WMVO05] Robert J. Webster III, Todd E.

Murphy, Lawton N. Verner, and Allison M. Okamura. A novel two-dimensional tactile slip display: design, kinematics and perceptual experiments. ACM Transactions on Applied Perception, 2(2):150–165, April 2005. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Williams:2007:FSS

[WNW^{+}07] Betsy Williams, Gayathri Narasimham, Claire Westerman, John Rieser, and Bobby Bodenheimer. Functional similarities in spatial representations between real and virtual environments. ACM Transactions on Applied Perception, 4 (2):12:1–12:??, July 2007. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wijntjes:2010:PPS

[WP10] Maarten W. A. Wijntjes and Sylvia C. Pont. Pointing in pictorial space: Quantifying the perceived relative depth structure in mono and stereo images of natural scenes. ACM Transactions on Applied Perception, 7(4):24:1–24:??, July 2010. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wang:2014:OGL

[WPDH14] Rui I. Wang, Brandon Pelfrey, Andrew T. Duchowski, and Donald H. House. Online 3D gaze localization on stereoscopic displays. ACM Transactions on Applied Perception, 11 (1):3:1–3:??, April 2014. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Weier:2018:FDF

[WRHS18] Martin Weier, Thorsten Roth, André Hinkenjann, and Philipp Slusallek. Foveated depth-of-field filtering in head-mounted displays. ACM Transactions on Applied Perception, 15(4): 26:1–26:??, October 2018. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3238301.

Wang:2016:AIH

[WTWN16] Yingying Wang, Jean E. Fox Tree, Marilyn Walker, and Michael Neff. Assessing the impact of hand motion on virtual character personality. ACM Transactions on Applied Perception, 13(2):9:1–9:??, March 2016. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wilkie:2011:MLC

[WWA11] Richard M. Wilkie, John P.

Wann, and Robert S. Allison. Modeling locomotor control: The advantages of mobile gaze. ACM Transactions on Applied Perception, 8(2): 9:1–9:??, January 2011. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Wang:2023:HLB

[WZA^{+}23] Yuanhao Wang, Qian Zhang, Celine Aubuchon, Jovan Kemp, Fulvio Domini, and James Tompkin. On human-like biases in convolutional neural networks for the perception of slant from texture. ACM Transactions on Applied Perception, 20 (4):15:1–15:??, October 2023. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3613451.

Yu:2004:MLI

[YB04] Chen Yu and Dana H. Ballard.

A multimodal learning interface for grounding spoken language in sensory perceptions. ACM Transactions on Applied Perception, 1(1):57–80, July 2004. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Yildiz:2013:FAP

[YBC13] Zeynep Cipiloglu Yildiz, Abdullah Bulbul, and Tolga Capin. A framework for applying the principles of depth perception to information visualization. ACM Transactions on Applied Perception, 10(4):19:1–19:??, October 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Yu:2009:PIA

[YCK^{+}09] Insu Yu, Andrew Cox, Min H. Kim, Tobias Ritschel, Thorsten Grosch, Carsten Dachsbacher, and Jan Kautz. Perceptual influence of approximate visibility in indirect illumination. ACM Transactions on Applied Perception, 6(4):24:1– 24:??, September 2009. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Yoshioka:2025:UIF

[YNYK25] Tatsuya Yoshioka, Takuji Narumi, Chi-Lan Yang, and Hideaki Kuzuoka. Understanding the impact of feedback on players’ motivation when experiencing defeat in online competitive games. ACM Transactions on Applied Perception, 22(4):16:1–16:20, October 2025.

CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Yalcin:2025:VES

[YO25] Ipek Yalçin and Nilgün Olguntürk. Visual experience of space: Relating textual descriptions of perceived atmosphere to luminance contrast metrics. ACM Transactions on Applied Perception, 22(3):10:1– 10:??, July 2025. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Yamac:2025:UIV

[YON25] Goksu Yamac, Carol O’Sullivan, and Michael Neff. Understanding the impact of visual and kinematic information on the perception of physicality errors. ACM Transactions on Applied Perception, 22(1):1:1– 1:??, January 2025. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3660636.

Yang:2024:CTE

[YWL^{+}24] Bailin Yang, Tianxiang Wei, Frederick W. B. Li, Xiaohui Liang, Zhigang Deng, and Yili Fang. Color theme evaluation through user preference modeling. ACM Transactions on Applied Perception, 21(3): 12:1–12:??, July 2024. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3665329.

Ziat:2012:EVM

[ZAAC12] Mounia Ziat, Carmen Au, Amin Haji Abolhassani, and James J. Clark. Enhancing visuospatial map learning through action on cellphones. ACM Transactions on Applied Perception, 9(1):5:1– 5:??, March 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Zannoli:2017:PCC

[ZB17] Marina Zannoli and Martin S. Banks. The perceptual consequences of curved screens. ACM Transactions on Applied Perception, 15(1):6:1–6:??, November 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Zana:2006:FRB

[ZC06] Yossi Zana and Roberto M. Cesar, Jr. Face recognition based on polar frequency features. ACM Transactions on Applied Perception, 3(1):62–82, January 2006. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Ziemek:2012:EEO

[ZCRTW12] Tina Ziemek, Sarah Creem-Regehr, William Thompson, and Ross Whitaker. Evaluating the effectiveness of orientation indicators with an awareness of individual differences. ACM Transactions on Applied Perception, 9(2):7:1–7:??, June 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Zibrek:2015:EEM

[ZHRM15] Katja Zibrek, Ludovic Hoyet, Kerstin Ruhland, and Rachel Mcdonnell. Exploring the effect of motion type and emotions on the perception of gender in virtual humans. ACM Transactions on Applied Perception, 12 (3):11:1–11:??, July 2015. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Zhan:2013:MDF

[ZLO13] Ce Zhan, Wanqing Li, and Philip Ogunbona. Measuring the degree of face familiarity based on extended NMF. ACM Transactions on Applied Perception, 10(2):8:1–8:??, May 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Zhang:2019:PBC

[ZLQ^{+}19] Xiao Zhang, Yongqiang Lyu, Tong Qu, Pengfei Qiu, Xiaomin Luo, Jingyu Zhang, Shunjie Fan, and Yuanchun Shi. Photoplethysmogrambased cognitive load assessment using multi-feature fusion model. ACM Transactions on Applied Perception, 16 (4):19:1–19:??, September 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3340962.

Zeng:2024:CHG

[ZLWZ24] Yao Zeng, Xiaoyu Liu, Yijun Wang, and Junsong Zhang. Color hint-guided ink wash painting colorization with ink style prediction mechanism. ACM Transactions on Applied Perception, 21(3):10:1–10:??, July 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3657637.

Zibrek:2019:PIP

[ZMM19] Katja Zibrek, Sean Martin, and Rachel McDonnell. Is photorealism important for perception of expressive virtual humans in virtual reality? ACM Transactions on Applied Perception, 16 (3):14:1–14:??, September 2019. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3349609.

Zibrek:2020:EGA

[ZNO^{+}20] Katja Zibrek, Benjamin Niay, Anne-Hélène Olivier, Ludovic Hoyet, Julien Pettre, and Rachel McDonnell. The effect of gender and attractiveness of motion on proximity in virtual reality. ACM Transactions on Applied Perception, 17(4):14:1– 14:15, December 2020. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic). URL

https://dl.acm.org/doi/10.
1145/3419985.

Zhang:2012:MAV

[ZNWK12] Ruimin Zhang, Anthony Nordman, James Walker, and Scott A. Kuhl. Minification affects verbal- and actionbased distance judgments differently in head-mounted displays. ACM Transactions on Applied Perception, 9(3):14:1– 14:??, July 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Zhang:2015:DFA

[ZOH^{+}15] Tingting Zhang, Louise O’hare, Paul B. Hibbard, Harold T. Nefs, and Ingrid Heynderickx. Depth of field affects perceived depth in stereographs. ACM Transactions on Applied Perception, 11(4):18:1–18:??, January 2015. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Zhang:2024:GSG

[ZWL^{+}24] Yunxiang Zhang, Nan Wu, Connor Z. Lin, Gordon Wetzstein, and Qi Sun. Gaze-Fusion: Saliency-guided image generation. ACM Transactions on Applied Perception, 21 (4):14:1–14:??, October 2024. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3694969.

Zhang:2022:CWP

[ZYJ^{+}22] Junsong Zhang, Zuyi Yang, Linchengyu Jin, Zhitang Lu, and Jinhui Yu. Creating word paintings jointly considering semantics, attention, and aesthetics. ACM Transactions on Applied Perception, 19(3): 13:1–13:??, July 2022. CO-DEN ???? ISSN 1544-3558 (print), 1544-3965 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3539610.

Zhang:2017:CAE

[ZYZ^{+}17] Jiajing Zhang, Jinhui Yu, Kang Zhang, Xianjun Sam Zheng, and Junsong Zhang. Computational aesthetic evaluation of logos. ACM Transactions on Applied Perception, 14(3):20:1– 20:??, July 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Zhao:2013:API

[ZZ13] Mingtian Zhao and Song-Chun Zhu. Abstract painting with interactive control of perceptual entropy. ACM Transactions on Applied Perception, 10(1):5:1– 5:??, February 2013. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).